### v1

# Complexity of deep computations via topology of function spaces

Abstract.
We use topological methods to study complexity of deep computations and limit computations. We use topology of function spaces, specifically, the classification Rosenthal compacta, to identify new complexity classes. We use the language of model theory, specifically, the concept of independence from Shelah’s classification theory, to translate between topology and computation. We use the theory of Rosenthal compacta to characterize approximablility of deep computations, both deterministically and probabilistically.
Key words and phrases:
Deep computations, deep equilibrium models, deep learning, physics-informed networks, computational complexity, independence property, NIP, infinite Ramsey theory, Baire class 1 functions, Rosenthal compacta, Todorčević trichotomy, Bourgain-Fremlin-Talagrand
2000 Mathematics Subject Classification:
54H30, 68T27, 68T07, 03C98, 03D15, 05D10
1Department of Mathematics, University of Texas at San Antonio
2Department of Mathematics, University of Toronto
Introduction
In this paper we study limit behavior of real-valued computations as the values of certain parameters of the computation model tend towards infinity, or towards zero, or towards some other fixed value, e.g., the depth of a neural network tending to infinity, or the time interval between layers of the network tending toward zero. Recently, particular cases of this situation have attracted considerable attention in deep learning research (e.g., Neural Ordinary Differential Equations [Chen:2018], Physics-Informed Neural Networks [Raissi-Perdikaris-Karniadakis:2019], and deep equilibrium models [Bai:2019], among others). In this paper, we combine ideas of topology, measure theory, and model theory to study these limit phenomena from a unified viewpoint.
Informed by model theory, to each computation in a given computation model, we associate a continuous real-valued function, called the type of the computation, that describes the logical properties of this computation with respect to the rest of the model. This allows us to view computations in any given computational model as elements of a space of real-valued functions, which is called the space of types of the model. The idea of embedding models of theories into their type spaces is central in model theory. In the context of this paper, the embedding of computations into spaces of types allows us to utilize the vast theory of topology of function spaces, known as $C_{p}$-theory, to obtain results about complexity of topological limits of computations. As we shall indicate next, recent classification results for spaces of functions provide an elegant and powerful machinery to classify computations according to their levels of “tameness” or “wildness”, with the former corresponding roughly to polynomial approximability and the latter to exponential approximability. The viewpoint of spaces of types, which we have borrowed from model theory, thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: In topology, the classification of Rosenthal compacta pioneered by Todorčević [Todorcevic_1999_CompactSubsetsBaire]; in logic, the classification of theories developed by Shelah [Shelah:1990]; and in statistical learning, the notion of PAC learning and VC dimension pioneered by Vapkins and Chervonenkis [Vapnik-Chervonenkis:1974, Vapnik-Chervonenkis:1971].
In a previous paper [alva2024approximability], we introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [Bai:2019]). There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
In [alva2024approximability], we proved a new “tame vs wild” (i.e., polynomial vs exponential) dichotomy for complexity of deep computations by invoking a classical result of Grothendieck from the 50s [Grothendieck:1952]. Under our model-theoretic Rosetta stone, polynomial approximability in the sense of computation becomes identified with the notion of continuous extendability in the sense of topology, and with the notions of stability and type definability in the sense of model theory.
In this paper, we follow a more general approach, i.e., we view deep computations as pointwise limits of continuous functions. In topology, functions that arise as the pointwise limit of a sequence of continuous functions are called functions of the first Baire class, or Baire class 1 functions, or Baire-1 for short; Baire class 1 form a step above simple continuity in the hierarchy of functions studied in real analysis (Baire class 0 functions being continuous functions). Intuitively, Baire-1 functions represent functions with “controlled” discontinuities, so they are crucial in topology and set theory.
We prove a new “tame vs wild” Ramsey-theoretic dichotomy for complexity of general deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT_1978_PCompactBaire], and a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [Todorcevic_1999_CompactSubsetsBaire].
Todorčević’s trichotomy regards Rosenthal compacta; these are special classes of topological spaces, defined as compact spaces that can be embedded (homeomorphically identified as a subset) within the space of Baire class 1 functions on some Polish (separable, complete metric) space, under the pointwise convergence topology. Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways, and since the late 70’s, they have played a crucial role for understanding complexity of structures of functional analysis, especially Banach spaces. Todorčević’s trichotomy has been utilized to settle longstanding problems in topological dynamics and topological entropy [glasner2022tame].
Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of “Non Independence Property” (known as “NIP”) in model theory, identified by Shelah [Shelah:1971, Shelah:1990], and to the concept of Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory identified by Valiant [Valiant:1984].
Going beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [argyros2008rosenthal]. Argyros, Dodos and Kanellopoulos identified seven fundamental “prototypes” of separable Rosenthal compacta, and proved that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes. They showed that if a separable Rosenthal compactum is not hereditarily separable, then it must contain an uncountable discrete subspace of the size of the continuum.
We believe that the results presented in this paper show practitioners of computation, or topology, or descriptive set theory, or model theory, how classification invariants used in their field translate into classification invariants of other fields. However, in the interest of accessibility, we do not assume previous familiarity with high-level topology or model theory, or computing. The only technical prerequisite of the paper is undergraduate-level topology and measure theory. The necessary topological background beyond undergraduate topology is covered in section 1.
In section 1, we present the basic topological and combinatorial preliminaries, and in section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed). Section 3 is devoted to the classification of deep computations, and the final section, section 4, presents the probabilistic viewpoint.
Throughout the paper, we focus on classical computation; however, by refining the model-theoretic tools, the results presented here can be extended to the realm of quantum computation and open quantum systems. This extension will be addressed in a forthcoming paper.
1. General topological preliminaries: From continuity
to Baire class 1
In this section we present the preliminaries from general topology and function space theory. We include some of the proofs for completeness, but the reader familiar with these topics may skip them.
Recall that a subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets. Note that in a metrizable space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
A Polish space is a separable and completely metrizable topological space. The most important examples are the reals $\mathbb{R}$, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), and the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology). Countable products of Polish spaces are Polish; this includes spaces like $\mathbb{R}^{\mathbb{N}}$, the space of sequences of real numbers.
In this paper, we shall often discuss subspaces, and so there is a pertinent subtlety of the definitions worth mentioning: completely metrizable space is not the same as complete metric space; for an illustrative example, the interval $(0,1)$ with the metric inherited from the reals not complete, but it is Polish since it is homeomorphic to the real line. Being Polish is a topological property while being metrically complete is not.
The following result is a cornerstone of descriptive set theory, closely tied to the work of Wacław Sierpiński and Kazimierz Kuratowski, with proofs often built upon their foundations and formalized later, notably involving Stefan Mazurkiewicz’s work on complete metric spaces.
Fact 1.1.
A subset $A$ of a Polish space $X$ is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set. In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
Given two topological spaces $X$ and $Y$ we denote by $C_{p}(X,Y)$ the set of all continuous functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence. When $Y=\mathbb{R}$, we denote this collection simply as $C_{p}(X)$. A natural question is, how do topological properties of $X$ translate into $C_{p}(X)$ and vice versa? These questions, and in general the study of these spaces, are the concern of $C_{p}$-theory, an active field of research in general topology which was pioneered by A. V. Arhangel’skiĭ and his students in the 1970’s and 1980’s [Arkhangelskii:1992]. This field has found many applications in model theory and functional analysis. For a recent survey, see [tkachuk2011cp].
A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions. If $X$ nd $Y$ are topological spaces, the Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$. As above, in the special case $Y=\mathbb{R}$ we denote $B_{1}(X,Y)$ as $B_{1}(X)$. Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)$. The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles. His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ (equivalently, every open subset of $X$ is a $G_{\delta}$). Note that every metrizable space is perfectly normal.
The following fact was established by Baire in his 1899 thesis. A proof can be found in Section 10 of [Todorcevic_1997_TopicsTop].
Fact 1.2 (Baire).
If $X$ is perfectly normal, then the following conditions are equivalent for a function $f:X\to\mathbb{R}$:
  • $f$ is a Baire class 1 function, that is, $f$ is a pointwise limit of continuous functions.
  • $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ whenever $U\subseteq Y$ is open.
  • For every closed $F\subseteq X$, the restriction $f|_{F}$ has a point of continuity.
Moreover, if $X$ is Polish and $f\notin B_{1}(X)$, then there exists countable $D_{0},D_{1}\subseteq X$ and reals $a<b$ such that
$D_{0}\subseteq f^{-1}(-\infty,a],\quad D_{1}\subseteq f^{-1}[b,\infty),\quad\overline{D_{0}}=\overline{D_{1}}.$
A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact. Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) have been objects of interest for researchers in Analysis and Topological Dynamics. We begin with the following well-known result. Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}>0$ such that $|f(x)|<M_{x}$ for all $f\in A$. We include a proof for the reader’s convenience:
Lemma 1.3.
Let $X$ be a Polish space and $A\subseteq B_{1}(X)$ be pointwise bounded. The following are equivalent:
  1. (i)
    AA is relatively compact in $B_{1}(X)$.
  2. (ii)
    AA is relatively countably compact in $B_{1}(X)$, i.e., every countable subset of $A$ has an accumulation point in $B_{1}(X)$.
  3. (iii)
    $\overline{A}\subseteq B_{1}(X)$, where $\overline{A}$ denotes the closure in $\mathbb{R}^{X}$.
Proof.
Since $A$ is pointwise bounded, for each $x\in X$, fix $M_{x}>0$ such that $|f(x)|\leq M_{x}$ for every $f\in A$.
[…]
By relative countable compactness of $A$, there is an accumulation point $g\in B_{1}(X)$ of $\{f_{n}\}_{n\in\mathbb{N}}$. It is straightforward to show that since $f$ and $g$ agree on $D_{0}\cup D_{1}$, $g$ does not have a point of continuity on the closed set $\overline{D_{0}}=\overline{D_{1}}$, which contradicts Fact 1.2.
[…]
1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
In metrizable spaces, points of closure of some subset can always be approximated by points inside the set, via a convergent sequence. For more complicated spaces, such as $C_{p}(X)$, this fails in remarkable ways. To see an example, consider the Cantor space $X=2^{\mathbb{N}}$, and for each $n\in\mathbb{N}$ define $p_{n}:X\to\{0,1\}$ by $p_{n}(x)=x(n)$ for each $x\in X$. Then $p_{n}$ is continuous for each $n$, but one can show (see Chapter 1.1 of [Todorcevic_1997_TopicsTop] for details) that the only continuous functions in the closure of $\{p_{n}\}_{n\in\mathbb{N}}$ are the functions $p_{n}$ themselves; moreover, none of the subsequences of $\{p_{n}\}_{n\in\mathbb{N}}$ converge. In some sense, this example is the worst possible scenario for convergence. The topological space obtained from this closure is well-known: it is the Stone-Čech compactification of the discrete space of natural numbers, or $\beta\mathbb{N}$ for short, and it is an important object of study in general topology.
The following theorem, established by Haskell Rosenthal in 1974, is fundamental in functional analysis, and describes a sharp division in the behavior of sequences in a Banach spaces. I
Theorem 1.4 (Rosenthal’s Dichotomy, [Rosenthal:1974]).
If $X$ is Polish and $\{f_{n}\}\subseteq C_{p}(X)$ is pointwise bounded, then either $\{f_{n}\}_{n\in\mathbb{N}}$ contains a convergent subsequence or a subsequence whose closure (in $\mathbb{R}^{X}$) is homeomorphic to $\beta\mathbb{N}$.
In other words, a pointwise bounded set of continuous functions either contains a convergent subsequence, or a subsequence whose closure is essentially the same as the example mentioned in the previous paragraphs (the “wildest” possible scenario). Note that in the preceding example, the functions are trivially pointwise bounded in $\mathbb{R}^{X}$ as the functions can only take values 0 and $1$.
The genesis of Theorem 1.4 was Rosenthal’s $\ell_{1}$ theorem, which states that the only reason why Banach space can fail to have an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences) is the presence of a bounded sequence with no weakly Cauchy subsequence. The theorem is famous for connecting diverse areas of mathematics, namely, Banach space geometry, Ramsey theory, set theory, and topology of function spaces.
As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, we find a similar dichotomy: Either every point of closure of the set of functions will be a Baire class 1 function, or there is a sequence inside the set that behaves in the wildest possible way. The theorem is usually not phrased as a dichotomy, but rather as an equivalence:
Theorem 1.5 (“The BFT Dichotomy”. Bourgain-Fremlin-Talagrand [BFT_1978_PCompactBaire, Theorem 4G]).
Let $X$ be a Polish space and $A\subseteq C_{p}(X)$ be pointwise bounded. The following are equivalent:
  1. (i)
    AA is relatively compact in $B_{1}(X)$, i.e., $\overline{A}\subseteq B_{1}(X)$.
  2. (ii)
    For every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that _nIf_n^-1(-,a]∩⋂_nIf_n^-1[b,)=.
Definition 1.6.
We shall say that a set $A\subseteq\mathbb{R}^{X}$ satisfies the Independence Property, or IP for short, if it satisfies the following condition: There exists every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and $a<b$ such that for every pair of disjoint sets $E,F\subseteq\mathbb{N}$, we have
$\bigcap_{n\in E}f_{n}^{-1}(-\infty,a]\cap\bigcap_{n\in F}f_{n}^{-1}[b,\infty)\neq\emptyset.$
If $A$ satisfies the negation of this condition, we will say that $A$ satisfies NIP, or that it has the NIP.
Remark 1.7.
Note that if $X$ is compact and $A\subseteq C_{p}(X)$, then $A$ satisfies the NIP if and only if for every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and for every $a<b$ there is $I\subseteq\mathbb{N}$ such that
$\bigcap_{n\in I}f_{n}^{-1}(-\infty,a]\cap\bigcap_{n\notin I}f_{n}^{-1}[b,\infty)=\emptyset.$
To summarize, the particular case of Theorem 1.5 for $X$ compact can be stated in the following way:
Theorem 1.8.
Let $X$ be a compact Polish space. Then, for every pointwise bounded $A\subseteq C_{p}(X)$, one and exactly one of the following two conditions must hold:
  1. (i)
    $\overline{A}\subseteq B_{1}(X)$.
  2. (ii)
    AA has NIP.
The Independence Property was first isolated by Saharon Shelah in model theory as a dividing line between theories whose models are “tame” (corresponding to NIP) and theories whose models are “wild” (corresponding to IP). See [Shelah:1971, Definition 4.1],[Shelah:1990]. We will discuss this dividing line in more detail in the next section.
1.2. NIP as a universal dividing line between polynomial and exponential complexity
The particular case of the BFT dichotomy (Theorem 1.5) when $A$ consists of $\{0,1\}$-valued (i.e., $\{\text{Yes},\text{No}\}$-valued) strings was discovered independently, around 1971-1972 in many foundational contexts related to polynomial (“tame”) vs exponential (“wild”) complexity: In model theory, by Saharon Shelah [Shelah:1971],[Shelah:1990],in combinatorics, by Norbert Sauer [Sauer:1972], and Shelah [Shelah:1972, Shelah:1990], and in statistical learning, by Vladimir Vapnik and Alexey Chervonenkis [Vapnik-Chervonenkis:1971, Vapnik-Chervonenkis:1974].
In model theory:
Shelah’s classification theory is a foundational program in mathematical logic devised to categorize first-order theories based on the complexity and structure of their models. A theory $T$ is considered classifiable in Shelah’s sense if the number of non-isomorphic models of $T$ of a given cardinality can be described by a bounded number of numerical invariants. In contrast, a theory $T$ is unclassifiable if the number of models of $T$ of a given cardinality is the maximum possible number. A key fact is that the number of models of $T$ is directly impacted by the number of types over sets of parameters in models of $T$; a controlled number of types is a characteristic of a classifiable theory.
In Shelah’s classification program [Shelah:1990], theories without the independence property (called NIP theories, or dependent theories) have a well-behaved, “tame” structure; the number of types over a set of parameters of size $\kappa$ of such a theory is of polynomially or similar “slow” growth on $\kappa$. In contrast, Theories with the Independence Property (called IP theories) are considered “intractable” or “wild”. A theory with the Independence Property produces the maximum possible number of types over a set of parameters; for a set of parameters of cardinality $\kappa$, the theory will have $2^{2^{\kappa}}$-many distinct types.
In combinatorics:
Sauer [Sauer:1972] and Shelah [Shelah:1972] proved the following: If $\mathscr{F}=\{S_{0},S_{1},\dots\}$ is a family of subsets of some infinite set $S$, then either for every $n\in\mathbb{N}$, there is either a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A):i\in\mathbb{N}\}|=2^{n}$ (yielding exponential complexity), or there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
$|\{S_{i}\cap A):i\in\mathbb{N}\}|\leq\sum_{i=0}^{N-1}\binom{|A|}{i}\approx O(|A|^{N})$
(yielding polynomial complexity). This answered a question of Erdős.
In machine learning:
Readers familiar with statistical learning may recognize the Sauer-Shelah lemma as the dichotomy discovered and proved slightly earlier (1971) by Vapknis and Chervonenkis [Vapnik-Chervonenkis:1971, Vapnik-Chervonenkis:1974] to address the problem of uniform convergence in statistics. The least integer $N$ given by the preceding paragraph, when it exists, is called the VC-dimension of $\mathscr{F}$. This is a core concept in machine learning. If such an integer $N$ does not exist, we say that the VC-dimension of $\mathscr{F}$ is infinite. The lemma provides upper bounds on the number of data points (sample size $m$) needed to learn a concept class with VC dimension $d\in\mathbb{N}$ by showing this number grows polynomially with $m$ and $d$ (namely, $\sum_{i=0}^{d}{m\choose i}\approx O(m^{d})$), not exponentially. The Fundamental Theorem of Statistical Learning states that a hypothesis class is PAC-learnable (PAC stands for “Probably Approximately Correct”) if and only if its VC dimension is finite.
1.3. Rosenthal compacta
The comprehensiveness of Theorem 1.5, attested by the examples outlined in the preceding section, led to the following definition (isolated by Gilles Godefroy [Godefroy:1980]):
Definition 1.9.
A Rosenthal compactum is a compact Hausdorff topological space $K$ that can be topologically embedded as a compact subset into the space of all functions of the first Baire class on some Polish space $X$, equipped with the topology of pointwise convergence.
Rosenthal compacta are characterized by significant topological and dynamical tameness properties. They play an important role in functional analysis, measure theory, dynamical systems, descriptive set theory, and model theory. In this paper, we introduce their applicability in deep computation. For this, we shall first focus on countable languages, which is the theme of the next subsection.
1.4. The special case $B_{1}(X,\mathbb{R}^{\mathcal{P}})$ with $\mathcal{P}$ countable.
Our goal now is to characterize relatively compact subsets of $B_{1}(X,Y)$ for the particular case when $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable. Given $P\in\mathcal{P}$ we denote the projection map onto the $P$-coordinate by $\pi_{P}:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}$. From a high-level topological interpretation, the next lemma states that, in this context, the spaces $\mathbb{R}$ and $\mathbb{R}^{\mathcal{P}}$ are really not that different, and that if we understand the Baire class 1 functions of one space, then we also understand the functions of both.
Lemma 1.10.
Let $X$ be a Polish space and $\mathcal{P}$ be a countable set. Then, $f\in B_{1}(X,\mathbb{R}^{\mathcal{P}})$ if and only if $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$.
Proof.
Only one implication needs a proof. Suppose that $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$. Let $V$ be a basic open subset of $\mathbb{R}^{\mathcal{P}}$. That is, there exists a finite $\mathcal{P}^{\prime}\subseteq\mathcal{P}$ such that $V=\bigcap_{P\in\mathcal{P}^{\prime}}\pi_{P}^{-1}[U_{P}]$ where $U_{P}$ is open in $\mathbb{R}$. Then,
$f^{-1}[V]=\bigcap_{P\in\mathcal{P}^{\prime}}(\pi_{P}\circ f)^{-1}[U_{P}]$
is an $F_{\sigma}$ set. Since $\mathcal{P}$ is countable, $\mathbb{R}^{\mathcal{P}}$ is second countable so every open set $U$ in $\mathbb{R}^{\mathcal{P}}$ is a countable union of basic open sets. Hence, $f^{-1}[U]$ is $F_{\sigma}$. ∎
Below we consider $\mathcal{P}$ with the discrete topology. For each $f:X\rightarrow\mathbb{R}^{\mathcal{P}}$ denote $\hat{f}(P,x):=\pi_{P}\circ f(x)$ for all $(P,x)\in\mathcal{P}\times X$. Similarly, for each $g:\mathcal{P}\times X\rightarrow\mathbb{R}$ denote $\check{g}(x)(P):=g(P,x)$. Given $A\subseteq(\mathbb{R}^{\mathcal{P}})^{X}$, we denote $\hat{A}$ as the set of all $\hat{f}$ such that $f\in A$. Note that the map $\left(\mathbb{R}^{\mathcal{P}}\right)^{X}\to\mathbb{R}^{\mathcal{P}\times X}$ given by $f\mapsto\hat{f}$ is a homeomorphism and its inverse is given by $g\mapsto\check{g}$.
Lemma 1.11.
Let $X$ be a Polish space and $\mathcal{P}$ be countable. Then, $f\in B_{1}(X,\mathbb{R}^{\mathcal{P}})$ if and only if $\hat{f}\in B_{1}(\mathcal{P}\times X)$.
Proof.
($\Rightarrow$) By Lemma 1.10, given an open set of reals $U$, we have $f^{-1}[\pi_{P}^{-1}[U]]$ is $F_{\sigma}$ for every $P\in\mathcal{P}$. Given that $\mathcal{P}$ is a discrete countable space, we observe that
$\hat{f}^{-1}[U]=\bigcup_{P\in\mathcal{P}}\left(\{P\}\times f^{-1}[\pi_{P}^{-1}[U]]\right)$
is an $F_{\sigma}$ as well.
($\Leftarrow$) By lemma 1.10 it suffices to show that $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$. Fix an open $U\subseteq\mathbb{R}$. Write $\hat{f}^{-1}[U]=\bigcup_{n\in\mathbb{N}}F_{n}$ where $F_{n}$ is closed in $\mathcal{P}\times X$. Then,
$(\pi_{P}\circ f)^{-1}[U]=\bigcup_{n\in\mathbb{N}}\{x\in X:(P,x)\in F_{n}\}$
which is $F_{\sigma}$. ∎
Given $A\subseteq Y^{X}$ and $K\subseteq X$ we write $A|_{K}:=\{f|_{K}:f\in A\}$, i.e., the set of all restrictions of functions in $A$ to $K$. The following Theorem is a slightly more general version of Theorem 1.5.
Theorem 1.12.
Assume that $\mathcal{P}$ is countable, $X$ is a Polish space, and $A\subseteq C_{p}(X,\mathbb{R}^{\mathcal{P}})$ is such that $\pi_{P}\circ A$ is pointwise bounded for all $P\in\mathcal{P}$. The following are equivalent for every compact $K\subseteq X$:
  1. (i)
    $\overline{A|_{K}}\subseteq B_{1}(K,\mathbb{R}^{\mathcal{P}})$.
  2. (ii)
    $\pi_{P}\circ A|_{K}$ satisfies the NIP for every $P\in\mathcal{P}$.
Proof.
[…]
$(\mathcal{P}\times K)\cap\bigcap_{n\in I}\hat{f_{n}}^{-1}(-\infty,a]\cap\bigcap_{n\notin I}\hat{f_{n}}^{-1}[b,\infty)=\emptyset$
Hence,
$K\cap\bigcap_{n\in I}(\pi_{P}\circ f_{n})^{-1}(-\infty,a]\cap\bigcap_{n\notin I}(\pi_{P}\circ f_{n})^{-1}[b,\infty)=\emptyset$
[…]
$K\cap\bigcap_{n\in E}(\pi_{P}\circ f_{n})^{-1}(-\infty,a]\cap\bigcap_{n\in F}(\pi_{P}\circ f_{n})^{-1}[b,\infty)=\emptyset$
Thus, $\pi_{P}\circ A|_{L}$ satisfies the NIP.
[…]
Lastly, a simple but useful lemma that helps understand when we restrict a set of functions to a specific subspace of the domain space, we may always assume that the subspace is closed, as replacing the subspace by its closure has no effect on NIP.
Lemma 1.13.
Assume that $X$ is Hausdorff and that $A\subseteq C_{p}(X)$. The following are equivalent for every $L\subseteq X$:
  1. (i)
    $A_{L}$ satisfies the NIP.
  2. (ii)
    $A|_{\overline{L}}$ satisfies the NIP.
Proof.
[…]
$\overline{L}\cap\bigcap_{n\in E}f_{n}^{-1}(-\infty,a]\cap\bigcap_{n\in F}f_{n}^{-1}[b,\infty)\neq\emptyset.$
Pick $a^{\prime}<b^{\prime}$ such that $a<a^{\prime}<b^{\prime}<b$. Then, for any finite disjoint $E,F\subseteq\mathbb{N}$ we can choose
$x\in\overline{L}\cap\bigcap_{n\in E}f_{n}^{-1}(-\infty,a^{\prime})\cap\bigcap_{n\in F}f_{n}^{-1}(b^{\prime},\infty)$
By definition of closure:
$L\cap\bigcap_{n\in E}f_{n}^{-1}(-\infty,a^{\prime}]\cap\bigcap_{n\in F}f_{n}^{-1}[b^{\prime},\infty)\neq\emptyset.$
This contradicts (i). ∎
2. Compositional computation structures: A structural approach
to floating-point computation
In this section, we connect function spaces with floating point computation. We start by summarizing some basic concepts from [alva2024approximability].
A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates. For a state $v\in L$, the type of $v$ is defined as the indexed family
$\operatorname{tp}(v)=(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}.$
For each $P\in\mathcal{P}$, we call the value $P(v)$ the $P$-th feature of $v$. A transition of a computation states structure $(L,\mathcal{P})$ is a map $f:L\to L$.
Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable. We think of each state $v\in L$ as being uniquely characterized by its type $\operatorname{tp}(v)$. Thus, in practice, we identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$. A typical case will be when $L=\mathbb{R}^{\mathbb{N}}$ or $L=\mathbb{R}^{n}$ for some positive integer $n$ and there is a predicate $P_{i}(v)=v_{i}$ for each of the coordinates $v_{i}$ of $v$. We regard the space of types as a topological space, endowed with the topology of pointwise convergence inherited from $\mathbb{R}^{\mathcal{P}}$. In particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
Definition 2.1.
Given a computation states structure $(L,\mathcal{P})$, any element of $\mathbb{R}^{\mathcal{P}}$ in the image of $L$ under the map $v\mapsto\operatorname{tp}(v)$ will be called a realized type. The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of types of $(L,\mathcal{P})$, denoted $\mathcal{L}$. Elements of $\mathcal{L}\setminus L$ will be called unrealized types.
In traditional, compact-valued, model theory, the space of types of a structure is viewed as a sort of compactification of the structure, and the compactness of type spaces plays a central role. However, here we are dealing with real-valued structures, and the space $\mathcal{L}$ defined above is not necessarily compact. To bypass this obstacle, we follow the idea introduced in [alva2024approximability] of covering $\mathcal{L}$ by “thin” compact subspaces that we call shards. The formal definition of shard is next.
Definition 2.2.
A sizer is a tuple $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers indexed by $\mathcal{P}$. Given a sizer $r_{\bullet}$, we define the $r_{\bullet}$-shard as:
$L[r_{\bullet}]=L\cap\prod_{P\in\mathcal{P}}[-r_{P},r_{P}].$
For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$. We define $\mathcal{L}_{\text{sh}}$, as the union of all type-shards.
2.1. Compositional Computation Structures
Definition 2.3.
A Compositional Computation Structure (CCS) is a triple $(L,\mathcal{P},\Gamma)$, where
  • $(L,\mathcal{P})$ is a computation states structure, and
  • $\Gamma\subseteq L^{L}$ is a semigroup under composition.
The elements of the semigroup $\Gamma$ are called the computations of the structure $(L,\mathcal{P},\Gamma)$.
If $\Delta\subseteq\Gamma$, we say that $\Delta\subseteq\Gamma$ is $R$-confined if $\gamma|_{L[r_{\bullet}]}:L[r_{\bullet}]\to L[r_{\bullet}]$ for every $r_{\bullet}\in R$ and $\gamma\in\Delta$. Elements in $\overline{\Gamma}\subseteq\mathcal{L}_{\text{sh}}$ are called (real-valued) deep computations or ultracomputations.
For a CCS $(L,\mathcal{P},\Gamma)$, we regard the elements of $\Gamma$ as “standard” finitary computations and the elements of $\overline{\Gamma}$, i.e., deep computations, as possibly infinitary limits of standard computations. The main goal of this paper is to study the computability, definability and computational complexity of deep computations. Since ultracomputations are defined through a combination of topological concepts (namely, topological closure) and structural and model-theoretic concepts (namely, models and types), we will import technology from both topology and model theory.
2.2. Computability and definability of deep computations and the Extendibility Axiom
Let $f:L\to\mathcal{L}$ be a function that maps each input state type $(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ into an output state type $(P\circ f(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$.
  1. (1)
    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{L}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ by a finite number of applications of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
    $|\varphi_{Q,\,K,\varepsilon}(v)-Q{\circ}f(v)|\leq\epsilon,\qquad\text{for all $v\in K$}.$
    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model the theory of real-valued structures first introduced in [Iovino:1994] for model theory of functional analysis and now standard in model theory (see [Keisler:2023]). The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
  2. (2)
    We will say that $f$ is computable if it is definable in the sense defined above under (1), but without the use of the sup/inf operators; in other words, if for every choice of $Q,K,\varepsilon$, the approximation function $\varphi_{Q,K,\varepsilon}:L\to\mathcal{L}$ can be constructed without any use of sup or inf operators. This is quantifier-free definability (i.e., definability as given by the preceding paragraph, but without use of quantifiers), which, from a logic viewpoint, corresponds to computability (the presence of the quantifiers $\exists$ and $\forall$ are the reason behind the undecidability of first-order logic).
It is shown in [alva2024approximability] that:
  1. (1)
    For a definable $f:L\to\mathcal{L}$, the approximating functions $\varphi_{Q,\,K,\varepsilon}$ may be taken to be polynomials of the input features, and
  2. (2)
    Definable transforms $f:L\to\mathcal{L}$ are precisely those that extend to continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$.
To summarize, a function $f:L\to\mathcal{L}$ is computable if and only if it is definable if and only if it is polynomially approximable if and only it can be extended to a continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$. This motivates the following definition.
Definition 2.4.
We say that a CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendability Axiom if for all $\gamma\in\Gamma$, there is $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous. We refer to $\tilde{\gamma}$ as a free extension of $\gamma$.
By the preceding remarks, the Extendability Axiom says that the elements of the semigroup $\Gamma$ are finitary computations. For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$. For any $\Delta\subseteq\Gamma$, let $\tilde{\Delta}$ denote $\{\tilde{\gamma}\colon\gamma\in\Delta\}$.
For a more detailed discussion of the Extendability Axiom, we refer the reader to [alva2024approximability].
For an illustrative example, we can frame Newton’s polynomial root approximation method in the context of a CCS (see Example 5.6 of [alva2024approximability] for details) as follows. Begin by considering the extended complex numbers $\hat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ with the usual Riemann sphere topology that makes it into a compact space (where unbounded sequences converge to $\infty$). In fact, not only is this space compact, but it is covered by the shard given by the sizer $(1,1,1)$ (the unit sphere is contained in the cube $1,1]^{3}$). The space $\hat{\mathbb{C}}$ is homeomorphic to the usual unit sphere $S^{2}:=\{(x,y,z):x^{2}+y^{2}+z^{2}=1\}$ of $\mathbb{R}^{3}$, by means of the stereographic projection and its inverse $\hat{\mathbb{C}}\to S^{2}$. This function is regarded as a triple of predicates $x,y,z:\hat{\mathbb{C}}\to[-1,1]$ where each will map an extended complex number to its corresponding real coordinate on the cube $1,1]^{3}$. Now fix the cubic complex polynomial $p(s):=s^{3}-1$, and consider the map which performs one step in Newton’s method at a particular (extended) complex number $s$, for finding a root of $p$, $\gamma_{p}:\hat{\mathbb{C}}\to\hat{\mathbb{C}}$. The explicit inner workings of $\gamma_{p}$ are irrelevant for this example, except for the fact that it is a continuous mapping. It follows that $(S^{2},\{x,y,z\},\{\gamma_{p}^{k}:k\in\mathbb{N}\})$ is a CCS. The idea is that repeated applications of $\gamma_{p}(s),\gamma_{p}\circ\gamma_{p}(s),\gamma_{p}\circ\gamma_{p}\circ\gamma_{p}(s),\dots$ would approximate a root of $p$ provided $s$ was a good enough initial guess.
3. Classifying deep computations
3.1. NIP, Rosenthal compacta, and deep computations
Under what conditions are deep computations Baire class 1, and thus well-behaved according to our framework, on type-shards? The following theorem says that, under the assumption that $\mathcal{P}$ is countable, the space of deep computations is a Rosenthal compactum (when restricted to shards) if and only if the set of computations satisfies the NIP feature by feature. Hence, we can import the theory of Rosenthal compacta into this framework of deep computations.
Theorem 3.1.
Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.3) satisfying the Extendability Axiom (Definition 2.4) with $\mathcal{P}$ countable. Let $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be $R$-confined. The following are equivalent.
  1. (i)
    $\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$.
  2. (ii)
    $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and $r_{\bullet}\in R$; that is, for all $P\in\mathcal{P}$, $r_{\bullet}\in R$, $a<b$, $\{\gamma_{n}\}_{n\in\mathbb{N}}\subseteq\Delta$ there are finite disjoint $E,F\subseteq\mathbb{N}$ such that L[r_]∩⋂_nE(π_Pγ_n)^-1(-,a]∩⋂_nF(π_Pγ_n)^-1[b,)=.
Moreover, if any (hence all) of the preceding conditions hold, then every deep computation $f\in\overline{\Delta}$ can be extended to a Baire-1 function on shards, i.e., there is $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ such that $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$. In particular, on each shard every deep computation is the pointwise limit of a countable sequence of computations.
Proof.
Since $\mathcal{P}$ is countable, $\mathcal{L}[r_{\bullet}]\subseteq\mathbb{R}^{\mathcal{P}}$ is Polish. Also, the Extendability Axiom implies that $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a pointwise bounded set of continuous functions for all $P\in\mathcal{P}$. Hence, Theorem 1.12 and Lemma 1.13 prove the equivalence of (i) and (ii). If (i) holds and $f\in\overline{\Delta}$, then write $f=\mathcal{U}{\rm lim}_{i}\gamma_{i}$ as an ultralimit. Define $\tilde{f}:=\mathcal{U}{\rm lim}_{i}\tilde{\gamma}_{i}$. Hence, for all $r_{\bullet}\in R$ we have $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$. That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where topological closures coincide with sequential closures, see Theorem 3F in [BFT_1978_PCompactBaire] or Theorem 4.1 in [debs2013rosenthal]). ∎
3.2. The Todorčević trichotomy and levels of PAC learnability
Given a countable set $\Delta$ of computations satisfying the NIP on features and shards (condition (ii) of Theorem 3.1) we have that $\overline{\tilde{\Delta}_{\mathcal{L}[r_{\bullet}]}}$ (for a fixed sizer $r_{\bullet}$) is a separable Rosenthal compactum (see Definition 1.9). Todorčević proved a remarkable trichotomy for Rosenthal compacta [Todorcevic_1999_CompactSubsetsBaire], and later Argyros, Dodos, Kanellopoulos [argyros2008rosenthal] proved an heptachotomy that refined Todorčević’s classification. In this section, inspired by the work of Glasner and Megrelishvili [glasner2022tame], we study ways in which this classification allows us to obtain different levels of PAC-learnability and NIP.
Recall that a topological space $X$ is hereditarily separable if every subspace is separable, and that $X$ is first countable if every point in $X$ has a countable local basis. Every separable metrizable space is hereditarily separable, and R. Pol proved that every hereditarily separable Rosenthal compactum is first countable (see section 10 of [debs2013rosenthal]). This suggests the following definition:
Definition 3.2.
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom and $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be an $R$-confined countable set of computations satisfying the NIP on shards and features (condition (ii) in Theorem 3.1). We say that $\Delta$ is:
  1. (i)
    NIP1 if $\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}$ is first countable for every $r_{\bullet}\in R$.
  2. (ii)
    NIP2 if $\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}$ is hereditarily separable for every $r_{\bullet}\in R$.
  3. (iii)
    NIP3 if $\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}$ is metrizable for every $r_{\bullet}\in R$.
[…]
We now present some separable and non-separable examples of Rosenthal compacta.
Examples 3.3.
  1. (1)
    Alexandroff compactification of a discrete space of size continuum. For each $a\in 2^{\mathbb{N}}$ consider the map $\delta_{a}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ given by $\delta_{a}(x)=1$ if $x=a$ and $\delta_{a}(x)=0$ otherwise. Let $A(2^{\mathbb{N}})=\{\delta_{a}:a\in 2^{\mathbb{N}}\}\cup\{0\}$, where 0 is the zero map. Notice that $A(2^{\mathbb{N}})$ is a compact subset of $B_{1}(2^{\mathbb{N}})$, in fact $\{\delta_{a}:a\in 2^{\mathbb{N}}\}$ is a discrete subspace of $B_{1}(2^{\mathbb{N}})$ and its pointwise closure is precisely $A(2^{\mathbb{N}})$. Hence, this is a Rosenthal compactum which is not first countable. Notice that this space is also not separable.
  2. (2)
    Extended Alexandroff compactification. For each finite binary sequence $s\in 2^{<\mathbb{N}}$, let $v_{s}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $v_{s}(x)=1$ if $x$ extends $s$ and $v_{s}(x)=0$ otherwise. Let $\hat{A}(2^{\mathbb{N}})$ be the pointwise closure of $\{v_{s}:s\in 2^{<\mathbb{N}}\}$, i.e., $\hat{A}(2^{\mathbb{N}})=A(2^{\mathbb{N}})\cup\{v_{s}:s\in 2^{<\mathbb{N}}\}$. Note that this space is a separable Rosenthal compactum which is not first countable.
  3. (3)
    Split Cantor. Let << be the lexicographic order in the space of infinite binary sequences, i.e., $2^{\mathbb{N}}$. For each $a\in 2^{\mathbb{N}}$ let $f_{a}^{-}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{-}(x)=1$ if $x<a$ and $f_{a}^{-}(x)=0$ otherwise. Let $f_{a}^{+}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{+}(x)=1$ if $x\leq a$ and $f_{a}^{+}(x)=0$ otherwise. The split Cantor is the space $S(2^{\mathbb{N}})=\{f_{a}^{-}:a\in 2^{\mathbb{N}}\}\cup\{f_{a}^{+}:a\in 2^{\mathbb{N}}\}$. This is a separable Rosenthal compactum. One example of a countable dense subset is the set of all $f_{a}^{+}$ and $f_{a}^{-}$ where $a$ is an infinite binary sequence that is eventually constant. Moreover, it is hereditarily separable, but it is not metrizable.
  4. (4)
    Alexandroff Duplicate. Let $K$ be any compact metric space and consider the Polish space $X=C(K)\sqcup K$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$. For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={x(a),x∈C(K)0,x∈K g_a^1(x)={x(a),x∈C(K)δ_a(x),x∈K Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$. Notice that $D(K)$ is a first countable Rosenthal compactum. It is not separable if $K$ is uncountable. The interesting case will be when $K=2^{\mathbb{N}}$.
  5. (5)
    Extended Alexandroff Duplicate of the split Cantor. For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s. Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0,x¡a_t1/2,a_t≤x ≤b_t1,b_t¡x Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$. Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable. In fact, it contains an uncountable discrete subspace (see Theorem 5 in [Todorcevic_1999_CompactSubsetsBaire]).
Theorem 3.4 (Todorčević’s Trichotomy, [Todorcevic_1999_CompactSubsetsBaire], Theorem 3 in [argyros2008rosenthal]).
Let $K$ be a separable Rosenthal Compactum.
  1. (i)
    If $K$ is hereditarily separable but non-metrizable, then $S(2^{\mathbb{N}})$ embeds into $K$.
  2. (ii)
    If $K$ is first countable but not hereditarily separable, then either $D(2^{\mathbb{N}})$ or $\hat{D}(S(2^{\mathbb{N}}))$ embeds into $K$.
  3. (iii)
    If $K$ is not first countable, then $A(2^{\mathbb{N}})$ embeds into $K$.
We thus have the following classification:
$K$ is separable Rosenthal compactumKK is metrizableKK is not metrizableKK is hereditarily separable (copy of $S(I)$)        $K$ is not hereditarily separableKK is first countable (copy of $D(2^{\mathbb{N}})$ or $\hat{D}(S(I))$)$K$ is not first countable (copy of $A(2^{\mathbb{N}})$)
The definitions provided here for NIPi ($i=1,2,3$) are topological. This raises the following question:
Question 3.5.
Is there a non-topological characterization for NIPi, $i=1,2,3$?
3.3. The Argyros-Dodos-Kanellopoulos heptachotomy, and approximability of deep computation by minimal classes
In the three separable three cases given in 3.3, namely, ($\hat{A}(2^{\mathbb{N}})$, $S(2^{\mathbb{N}})$ and $\hat{D}(S(2^{\mathbb{N}}))$), the countable dense subsets are indexed by the binary tree $2^{<\mathbb{N}}$. This choice of index is useful for two reasons:
  1. (1)
    Our emphasis is computational. Elements of $2^{<\mathbb{N}}$ represent finite bitstrings, i.e., standard computations, while Rosenthal compacta represent deep computations, i.e., limits of finite computations. Mathematically, deep computations are pointwise limits of standard computations. Howwever, computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
  2. (2)
    The Ramsey theory of perfect subsets of the Cantor space $2^{\mathbb{N}}$ can be imported to analyze the behavior of the accumulation points. Since $2^{<\mathbb{N}}$ is countable, we can always choose this index for the countable dense subsets. This is done in [argyros2008rosenthal].
Definition 3.6.
Let $X$ be a Polish space.
  1. (1)
    If $I$ is a countable and $\{f_{i}:i\in I\}\subseteq\mathbb{R}^{X}$, $\{g_{i}:i\in I\}\subseteq\mathbb{R}^{X}$ are two pointwise families by $I$, we say that $\{f_{i}:i\in\ I\}$ and $\{g_{i}:i\in I\}$ are equivalent if and only if the map $f_{i}\mapsto g_{i}$ is extended to a homeomorphism from $\overline{\{f_{i}:i\in I\}}$ to $\overline{\{g_{i}:i\in I\}}$.
  2. (2)
    If $\{f_{t}:t\in 2^{<\mathbb{N}}\}$ is a pointwise bounded family, we say that $\{f_{t}:t\in 2^{<\mathbb{N}}\}$ is minimal if and only if for every dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$, $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to $\{f_{t}:t\in 2^{<\mathbb{N}}\}$.
One of the main results in [argyros2008rosenthal] is that, up to equivalence, there are seven minimal families of Rosenthal compacta and that for every relatively compact $\{f_{t}:t\in 2^{<\mathbb{N}}\}\subseteq B_{1}(X)$ there is a dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ such that $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to one of the minimal families. We shall describe the seven minimal families next. We follow the same notation as in [argyros2008rosenthal]. For any node $t\in 2^{<\mathbb{N}}$, let us denote by $t^{\frown}0^{\infty}$ ($t^{\frown}1^{\infty}$) the infinite binary sequence starting with $t$ and continuing will all 0’s (respectively, all $1$’s). Fix a regular dyadic subtree $R=\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$ (i.e., a dyadic subtree such that every level of $R$ is contained in a level of $2^{<\mathbb{N}}$) with the property that for all $s,s^{\prime}\in R$, $s^{\frown}0^{\infty}\neq s^{\prime\frown}0^{\infty}$ and $s^{\frown}1^{\infty}\neq s^{\prime\frown}1^{\infty}$. Given $t\in 2^{<\mathbb{N}}$, let $v_{t}$ be the characteristic function of the set $\{x\in 2^{\mathbb{N}}:x\text{ extends }t\}$. Let << be the lexicographic order in $2^{\mathbb{N}}$. Given $a\in 2^{\mathbb{N}}$, let $f^{+}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a\leq x\}$ and let $f^{-}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a<x\}$. Given two maps $f,g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ we denote by $(f,g):2^{\mathbb{N}}\sqcup 2^{\mathbb{N}}\rightarrow\mathbb{R}$ the function which is $f$ on the first copy of $2^{\mathbb{N}}$ and $g$ on the second copy of $2^{\mathbb{N}}$.
  1. (1)
    $D_{1}=\{\frac{1}{|t|+1}v_{t}:t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{1}}=A(2^{\mathbb{N}})$.
  2. (2)
    $D_{2}=\{s_{t}^{\frown}0^{\infty}:t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{2}}=2^{\leq N}$.
  3. (3)
    $D_{3}=\{f^{+}_{s_{t}^{\frown}0^{\infty}}:t\in 2^{<\mathbb{N}}\}$. This is a discrete in $\overline{D_{3}}=S(2^{\mathbb{N}})$.
  4. (4)
    $D_{4}=\{f^{-}_{s_{t}^{\frown}1^{\infty}}:t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{4}}=S(2^{\mathbb{N}})$.
  5. (5)
    $D_{5}=\{v_{t}:t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{5}}=\hat{A}(2^{\mathbb{N}})$.
  6. (6)
    $D_{6}=\{(v_{s_{t}},s_{t}^{\frown}0^{\infty}):t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{6}}=\hat{D}(2^{\mathbb{N}})$.
  7. (7)
    $D_{7}=\{(v_{s_{t}},x^{+}_{s_{t}^{\frown}0^{\infty}}):t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{7}}=\hat{D}(S(2^{\mathbb{N}}))$
Theorem 3.7 (Heptachotomy of minimal families, Theorem 2 in [argyros2008rosenthal]).
Let $X$ be Polish. For every relatively compact $\{f_{t}:t\in 2^{<\mathbb{N}}\}\subseteq B_{1}(X)$, there exists $i=1,2,\dots,7$ and a regular dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$ such that $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to $D_{i}$. Moreover, all $D_{i}$ are minimal and mutually non-equivalent.
4. Measure-theoretic versions of NIP and universal Monte Carlo computability of deep computations
The countability assumption on $\mathcal{P}$ played a crucial role in the proof of Theorem 1.12 , as it makes $\mathbb{R}^{\mathcal{P}}$ a Polish space. In this section, we do not assume that $\mathcal{P}$ is countable. We replace deterministic computability by measure-theoretic (‘Monte Carlo’) computability.
4.1. A measure-theoretic version of NIP
Recall that the raison d’être of the class of Baire-1 functions is to have a class that contains the continuous functions but is closed under pointwise limits, and that for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$ (see Fact 1.2). This motivates the following definition:
Definition 4.1.
Given a Hausdorff space $X$ and a measurable space $(Y,\Sigma)$, we say that $f:X\rightarrow Y$ is universally measurable (with respect to $\Sigma$) if $f^{-1}(E)$ is Borel for every $E\in\Sigma$, i.e., $f^{-1}(E)$ is $\mu$-measurable for every Radon measure $\mu$ on $X$. When $Y=\mathbb{R}$ we will always take $\Sigma=\mathcal{B}(\mathbb{R})$, the Borel $\sigma$-algebra of $\mathbb{R}$.
Remark 4.2.
A function $f:X\rightarrow\mathbb{R}$ is universally measurable if and only if $f^{-1}(U)$ is $\mu$-measurable for every Radon probability measure $\mu$ on $X$ and every open set $U\subseteq\mathbb{R}$.
Intuitively, a function is universally measurable if it is “measurable no matter which reasonable way you try to measure things on its domain”. The concept of universal measurability emerged from work of Kallianpur and Sazonov, in the late 1950’s and 1960s, with later developments by Blackwell, Darst, and others, building on earlier ideas of Gnedenko and Kolmogorov from the 1950s. See [Pap:2002, Chapters 1 and 2].
Notation 4.3.
Following [BFT_1978_PCompactBaire], the collection of all universally measurable real-valued functions will be denoted by $M_{r}(X)$.
In the context of deep computations, we will be interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself. There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra, i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$. Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller. We will use the cylinder $\sigma$-algebra to define universally measurable maps $f:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$. The reason for this choice is the following characterization:
Lemma 4.4.
Let $X$ be a Hausdorff space and $Y=\prod_{i\in I}Y_{i}$ be any product of measurable spaces $(Y_{i},\Sigma_{i})$ for $i\in I$. Let $\Sigma_{Y}$ be the cylinder $\sigma$-algebra generated by the measurable spaces $(Y_{i},\Sigma_{i})$. Let $f:X\rightarrow Y$. The following are equivalent:
  1. (i)
    $f:X\rightarrow Y$ is universally measurable (with respect to $\Sigma_{Y}$).
  2. (ii)
    $\pi_{i}\circ f:X\rightarrow Y_{i}$ is universally measurable (with respect to $\Sigma_{i}$) for all $i\in I$.
Proof.
[…]
The preceding lemma says that a transition map is universally measurable if and only if it is universally measurable on all its features; in other words, we can check measurability of a transition just by checking measurability feature by feature. We will denote by $M_{r}(X,\mathbb{R}^{\mathcal{P}})$ the collection of all universally measurable functions $f:X\rightarrow\mathbb{R}^{\mathcal{P}}$ (with respect to the cylinder $\sigma$-algebra), endowed with the topology of pointwise convergence.
We will need the following result about NIP and universally measurable functions:
Theorem 4.5 (Bourgain-Fremlin-Talagrand, Theorem 2F in [BFT_1978_PCompactBaire]).
Let $X$ be a Hausdorff space and $A\subseteq C(X)$ be pointwise bounded. The following are equivalent:
  1. (i)
    $\overline{A}\subseteq M_{r}(X)$.
  2. (ii)
    For every compact $K\subseteq X$, $A|_{K}$ satisfies the NIP.
  3. (iii)
    For every Radon measure $\mu$ on $X$, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$, i.e., every countable subset of $A$ has an accumulation point in $\mathscr{M}^{0}(X,\mu)$.
4.2. Universal Monte Carlo computability of deep computations
We now wish to define the concept of a deep computation being computable except on a set of arbitrarily small measure “no matter which reasonable way you try to measure things on its domain” (see the remarks following definition 4.1). This is the concept of universal Monte Carlo computability defined below (Definition 4.6). To motivate the definition, we need to recall two facts:
  1. (1)
    Littlewoood’s second principle states that every Lebesgue measurable function is “nearly continuous”. The formal version of this, which is Luzin’s theorem, states that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ be a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable ) equipped with a Borel algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|F$ is continuous.
  2. (2)
    Computability of deep computations can is characterized in terms of continuous extendibility of computations. This is at the core of [alva2024approximability].
These two facts motivate the following definition:
Definition 4.6.
Let $(L,\mathcal{P},\Gamma)$ be a CCS. We say that a transition $f:L\rightarrow L$ is universally Monte Carlo computable if and only if there exists $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ extending $f$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that the restriction $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow\mathcal{L}[s_{\bullet}]$ is universally measurable, i.e., $\pi_{P}\circ\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow[-s_{P},s_{P}]$ is $\mu$-measurable for every Radon probability measure $\mu$ on $\mathcal{L}[r_{\bullet}]$.
4.3. Bourgain-Fremlin-Talagrand, NIP, and universal Monte Carlo computability of deep computations
Theorem 4.5 immediately yields the following.
Theorem 4.7.
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom. Let $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be $R$-confined. If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and all $r_{\bullet}\in R$, then every deep computation is universally Monte Carlo computable.
Proof.
By the Extendability Axiom, Theorem 4.5 and Lemma 1.13 we have $\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$. Let $f\in\overline{\Delta}$ be a deep computation. Write $f=\mathcal{U}\lim_{i}\gamma_{i}$ as an ultralimit of computations in $\Delta$. Define $\tilde{f}:=\mathcal{U}\lim_{i}\tilde{\gamma}_{i}$. Then, for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$ $\pi_{P}\circ\tilde{\gamma}_{i}|_{\mathcal{L}[r_{\bullet}]}\in M_{r}(\mathcal{L}[r_{\bullet}])$ for all $i$, so $\pi_{P}\circ f|_{\mathcal{L}[r_{\bullet}]}\in\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$. ∎
Question 4.8.
Under the same assumptions of the preceding theorem, suppose that every deep computation of $\Delta$ is universally Monte Carlo computable. Must $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ have the NIP for all $P\in\mathcal{P}$ and all $r_{\bullet}\in R$?
4.4. Talagrand stability, Fremlin’s dichotomy, NIP, and universal Monte Carlo computability of deep computations
There is another notion closely related to NIP, introduced by Talagrand in [talagrand1984pettis] while studying Pettis integration. Suppose that $X$ is a compact Hausdorff space and $A\subseteq\mathbb{R}^{X}$. Let $\mu$ be a Radon probability measure on $X$. Given a $\mu$-measurable set $E\subseteq X$, a positive integer $k$ and real numbers $a<b$. we write:
$D_{k}(A,E,a,b)=\bigcup_{f\in A}\{x\in E^{2k}:f(x_{2i})\leq a,\hskip 2.84526ptf(x_{2i+1})\geq b\hskip 2.84526pt{\text{ for all }i<k}\}$
We say that $A$ is Talagrand $\mu$-stable if and only if for every $\mu$-measurable set $E\subseteq X$ of positive measure and for every $a<b$ there is $k\geq 1$ such that
$(\mu^{2k})^{*}(D_{k}(A,E,a,b))<(\mu(E))^{2k},$
where $\mu*$ denotes the outer measure (we work with outer since the sets $D_{k}(A,E,a,b)$ need not be $\mu$-measurable). This is certainly the case when $A$ is a countable set of continuous (or $\mu$-measurable) functions.
Notation 4.9.
For a measure $\mu$ on a set $X$, the set of all $\mu$-measurable functions will denoted by $\mathscr{M}^{0}(X,\mu)$.
The following lemma establishes that Talagrand stability is a way to ensure that deep computations are definable by measurable functions.
Lemma 4.10.
If $A$ is Talagrand $\mu$-stable, then $\overline{A}$ is also Talagrand $\mu$-stable and $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$.
Proof.
First, observe that a subset of a $\mu$-stable set is $\mu$-stable. To show that $\overline{A}$ is $\mu$-stable, observe that $D_{k}(\overline{A},E,a,b)\subseteq D_{k}(A,E,a^{\prime},b^{\prime})$ where $a<a^{\prime}<b^{\prime}<b$ and $E$ is a $\mu$-measurable set with positive measure. It suffices to show that $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$. Suppose that there exists $f\in\overline{A}$ such that $f\notin\mathscr{M}^{0}(X,\mu)$. By a characterization of measurable functions (see 413G in [fremlin2003vol4]), there exists a $\mu$-measurable set $E$ of positive measure and $a<b$ such that $\mu^{*}(P)=\mu^{*}(Q)=\mu(E)$ where $P=\{x\in E:f(x)\leq a\}$ and $Q=\{x\in E:f(x)\geq b\}$. Then, for any $k\geq 1$: $(P\times Q)^{k}\subseteq D_{k}(\{f\},E,a,b)$ so $(\mu^{2k})^{*}(D_{k}(\{f\},E,a,b))=(\mu^{*}(P)\mu^{*}(Q))^{k}=(\mu(E))^{2k}$. Thus, $\{f\}$ is not $\mu$-stable, but we argued before that a subset of a $\mu$-stable set must be $\mu$-stable. ∎
We say that $A$ is universally Talagrand stable if $A$ is Talagrand $\mu$-stable for every Radon probability measure $\mu$ on $X$. An argument similar to the proof of 4.5, yields the following:
Theorem 4.11.
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom. If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ is universally Talagrand stable for all $P\in\mathcal{P}$ and all sizers $r_{\bullet}$, then every deep computation is universally Monte Carlo computable.
It is then natural to ask: what is the relationship between Talagrand stability and the NIP? The following dichotomy will be useful.
Lemma 4.12 (Fremlin’s Dichotomy, 463K in [fremlin2003vol4]).
If $(X,\Sigma,\mu)$ is a perfect $\sigma$-finite measure space (in particular, for $X$ compact and $\mu$ a Radon probability measure on $X$) and $\{f_{n}:n\in\mathbb{N}\}$ be a sequence of real-valued measurable functions on $X$, then one (and only one) of the following conditions holds:
  1. (i)
    $\{f_{n}:n\in\mathbb{N}\}$ has a subsequence that converges $\mu$-almost everywhere,
  2. (ii)
    $\{f_{n}:n\in\mathbb{N}\}$ has a subsequence with no $\mu$-measurable accumulation point in $\mathbb{R}^{X}$.
The preceding lemma can be considered as a measure-theoretic version of Rosenthal’s dichotomy. Combining this dichotomy with Theorem 4.5, we get the following result:
Theorem 4.13.
Let $X$ be a Hausdorff space and $A\subseteq C(X)$ be pointwise bounded. The following are equivalent:
  1. (i)
    $\overline{A}\subseteq M_{r}(X)$.
  2. (ii)
    For every compact $K\subseteq X$, $A|_{K}$ satisfies the NIP.
  3. (iii)
    For every Radon measure $\mu$ on $X$, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$, i.e., every countable subset of $A$ has an accumulation point in $\mathscr{M}^{0}(X,\mu)$.
  4. (iv)
    For every Radon measure $\mu$ on $X$ and every sequence $\{f_{n}:n\in\mathbb{N}\}\subseteq A$, there is a subsequence that converges $\mu$-almost everywhere.
Proof.
Notice that the equivalence (i)-(iii) is Theorem 4.5. Notice that the equivalence of (iii) and (iv) is Fremlin’s Dichotomy (Theorem 4.12). ∎
Finally, it is natural to ask what the connection is between Talagrand stability and NIP.
Proposition 4.14.
Let $X$ be a compact Hausdorff space and $A\subseteq C(X)$ be pointwise bounded. If $A$ is universally Talagrand stable, then $A$ satisfies the NIP.
Proof.
By Theorem 4.5, it suffices to show that $A$ is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$ for all Radon probability measure $\mu$ on $X$. Since $A$ is Talagrand $\mu$-stable for any such $\mu$, we have $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$. In particular, $A$ is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$. ∎
Question 4.15.
Is the converse true?
The following two results suggest that the precise connection between Talagrand stability and NIP may be sensitive to set-theoretic axioms (even assuming countability of $A$).
Theorem 4.16 (Talagrand, Theorem 9-3-1(a) in [talagrand1984pettis]).
Let $X$ be a compact Hausdorff space and $A\subseteq M_{r}(X)$ be countable and pointwise bounded. Assume that $[0,1]$ is not the union of $\mathfrak{c}$ closed measure zero sets. If $A$ satisfies the NIP, then $A$ is universally Talagrand stable.
Theorem 4.17 (Fremlin, Shelah, [fremlin1993pointwise]).
It is consistent with the usual axioms of set theory that there exists a countable pointwise bounded set of Lebesgue measurable functions with the NIP which is not Talagrand stable with respect to Lebesgue measure.

### v2

@@
 Abstract.
-We use topological methods to study complexity of deep computations and limit computations.
-We use topology of function spaces, specifically, the classification Rosenthal compacta, to identify new complexity classes.
+We use topological methods to study the complexity of deep computations and limit computations.
+We use topology of function spaces, specifically, the classification of Rosenthal compacta, to identify new complexity classes.
 We use the language of model theory, specifically, the concept of independence from Shelah’s classification theory, to translate between topology and computation.
-We use the theory of Rosenthal compacta to characterize approximablility of deep computations, both deterministically and probabilistically.
+We use the theory of Rosenthal compacta to characterize approximability of deep computations, both deterministically and probabilistically.
 Key words and phrases:
@@
 Introduction
-In this paper we study limit behavior of real-valued computations as the values of certain parameters of the computation model tend towards infinity, or towards zero, or towards some other fixed value, e.g., the depth of a neural network tending to infinity, or the time interval between layers of the network tending toward zero.
-Recently, particular cases of this situation have attracted considerable attention in deep learning research (e.g., Neural Ordinary Differential Equations [Chen:2018], Physics-Informed Neural Networks [Raissi-Perdikaris-Karniadakis:2019], and deep equilibrium models [Bai:2019], among others).
-In this paper, we combine ideas of topology, measure theory, and model theory to study these limit phenomena from a unified viewpoint.
+In this paper we study asymptotic behavior of computations, e.g., the depth of a neural network tending to infinity, or the time interval between layers of a time-series network tending toward zero.
+Recently, particular cases of this concept have attracted considerable attention in deep learning research (e.g., Neural Ordinary Differential Equations [8], Physics-Informed Neural Networks [RPK19], and deep equilibrium models [4]).
+The formal framework introduced here provides a unified setting to study these limit phenomena from a foundational viewpoint.
 Informed by model theory, to each computation in a given computation model, we associate a continuous real-valued function, called the type of the computation, that describes the logical properties of this computation with respect to the rest of the model.
@@
 As we shall indicate next, recent classification results for spaces of functions provide an elegant and powerful machinery to classify computations according to their levels of “tameness” or “wildness”, with the former corresponding roughly to polynomial approximability and the latter to exponential approximability.
-The viewpoint of spaces of types, which we have borrowed from model theory, thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: In topology, the classification of Rosenthal compacta pioneered by Todorčević [Todorcevic_1999_CompactSubsetsBaire]; in logic, the classification of theories developed by Shelah [Shelah:1990]; and in statistical learning, the notion of PAC learning and VC dimension pioneered by Vapkins and Chervonenkis [Vapnik-Chervonenkis:1974, Vapnik-Chervonenkis:1971].
-In a previous paper [alva2024approximability], we introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [Bai:2019]).
+The viewpoint of spaces of types, which we have borrowed from model theory, thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: In topology, the classification of Rosenthal compacta pioneered by Todorčević [TOD99]; in logic, the classification of theories developed by Shelah [SHE90]; and in statistical learning, the notions of PAC learning and VC dimension pioneered by Vapnik and Chervonenkis [VC74, VC71].
+In a previous paper [ADI+24], we introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]).
 There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
-In [alva2024approximability], we proved a new “tame vs wild” (i.e., polynomial vs exponential) dichotomy for complexity of deep computations by invoking a classical result of Grothendieck from the 50s [Grothendieck:1952].
-Under our model-theoretic Rosetta stone, polynomial approximability in the sense of computation becomes identified with the notion of continuous extendability in the sense of topology, and with the notions of stability and type definability in the sense of model theory.
-In this paper, we follow a more general approach, i.e., we view deep computations as pointwise limits of continuous functions.
-In topology, functions that arise as the pointwise limit of a sequence of continuous functions are called functions of the first Baire class, or Baire class 1 functions, or Baire-1 for short; Baire class 1 form a step above simple continuity in the hierarchy of functions studied in real analysis (Baire class 0 functions being continuous functions).
-Intuitively, Baire-1 functions represent functions with “controlled” discontinuities, so they are crucial in topology and set theory.
-We prove a new “tame vs wild” Ramsey-theoretic dichotomy for complexity of general deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT_1978_PCompactBaire], and a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [Todorcevic_1999_CompactSubsetsBaire].
-Todorčević’s trichotomy regards Rosenthal compacta; these are special classes of topological spaces, defined as compact spaces that can be embedded (homeomorphically identified as a subset) within the space of Baire class 1 functions on some Polish (separable, complete metric) space, under the pointwise convergence topology.
-Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways, and since the late 70’s, they have played a crucial role for understanding complexity of structures of functional analysis, especially Banach spaces.
-Todorčević’s trichotomy has been utilized to settle longstanding problems in topological dynamics and topological entropy [glasner2022tame].
-Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of “Non Independence Property” (known as “NIP”) in model theory, identified by Shelah [Shelah:1971, Shelah:1990], and to the concept of Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory identified by Valiant [Valiant:1984].
-Going beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [argyros2008rosenthal].
-Argyros, Dodos and Kanellopoulos identified seven fundamental “prototypes” of separable Rosenthal compacta, and proved that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
-They showed that if a separable Rosenthal compactum is not hereditarily separable, then it must contain an uncountable discrete subspace of the size of the continuum.
+In [ADI+24], we proved a new “tame vs wild” (i.e., polynomially approximable vs nonpolynimially approximable) dichotomy for the complexity of deep computations by invoking a classical result of Grothendieck from the 1950s [GRO52].
+Under our model-theoretic Rosetta stone, the property of polynomial approximability of computations is identified with continuous extendibility in the sense of topology, and with the notions of stability and type definability in model theory.
+Simplest among deep computations are those arising as pointwise limits of (continuous) computations proper.
+In topology, the first Baire class, or Baire class 1 consists of functions (also called simply “Baire-1”) arising as pointwise limits of sequences of continuous functions.
+Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity away from the Baire class 0 which (by definition) consists of continuous functions.
+We prove a new “tame vs wild” Ramsey-theoretic dichotomy for complexity of general deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT78], and a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [TOD99].
+Todorčević’s trichotomy regards Rosenthal compacta; these are special classes of topological spaces, defined as compact spaces that can be embedded (homeomorphically identified as a subset) within the space of Baire class 1 functions on some Polish (separable, completely metrizable) space, under the topology of pointwise convergence.
+Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways; since the late 70’s, they have played a crucial role in understanding the complexity of structures of functional analysis, especially Banach spaces.
+Todorčević’s trichotomy has been utilized to settle longstanding problems in topological dynamics and topological entropy [GM22].
+It is noteworthy that Todorčević’s proof relies on sophisticated set-theoretic forcing and infinite Ramsey theory.
+At the time of writing this paper, decades after his original argument, no elementary proof has been found [TOD23, HT19].
+Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of “Non Independence Property” (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concept of Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory identified by Valiant [VAL84].
+Going beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]; they identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
+They also showed that if a separable Rosenthal compactum is not hereditarily separable, then it must contain an uncountable discrete subspace of the size of the continuum.
 We believe that the results presented in this paper show practitioners of computation, or topology, or descriptive set theory, or model theory, how classification invariants used in their field translate into classification invariants of other fields.
-However, in the interest of accessibility, we do not assume previous familiarity with high-level topology or model theory, or computing.
-The only technical prerequisite of the paper is undergraduate-level topology and measure theory.
+In the interest of accessibility, we do not assume the reader to have previous familiarity with advanced topology, model theory, or computing.
+The only technical prerequisites to read this paper are undergraduate-level topology and measure theory.
 The necessary topological background beyond undergraduate topology is covered in section 1.
-In section 1, we present the basic topological and combinatorial preliminaries, and in section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed).
-Section 3 is devoted to the classification of deep computations, and the final section, section 4, presents the probabilistic viewpoint.
-Throughout the paper, we focus on classical computation; however, by refining the model-theoretic tools, the results presented here can be extended to the realm of quantum computation and open quantum systems.
-This extension will be addressed in a forthcoming paper.
-1. General topological preliminaries: From continuity
-to Baire class 1
-In this section we present the preliminaries from general topology and function space theory. We include some of the proofs for completeness, but the reader familiar with these topics may skip them.
+In section 1, we present basic topological and combinatorial preliminaries, and in section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed).
+Section 3 is devoted to the classification of deep computations.
+Throughout the paper, our results pertain to classical models of computation (particularly computations involving real-valued quantities that are known and manipulated to a finite degree of precision).
+The final section, Section 4, introduces a probabilistic viewpoint, the development of which we intend to pursue in future research, extending the present framework to encompass non-deterministic and quantum computations.
+1. General topological preliminaries: From continuity to Baire class 1
+In this section we present some preliminaries from general topology and function space theory.
+In the interest of completeness, we some proofs that may be safely skipped by readers familiar with these topics.
 Recall that a subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets.
-Note that in a metrizable space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
-A Polish space is a separable and completely metrizable topological space.
-The most important examples are the reals $\mathbb{R}$, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), and the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology).
-Countable products of Polish spaces are Polish; this includes spaces like $\mathbb{R}^{\mathbb{N}}$, the space of sequences of real numbers.
-In this paper, we shall often discuss subspaces, and so there is a pertinent subtlety of the definitions worth mentioning: completely metrizable space is not the same as complete metric space; for an illustrative example, the interval $(0,1)$ with the metric inherited from the reals not complete, but it is Polish since it is homeomorphic to the real line.
-Being Polish is a topological property while being metrically complete is not.
-The following result is a cornerstone of descriptive set theory, closely tied to the work of Wacław Sierpiński and Kazimierz Kuratowski, with proofs often built upon their foundations and formalized later, notably involving Stefan Mazurkiewicz’s work on complete metric spaces.
+A space is metrizable if its topology agrees with the topology induced by some metric therein.
+Two such metrics inducing the same topology may induce quite different properties in the category of metric spaces.
+For example, the interval $(0,1)$ with the usual metric (as a subset) of the reals is not complete; however, $(0,1)$ is homeomorphic to the real line, which is complete with respect to the usual metric thereon.
+In a metrizable space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
+A Polish space is a separable and completely metrizable topological space, i.e., admitting some complete metric inducing its topology.
+Although other (possibly incomplete) metrics may induce the same topology, being Polish is a purely topological property.
+One of the most important Polish spaces is the real line $\mathbb{R}$; the others include the Cantor space $2^{\mathbb{N}}$ and the Baire space $\mathbb{N}^{\mathbb{N}}$.
+The class of Polish spaces is closed under countable topological products; in particular, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology), and the space $\mathbb{R}^{\mathbb{N}}$ of sequences of real numbers are all Polish.
+Recall that the product topology on these spaces is the topology of pointwise convergence: a sequence converges in the space if and only if it converges at each coordinate index.
 Fact 1.1.
-A subset $A$ of a Polish space $X$ is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set.
-In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
+A subset of a Polish space is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set. In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
+For a proof, see [ENG89, 4.3.24].
 Given two topological spaces $X$ and $Y$ we denote by $C_{p}(X,Y)$ the set of all continuous functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence.
-When $Y=\mathbb{R}$, we denote this collection simply as $C_{p}(X)$.
+The space $C_{p}(X,\mathbb{R})$ of continuous real functions on $X$ is denoted simply $C_{p}(X)$.
 A natural question is, how do topological properties of $X$ translate into $C_{p}(X)$ and vice versa?
-These questions, and in general the study of these spaces, are the concern of $C_{p}$-theory, an active field of research in general topology which was pioneered by A.
+This general question, and the study of these spaces in general, is the concern of $C_{p}$-theory, an active field of research in general topology which was pioneered by A.
 V.
-Arhangel’skiĭ and his students in the 1970’s and 1980’s [Arkhangelskii:1992].
+Arhangel’skiĭ and his students in the 1970’s and 1980’s [ARK92].
 This field has found many applications in model theory and functional analysis.
-For a recent survey, see [tkachuk2011cp].
+For a recent survey, see [TKA11].
 A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions.
-If $X$ nd $Y$ are topological spaces, the Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$.
-As above, in the special case $Y=\mathbb{R}$ we denote $B_{1}(X,Y)$ as $B_{1}(X)$.
+If $X$ and $Y$ are topological spaces, the space of Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$ (as above, $B_{1}(X,\mathbb{R})$ is denoted $B_{1}(X)$).
 Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)$.
@@
 His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
-A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ (equivalently, every open subset of $X$ is a $G_{\delta}$).
-Note that every metrizable space is perfectly normal.
-The following fact was established by Baire in his 1899 thesis. A proof can be found in Section 10 of [Todorcevic_1997_TopicsTop].
+A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ (equivalently, every open subset of $X$ is a $F_{\sigma}$).
+Every metrizable space (hence, every Polish space) is perfectly normal.
+A topological space $X$ is Baire if every countable intersection of dense open sets is dense.
+The Baire Category Theorem states that every Hausdorff compact or completely metrizable space (hence, every Polish space) is Baire.
+The following fact was established by Baire in his 1899 thesis. A proof can be found in [TOD97, Section 10].
 Fact 1.2 (Baire).
 If $X$ is perfectly normal, then the following conditions are equivalent for a function $f:X\to\mathbb{R}$:
+  • (1)
     $f$ is a Baire class 1 function, that is, $f$ is a pointwise limit of continuous functions.
+  • (2)
     $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ whenever $U\subseteq Y$ is open.
+If, mooreover, $X$ is Baire, then (1) and (2) are equivalent to:
+  • (3)
     For every closed $F\subseteq X$, the restriction $f|_{F}$ has a point of continuity.
-Moreover, if $X$ is Polish and $f\notin B_{1}(X)$, then there exists countable $D_{0},D_{1}\subseteq X$ and reals $a<b$ such that
+Moreover, if $X$ is Polish and $f\notin B_{1}(X)$, then there exist countable $D_{0},D_{1}\subseteq X$ and reals $a<b$ such that
 $D_{0}\subseteq f^{-1}(-\infty,a],\quad D_{1}\subseteq f^{-1}[b,\infty),\quad\overline{D_{0}}=\overline{D_{1}}.$
 A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact.
-Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) have been objects of interest for researchers in Analysis and Topological Dynamics.
+Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) are of interest analysis and topological dynamics.
 We begin with the following well-known result.
-Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}>0$ such that $|f(x)|<M_{x}$ for all $f\in A$.
+Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}\geq 0$ (a pointwise bound at $x$) such that $|f(x)|\leq M_{x}$ for all $f\in A$.
 We include a proof for the reader’s convenience:
@@
 Proof.
-Since $A$ is pointwise bounded, for each $x\in X$, fix $M_{x}>0$ such that $|f(x)|\leq M_{x}$ for every $f\in A$.
 […]
 […]
-By relative countable compactness of $A$, there is an accumulation point $g\in B_{1}(X)$ of $\{f_{n}\}_{n\in\mathbb{N}}$.
-It is straightforward to show that since $f$ and $g$ agree on $D_{0}\cup D_{1}$, $g$ does not have a point of continuity on the closed set $\overline{D_{0}}=\overline{D_{1}}$, which contradicts Fact 1.2.
 […]
 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
-In metrizable spaces, points of closure of some subset can always be approximated by points inside the set, via a convergent sequence.
-For more complicated spaces, such as $C_{p}(X)$, this fails in remarkable ways.
-To see an example, consider the Cantor space $X=2^{\mathbb{N}}$, and for each $n\in\mathbb{N}$ define $p_{n}:X\to\{0,1\}$ by $p_{n}(x)=x(n)$ for each $x\in X$.
-Then $p_{n}$ is continuous for each $n$, but one can show (see Chapter 1.1 of [Todorcevic_1997_TopicsTop] for details) that the only continuous functions in the closure of $\{p_{n}\}_{n\in\mathbb{N}}$ are the functions $p_{n}$ themselves; moreover, none of the subsequences of $\{p_{n}\}_{n\in\mathbb{N}}$ converge.
-In some sense, this example is the worst possible scenario for convergence.
-The topological space obtained from this closure is well-known: it is the Stone-Čech compactification of the discrete space of natural numbers, or $\beta\mathbb{N}$ for short, and it is an important object of study in general topology.
-The following theorem, established by Haskell Rosenthal in 1974, is fundamental in functional analysis, and describes a sharp division in the behavior of sequences in a Banach spaces. I
-Theorem 1.4 (Rosenthal’s Dichotomy, [Rosenthal:1974]).
-If $X$ is Polish and $\{f_{n}\}\subseteq C_{p}(X)$ is pointwise bounded, then either $\{f_{n}\}_{n\in\mathbb{N}}$ contains a convergent subsequence or a subsequence whose closure (in $\mathbb{R}^{X}$) is homeomorphic to $\beta\mathbb{N}$.
-In other words, a pointwise bounded set of continuous functions either contains a convergent subsequence, or a subsequence whose closure is essentially the same as the example mentioned in the previous paragraphs (the “wildest” possible scenario).
-Note that in the preceding example, the functions are trivially pointwise bounded in $\mathbb{R}^{X}$ as the functions can only take values 0 and $1$.
-The genesis of Theorem 1.4 was Rosenthal’s $\ell_{1}$ theorem, which states that the only reason why Banach space can fail to have an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences) is the presence of a bounded sequence with no weakly Cauchy subsequence.
-The theorem is famous for connecting diverse areas of mathematics, namely, Banach space geometry, Ramsey theory, set theory, and topology of function spaces.
-As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, we find a similar dichotomy: Either every point of closure of the set of functions will be a Baire class 1 function, or there is a sequence inside the set that behaves in the wildest possible way.
-The theorem is usually not phrased as a dichotomy, but rather as an equivalence:
-Theorem 1.5 (“The BFT Dichotomy”. Bourgain-Fremlin-Talagrand [BFT_1978_PCompactBaire, Theorem 4G]).
+In metrizable spaces, points of the closure of some subset can always be approximated by points in the set proper, via a convergent sequence.
+For more complicated spaces, such as $C_{p}$-spaces, this fails in remarkable ways.
+The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$, and one can show (e.g., [TOD97, Chapter 1.1]) that $\{p_{n}\}_{n\in\mathbb{N}}$ is closed, but has no convergent subsequences, in $Y=C_{p}(X)$: in a sense, such example exhibits the worst failure of sequential convergence possible.
+The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) realizes the Stone-Čech compactification $\beta\mathbb{N}$ of the discrete space of natural numbers; it is an important object of study in general topology.
+The following theorem, proved by Haskell Rosenthal in 1974, is fundamental in functional analysis and captures a sharp division in the behavior of sequences in a Banach spaces.
+Theorem 1.4 (Rosenthal’s Dichotomy, [ROS74]).
+If $X$ is Polish and $\{f_{n}\}\subseteq C_{p}(X)$ is pointwise bounded, then $\{f_{n}\}_{n\in\mathbb{N}}$ has a convergent subsequence, or a subsequence whose closure in $\mathbb{R}^{X}$ is homeomorphic to $\beta\mathbb{N}$.
+Rosenthal’s Dichotomy states that a pointwise bounded set of continuous functions contains either a convergent subsequence, or a subsequence whose closure is essentially the same as the example mentioned in the previous paragraphs (i.e., “wildest” possible).
+The genesis of this theorem was Rosenthal’s $\ell_{1}$-Theorem”, which states that a Banach space includes an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences), or else every bounded sequence therein is weakly Cauchy.
+The $\ell_{1}$-Theorem connects diverse areas: Banach space geometry, Ramsey theory, set theory, and topology of function spaces.
+As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, a dichotomy paralleling the $\ell_{1}$-Theorem holds: Either every point of the closure of a set of functions is a Baire class 1 function, or there is a sequence in the set behaving in the wildest possible way.
+This result is usually not phrased as a dichotomy, but rather as an equivalence as in Theorem 1.5 below.
+First, we introduce some useful notation. For any set $A\subseteq\mathbb{R}^{X}$ and any real $a$, define
+[…]
+[…]
+(In case $A=\emptyset$, we define $X^{\emptyset}_{\geq a}=X=X^{\emptyset}_{\leq a}$.) For any sequence $\{f_{n}\}\subseteq\mathbb{R}^{X}$ and $I\subseteq\mathbf{N}$, define $I^{\complement}\coloneqq\mathbb{N}\setminus I$ and $f_{I}\coloneqq\{f_{i}:i\in I\}$.
+Theorem 1.5 (“The BFT Dichotomy”. Bourgain-Fremlin-Talagrand [BFT78]).
 Let $X$ be a Polish space and $A\subseteq C_{p}(X)$ be pointwise bounded. The following are equivalent:
   1. (i)
-    AA is relatively compact in $B_{1}(X)$, i.e., $\overline{A}\subseteq B_{1}(X)$.
+    AA is relatively compact in $B_{1}(X)$.
   2. (ii)
-    For every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that _nIf_n^-1(-,a]∩⋂_nIf_n^-1[b,)=.
+    For every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{I}}\cap X_{\geq b}^{f_{I^{\complement}}}=\emptyset$.
+[…]
+The sets $X_{\leq a}^{f_{I}}$ and $X_{\geq b}^{f_{I^{\complement}}}$ appearing in condition Theorem 1.5(ii) are defined, respectively, in terms of $|I|$-many inequalities of the form $f_{i}(x)\leq a$, and $|I^{\complement}|$-many of the form $f_{j}(x)\geq b$.
+Thus, at least one of $X_{\leq a}^{f_{I}}$ and $X_{\geq b}^{f_{I^{\complement}}}$ is defined by the satisfaction of infinitely (countably) many inequalities.
+For our purposes, it is more natural to consider only finitely many inequalities at a time, which motivates the definitions below.
 Definition 1.6.
-We shall say that a set $A\subseteq\mathbb{R}^{X}$ satisfies the Independence Property, or IP for short, if it satisfies the following condition: There exists every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and $a<b$ such that for every pair of disjoint sets $E,F\subseteq\mathbb{N}$, we have
-$\bigcap_{n\in E}f_{n}^{-1}(-\infty,a]\cap\bigcap_{n\in F}f_{n}^{-1}[b,\infty)\neq\emptyset.$
-If $A$ satisfies the negation of this condition, we will say that $A$ satisfies NIP, or that it has the NIP.
-Remark 1.7.
-Note that if $X$ is compact and $A\subseteq C_{p}(X)$, then $A$ satisfies the NIP if and only if for every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and for every $a<b$ there is $I\subseteq\mathbb{N}$ such that
-$\bigcap_{n\in I}f_{n}^{-1}(-\infty,a]\cap\bigcap_{n\notin I}f_{n}^{-1}[b,\infty)=\emptyset.$
-To summarize, the particular case of Theorem 1.5 for $X$ compact can be stated in the following way:
+We say that a function collection $A\subseteq\mathbb{R}^{X}$ has the finitary No-Independence Property (NIP) if, for all sequences $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$, there exist finite disjoint sets $E,F\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}=\emptyset$.
+We say that such $E,F$ witness finitary NIP for $A$, $\{f_{n}\}$ and $a,b$.
+A set $A\subseteq\mathbb{R}^{X}$ has the finitary Independence Property (IP) if it does not have finitary NIP, i.e., if there exists a sequence $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$ such that for every pair of finite disjoint sets $E,F\subseteq\mathbb{N}$, we have $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}\neq\emptyset$.
+If the word “finite” is omitted in the above definitions, we obtain the definitions of countable NIP (weaker than finitary NIP) and countable IP (stronger than finitary IP), respectively.
+If we insist on witnesses $E,F\subseteq\mathbb{N}$ such that $F=E^{\complement}$, we call the respective properties “BFT-NIP” (even weaker than countable NIP) and “BFT-IP” (even stronger than countable IP).
+Thus, Theorem 1.5 becomes that statement, for pointwise bounded function collections $A\subseteq C_{p}(X)$, that $A$ is relatively compact in $B_{1}(X)$ if and only if $A$ has BFT-NIP.
+Unless otherwise unspecified, IP/NIP shall mean finitary IP/NIP henceforth.
+Proposition 1.7.
+If $X$ is compact and $A\subseteq C_{p}(X)$, then $A$ has BFT-NIP if and only if it has finitary NIP.
+(No pointwise boundedness is assumed of $A$.)
+Proof.
+Trivially (as per the preceding discussion), finitary NIP implies BFT-NIP.
+Reciprocally, assume that $X$ is compact and has finitary IP.
+Fix $A\subseteq C_{p}(X)$, a sequence $\{f_{n}\}\subseteq A$ and reals $r<s$.
+For any $I,J\subseteq\mathbb{N}$ (eventually disjoint in applications), write $X_{I,J}$ for $X^{f_{I}}_{\leq r}\cap X^{f_{J}}_{\geq s}$.
+For $I\subseteq I^{\prime}\subseteq\mathbb{N}$ and $J\subseteq J^{\prime}\subseteq\mathbb{N}$, we have $X_{I,J}\supseteq X_{I^{\prime},J^{\prime}}$; moreover, $X_{I,J}=\bigcap_{E\subseteq I,F\subseteq J}X_{E,F}$, where the index variables $E\subseteq I$, $F\subseteq J$ range over finite subsets of $I,J$, respectively.
+Clearly, $E,F\subseteq\mathbb{N}$ witness finitary NIP for $\{f_{n}\}$ if and only if $X^{E,F}=\emptyset$.
+Fix $I\subseteq\mathbb{N}$.
+Since $\{f_{n}\}\subseteq A\subseteq C_{p}(X)$ is a sequence of continuous functions, and $X$ is compact, the nested family $\{X_{E,F}:E\subseteq I,F\subseteq I^{\complement}\}$ consists of closed, thus compact, sets.
+Since $A$ has finitary IP by hypothesis, the nested family consists of nonempty sets, hence its intersection $X_{I,I^{\complement}}\neq\emptyset$ by compactness.
+This holds for arbitrary $\{f_{n}\}\subseteq A$ and $r<s$, so $A$ has BFT-IP.
+∎
 Theorem 1.8.
-Let $X$ be a compact Polish space. Then, for every pointwise bounded $A\subseteq C_{p}(X)$, one and exactly one of the following two conditions must hold:
+Let $X$ be a metrizable compact (hence Polish) space. For every pointwise bounded $A\subseteq C_{p}(X)$, the following properties are all equivalent:
   1. (i)
-    $\overline{A}\subseteq B_{1}(X)$.
+    AA is relatively compact in $B_{1}(X)$;
   2. (ii)
-    AA has NIP.
+    AA has BFT-NIP;
+  3. (iii)
+    AA has countable NIP;
+  4. (iv)
+    AA has finitary NIP.
+[…]
+Proof.
+Corollary of Theorem 1.5 and Proposition 1.7. ∎
+Theorem 1.8 may be stated as the following dichotomy (under the assumptions): either $A$ is relatively compact in $B_{1}(X)$, or $A$ has IP (in either sense).
 The Independence Property was first isolated by Saharon Shelah in model theory as a dividing line between theories whose models are “tame” (corresponding to NIP) and theories whose models are “wild” (corresponding to IP).
-See [Shelah:1971, Definition 4.1],[Shelah:1990].
+See [SHE71, Definition 4.1], [SHE90].
 We will discuss this dividing line in more detail in the next section.
 1.2. NIP as a universal dividing line between polynomial and exponential complexity
-The particular case of the BFT dichotomy (Theorem 1.5) when $A$ consists of $\{0,1\}$-valued (i.e., $\{\text{Yes},\text{No}\}$-valued) strings was discovered independently, around 1971-1972 in many foundational contexts related to polynomial (“tame”) vs exponential (“wild”) complexity: In model theory, by Saharon Shelah [Shelah:1971],[Shelah:1990],in combinatorics, by Norbert Sauer [Sauer:1972], and Shelah [Shelah:1972, Shelah:1990], and in statistical learning, by Vladimir Vapnik and Alexey Chervonenkis [Vapnik-Chervonenkis:1971, Vapnik-Chervonenkis:1974].
+The particular case of the BFT dichotomy (Theorem 1.5) when $A$ consists of $\{0,1\}$-valued (i.e., $\{\text{Yes},\text{No}\}$-valued) strings was discovered independently, around 1971-1972 in many foundational contexts related to polynomial (“tame”) vs exponential (“wild”) complexity: In model theory, by Saharon Shelah [SHE71], [SHE90], in combinatorics, by Norbert Sauer [SAU72], and Shelah [SHE72], [SHE90], and in statistical learning, by Vladimir Vapnik and Alexey Chervonenkis [VC71], [VC74].
 In model theory:
@@
 A key fact is that the number of models of $T$ is directly impacted by the number of types over sets of parameters in models of $T$; a controlled number of types is a characteristic of a classifiable theory.
-In Shelah’s classification program [Shelah:1990], theories without the independence property (called NIP theories, or dependent theories) have a well-behaved, “tame” structure; the number of types over a set of parameters of size $\kappa$ of such a theory is of polynomially or similar “slow” growth on $\kappa$.
+In Shelah’s classification program [SHE90], theories without the independence property (called NIP theories, or dependent theories) have a well-behaved, “tame” structure; the number of types over a set of parameters of size $\kappa$ of such a theory is of polynomially or similar “slow” growth on $\kappa$.
 In contrast, theories with the Independence Property (called IP theories) are considered “intractable” or “wild”.
@@
 In combinatorics:
-Sauer [Sauer:1972] and Shelah [Shelah:1972] proved the following: If $\mathscr{F}=\{S_{0},S_{1},\dots\}$ is a family of subsets of some infinite set $S$, then either for every $n\in\mathbb{N}$, there is either a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A):i\in\mathbb{N}\}|=2^{n}$ (yielding exponential complexity), or there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
-$|\{S_{i}\cap A):i\in\mathbb{N}\}|\leq\sum_{i=0}^{N-1}\binom{|A|}{i}\approx O(|A|^{N})$
-(yielding polynomial complexity). This answered a question of Erdős.
+Sauer [SAU72] and Shelah [SHE72] proved the following independently: Let $\mathscr{F}$ be a family of subsets of some set $S$.
+Either: for every $n\in\mathbb{N}$ there is a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A:i\in\mathbb{N}\}|=2^{n}$ ($\mathscr{F}$ has “exponential complexity”); or: there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
+$|\{S_{i}\cap A:i\in\mathbb{N}\}|\leq\sum_{i=0}^{N-1}\binom{|A|}{i}$.
+($\mathscr{F}$ has “polynomial complexity”).
+Clearly, any family $\mathscr{F}$ of subsets of a finite set $S$ has polynomial complexity.
+The “polynomial” name is justified: indeed, for fixed $N>0$, as a function of the size $|A|=m>0$, we have
+$\sum_{i=0}^{N-1}\binom{m}{i}\leq\sum_{i=0}^{N-1}\frac{m^{i}}{i!}\leq\left(\sum_{i=0}^{N-1}\frac{1}{i!}\right)\cdot m^{N-1}\cdot<e\cdot m^{N-1}=O\left(m^{N}\right).$
+(More precisely, the order of magnitude is $O(m^{N-1})$: polynomial in $m$ for $N$ fixed.)
 In machine learning:
-Readers familiar with statistical learning may recognize the Sauer-Shelah lemma as the dichotomy discovered and proved slightly earlier (1971) by Vapknis and Chervonenkis [Vapnik-Chervonenkis:1971, Vapnik-Chervonenkis:1974] to address the problem of uniform convergence in statistics.
-The least integer $N$ given by the preceding paragraph, when it exists, is called the VC-dimension of $\mathscr{F}$.
-This is a core concept in machine learning.
+Readers familiar with statistical learning may recognize the Sauer-Shelah lemma as the dichotomy discovered and proved slightly earlier (1971) by Vapnik and Chervonenkis [VC71, VC74] to address uniform convergence in statistics.
+The least integer $N$ given by the preceding paragraph, when it exists, is called the VC-dimension of $\mathscr{F}$; it is a core concept in machine learning.
 If such an integer $N$ does not exist, we say that the VC-dimension of $\mathscr{F}$ is infinite.
-The lemma provides upper bounds on the number of data points (sample size $m$) needed to learn a concept class with VC dimension $d\in\mathbb{N}$ by showing this number grows polynomially with $m$ and $d$ (namely, $\sum_{i=0}^{d}{m\choose i}\approx O(m^{d})$), not exponentially.
+The lemma provides upper bounds on the number of data points (sample size) needed to learn a concept class of known VC dimension $d$ up to a given admissible error in the statistical sense.
 The Fundamental Theorem of Statistical Learning states that a hypothesis class is PAC-learnable (PAC stands for “Probably Approximately Correct”) if and only if its VC dimension is finite.
 1.3. Rosenthal compacta
-The comprehensiveness of Theorem 1.5, attested by the examples outlined in the preceding section, led to the following definition (isolated by Gilles Godefroy [Godefroy:1980]):
+The universal classification implied by Theorem 1.5, as attested by the examples outlined in the preceding section, led to the following definition (by Gilles Godefroy [GOD80]):
 Definition 1.9.
-A Rosenthal compactum is a compact Hausdorff topological space $K$ that can be topologically embedded as a compact subset into the space of all functions of the first Baire class on some Polish space $X$, equipped with the topology of pointwise convergence.
-Rosenthal compacta are characterized by significant topological and dynamical tameness properties.
-They play an important role in functional analysis, measure theory, dynamical systems, descriptive set theory, and model theory.
-In this paper, we introduce their applicability in deep computation.
+A Rosenthal compactum is any topological space realized as a compact subset of the space $B_{1}(X)=B_{1}(X,\mathbb{R})$ (equipped with the topology of pointwise convergence) of all real functions of the first Baire class on some Polish space $X$.
+A Rosenthal compactum $K$ is necessarily Hausdorff since it is a topological subspace of the Hausdorff product space $\mathbb{R}^{X}$.
+Rosenthal compacta possess significant topological and dynamical tameness properties, and play an important role in functional analysis, measure theory, dynamical systems, descriptive set theory, and model theory.
+In this paper, we use them to study deep computations.
 For this, we shall first focus on countable languages, which is the theme of the next subsection.
 1.4. The special case $B_{1}(X,\mathbb{R}^{\mathcal{P}})$ with $\mathcal{P}$ countable.
-Our goal now is to characterize relatively compact subsets of $B_{1}(X,Y)$ for the particular case when $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable.
-Given $P\in\mathcal{P}$ we denote the projection map onto the $P$-coordinate by $\pi_{P}:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}$.
-From a high-level topological interpretation, the next lemma states that, in this context, the spaces $\mathbb{R}$ and $\mathbb{R}^{\mathcal{P}}$ are really not that different, and that if we understand the Baire class 1 functions of one space, then we also understand the functions of both.
-Lemma 1.10.
-Let $X$ be a Polish space and $\mathcal{P}$ be a countable set. Then, $f\in B_{1}(X,\mathbb{R}^{\mathcal{P}})$ if and only if $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$.
+Fix an arbitrary (at most) countable set $\mathcal{P}$ whose elements $P\in\mathcal{P}$ will be called predicate symbols or formal predicates.
+Our present goal is to characterize relatively compact subsets of $B_{1}(X,\mathbb{R}^{\mathcal{P}})$, where $X$ is always assumed to be a perfectly normal space (often a Polish space).
+The set $\mathcal{P}$ shall be considered discrete whenever regarded as a topological space.
+Since $C_{p}(X,\mathbb{R}^{\mathcal{P}})\subseteq B_{1}(X,\mathbb{R}^{\mathcal{P}})\subseteq(\mathbb{R}^{\mathcal{P}})^{X}$, the “ambient” space $(\mathbb{R}^{\mathcal{P}})^{X}$ is quite relevant.
+The product $X\times\mathcal{P}$ will be regarded as either a pointset, or as a topological product depending on context.
+We have natural homeomorphic identifications
+$(\mathbb{R}^{\mathcal{P}})^{X}\cong\mathbb{R}^{X\times\mathcal{P}}\cong(\mathbb{R}^{X})^{\mathcal{P}},$
+given by
+$\mathbb{R}^{X\times\mathcal{P}}\to(\mathbb{R}^{\mathcal{P}})^{X}:\varphi$ $\mapsto{\hat{}}\varphi$
+$\mathbb{R}^{X\times\mathcal{P}}\to(\mathbb{R}^{X})^{\mathcal{P}}:\varphi$ $\mapsto\varphi{\,\hat{}}$,
+where
+$\hat{}\varphi(x)\coloneqq\varphi(x,\cdot)\in\mathbb{R}^{\mathcal{P}},$ $\varphi\,\hat{}(P)\coloneqq\varphi(\cdot,P)\in\mathbb{R}^{X}.$
+Such identifications view $X$ and $\mathcal{P}$ as mere pointsets (the topology of $X$ in particular plays no role).
+For $x\in X$, define the “left projection” map
+$\lambda_{x}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{\mathcal{P}}:\varphi\mapsto\lambda_{x}(\varphi)\coloneqq\varphi(x,\cdot);$
+for $P\in\mathcal{P}$, the “right projection” map
+$\rho_{P}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{X}:\varphi\mapsto\varphi(\cdot,P).$
+For fixed $x\in X$ and $P\in\mathcal{P}$, we also have canonical projection maps
+$\pi_{x}:\mathbb{R}^{X}\to\mathbb{R}:f\mapsto f(x),$ $\pi_{P}:\mathbb{R}^{\mathcal{P}}\to\mathbb{R}:f\mapsto f(P).$
+[…]
+The Proposition below reduces the study of $\mathbb{R}^{\mathcal{P}}$-valued continuous or Baire-1 functions on $X$ to the special case of real-valued ones.
+Proposition 1.10.
+The identification $(\mathbb{R}^{\mathcal{P}})^{X}\cong\mathbb{R}^{X\times\mathcal{P}}\cong(\mathbb{R}^{X})^{\mathcal{P}}$ induces identifications
+$C_{p}(X,\mathbb{R}^{\mathcal{P}})\cong C_{p}(X\times\mathcal{P})\cong C_{p}(X)^{\mathcal{P}},$ $B_{1}(X,\mathbb{R}^{\mathcal{P}})\cong B_{1}(X\times\mathcal{P})\cong B_{1}(X)^{\mathcal{P}}.$
+(The cardinality of $\mathcal{P}$ plays no role.)
 Proof.
-Only one implication needs a proof.
-Suppose that $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$.
-Let $V$ be a basic open subset of $\mathbb{R}^{\mathcal{P}}$.
-That is, there exists a finite $\mathcal{P}^{\prime}\subseteq\mathcal{P}$ such that $V=\bigcap_{P\in\mathcal{P}^{\prime}}\pi_{P}^{-1}[U_{P}]$ where $U_{P}$ is open in $\mathbb{R}$.
-Then,
-$f^{-1}[V]=\bigcap_{P\in\mathcal{P}^{\prime}}(\pi_{P}\circ f)^{-1}[U_{P}]$
-is an $F_{\sigma}$ set.
-Since $\mathcal{P}$ is countable, $\mathbb{R}^{\mathcal{P}}$ is second countable so every open set $U$ in $\mathbb{R}^{\mathcal{P}}$ is a countable union of basic open sets.
-Hence, $f^{-1}[U]$ is $F_{\sigma}$.
+The identification of $C_{p}$-spaces follows trivially from the definition of topological product and the fact that $\mathcal{P}$ is discrete: a continuous map $X\to\mathbb{R}^{\mathcal{P}}$ is precisely a $\mathcal{P}$-tuple of continuous functions $X\to\mathbb{R}$, and these correspond to continuous functions $X\times\mathcal{P}\to\mathbb{R}$.
+The identification of Baire-1 spaces follows immediately, since it is defined in terms of the purely topological notion of limit (in the ambient space) of sequences of continuous functions.
 ∎
-Below we consider $\mathcal{P}$ with the discrete topology.
-For each $f:X\rightarrow\mathbb{R}^{\mathcal{P}}$ denote $\hat{f}(P,x):=\pi_{P}\circ f(x)$ for all $(P,x)\in\mathcal{P}\times X$.
-Similarly, for each $g:\mathcal{P}\times X\rightarrow\mathbb{R}$ denote $\check{g}(x)(P):=g(P,x)$.
-Given $A\subseteq(\mathbb{R}^{\mathcal{P}})^{X}$, we denote $\hat{A}$ as the set of all $\hat{f}$ such that $f\in A$.
-Note that the map $\left(\mathbb{R}^{\mathcal{P}}\right)^{X}\to\mathbb{R}^{\mathcal{P}\times X}$ given by $f\mapsto\hat{f}$ is a homeomorphism and its inverse is given by $g\mapsto\check{g}$.
-Lemma 1.11.
-Let $X$ be a Polish space and $\mathcal{P}$ be countable. Then, $f\in B_{1}(X,\mathbb{R}^{\mathcal{P}})$ if and only if $\hat{f}\in B_{1}(\mathcal{P}\times X)$.
-Proof.
-($\Rightarrow$) By Lemma 1.10, given an open set of reals $U$, we have $f^{-1}[\pi_{P}^{-1}[U]]$ is $F_{\sigma}$ for every $P\in\mathcal{P}$.
-Given that $\mathcal{P}$ is a discrete countable space, we observe that
-$\hat{f}^{-1}[U]=\bigcup_{P\in\mathcal{P}}\left(\{P\}\times f^{-1}[\pi_{P}^{-1}[U]]\right)$
-is an $F_{\sigma}$ as well.
-($\Leftarrow$) By lemma 1.10 it suffices to show that $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$.
-Fix an open $U\subseteq\mathbb{R}$.
-Write $\hat{f}^{-1}[U]=\bigcup_{n\in\mathbb{N}}F_{n}$ where $F_{n}$ is closed in $\mathcal{P}\times X$.
-Then,
-$(\pi_{P}\circ f)^{-1}[U]=\bigcup_{n\in\mathbb{N}}\{x\in X:(P,x)\in F_{n}\}$
-which is $F_{\sigma}$. ∎
 Given $A\subseteq Y^{X}$ and $K\subseteq X$ we write $A|_{K}:=\{f|_{K}:f\in A\}$, i.e., the set of all restrictions of functions in $A$ to $K$.
@@
 Theorem 1.11.
-Assume that $\mathcal{P}$ is countable, $X$ is a Polish space, and $A\subseteq C_{p}(X,\mathbb{R}^{\mathcal{P}})$ is such that $\pi_{P}\circ A$ is pointwise bounded for all $P\in\mathcal{P}$.
+Assume that $\mathcal{P}$ is countable, $X$ is a Polish space, and $A\subseteq C_{p}(X,\mathbb{R}^{\mathcal{P}})$ is pointwise bounded in the sense that $\pi_{P}\circ A$ ($\subseteq C_{p}(X)$) is pointwise bounded for every $P\in\mathcal{P}$.
 The following are equivalent for every compact $K\subseteq X$:
   1. (i)
-    $\overline{A|_{K}}\subseteq B_{1}(K,\mathbb{R}^{\mathcal{P}})$.
+    $A|_{K}$ is relatively compact in $B_{1}(K,\mathbb{R}^{\mathcal{P}})$;
   2. (ii)
-    $\pi_{P}\circ A|_{K}$ satisfies the NIP for every $P\in\mathcal{P}$.
+    $\pi_{P}\circ A|_{K}$ has NIP for every $P\in\mathcal{P}$.
 Proof.
-[…]
-$(\mathcal{P}\times K)\cap\bigcap_{n\in I}\hat{f_{n}}^{-1}(-\infty,a]\cap\bigcap_{n\notin I}\hat{f_{n}}^{-1}[b,\infty)=\emptyset$
-Hence,
-$K\cap\bigcap_{n\in I}(\pi_{P}\circ f_{n})^{-1}(-\infty,a]\cap\bigcap_{n\notin I}(\pi_{P}\circ f_{n})^{-1}[b,\infty)=\emptyset$
-[…]
-$K\cap\bigcap_{n\in E}(\pi_{P}\circ f_{n})^{-1}(-\infty,a]\cap\bigcap_{n\in F}(\pi_{P}\circ f_{n})^{-1}[b,\infty)=\emptyset$
-Thus, $\pi_{P}\circ A|_{L}$ satisfies the NIP.
-[…]
+Compact subsets $K\subseteq X$ are closed, hence also Polish. Therefore, the asserted equivalence follows from Theorems 1.5 and 1.7. ∎
 Lastly, a simple but useful lemma that helps understand when we restrict a set of functions to a specific subspace of the domain space, we may always assume that the subspace is closed, as replacing the subspace by its closure has no effect on NIP.
@@
 This contradicts (i). ∎
-2. Compositional computation structures: A structural approach
-to floating-point computation
-In this section, we connect function spaces with floating point computation. We start by summarizing some basic concepts from [alva2024approximability].
+2. Compositional Computation Structures: A structural approach to floating-point computations
+In this section, we connect function spaces with floating point computations. We start by summarizing some basic concepts from [ADI+24].
 A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates.
-For a state $v\in L$, the type of $v$ is defined as the indexed family
+For a state $v\in L$, the type of a state $v$ is the indexed family
 $\operatorname{tp}(v)=(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}.$
@@
 Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable.
-We think of each state $v\in L$ as being uniquely characterized by its type $\operatorname{tp}(v)$.
-Thus, in practice, we identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$.
-A typical case will be when $L=\mathbb{R}^{\mathbb{N}}$ or $L=\mathbb{R}^{n}$ for some positive integer $n$ and there is a predicate $P_{i}(v)=v_{i}$ for each of the coordinates $v_{i}$ of $v$.
-We regard the space of types as a topological space, endowed with the topology of pointwise convergence inherited from $\mathbb{R}^{\mathcal{P}}$.
-In particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
+Each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$.
+Important space states are $L=\mathbb{R}^{\mathbb{N}}$ and $L=\mathbb{R}^{n}$ for some positive integer $n$, endowed with predicate $P_{i}(v)=v_{i}$, one each for the $i$-th coordinate of $v$.
+We regard the space of types as a topological space, endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$.
+Via the identification $v\mapsto\operatorname{tp}(v)$, the states space $L$ is correspondingly topologized; in particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
 Definition 2.1.
-Given a computation states structure $(L,\mathcal{P})$, any element of $\mathbb{R}^{\mathcal{P}}$ in the image of $L$ under the map $v\mapsto\operatorname{tp}(v)$ will be called a realized type.
-The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of types of $(L,\mathcal{P})$, denoted $\mathcal{L}$.
+Given a computation states structure $(L,\mathcal{P})$, any element of $\mathbb{R}^{\mathcal{P}}$ in the image of $L$ under the map $v\mapsto\operatorname{tp}(v)$ will be called a realized (state) type.
+The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of types of $(L,\mathcal{P})$, denoted $\mathcal{L}$; elements $\xi\in\mathcal{L}$ are called state types.
 Elements of $\mathcal{L}\setminus L$ will be called unrealized types.
-In traditional, compact-valued, model theory, the space of types of a structure is viewed as a sort of compactification of the structure, and the compactness of type spaces plays a central role.
-However, here we are dealing with real-valued structures, and the space $\mathcal{L}$ defined above is not necessarily compact.
-To bypass this obstacle, we follow the idea introduced in [alva2024approximability] of covering $\mathcal{L}$ by “thin” compact subspaces that we call shards.
-The formal definition of shard is next.
+Intuitively, state types capture a notion of “ultrastate”.
+The traditional setting of discrete first-order logic may be regarded as the special case in which predicates are binary (such predicates may take only the values 0 and $1$, say).
+In such a traditional setting, the space of types of a structure is necessarily compact; for some purposes, it serves as a sort of compactification of the structure itself.
+Compactness in such setting is critical.
+However, predicates of a computational states structure take values in the non-compact space $\mathbb{R}$, so the corresponding type space $\mathcal{L}$ need not be compact.
+To bypass this obstacle, we follow the idea, introduced in [ADI+24], of covering $\mathcal{L}$ by “thin” compact subspaces called shards.
 Definition 2.2.
-A sizer is a tuple $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers indexed by $\mathcal{P}$. Given a sizer $r_{\bullet}$, we define the $r_{\bullet}$-shard as:
-$L[r_{\bullet}]=L\cap\prod_{P\in\mathcal{P}}[-r_{P},r_{P}].$
-For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$. We define $\mathcal{L}_{\text{sh}}$, as the union of all type-shards.
+A sizer is a tuple $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$.
+Given a sizer $r_{\bullet}$, let $\mathbb{R}^{[r_{\bullet}]}=\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of a states space $L$ be
+$L[r_{\bullet}]=L\cap\mathbb{R}^{[r_{\bullet}]}.$
+For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$ (a closed, hence compact subset of $\mathbb{R}^{[r_{\bullet}]}$).
+Let also $\mathcal{L}_{\text{sh}}$ be the union of all type-shards (as the sizer $r_{\bullet}$ varies).
+In general, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$ may be a proper subset. Equality holds in the important special case when $\mathcal{P}$ is countable (see *** below).
 2.1. Compositional Computation Structures
@@
   • $\Gamma\subseteq L^{L}$ is a semigroup under composition.
-The elements of the semigroup $\Gamma$ are called the computations of the structure $(L,\mathcal{P},\Gamma)$.
-If $\Delta\subseteq\Gamma$, we say that $\Delta\subseteq\Gamma$ is $R$-confined if $\gamma|_{L[r_{\bullet}]}:L[r_{\bullet}]\to L[r_{\bullet}]$ for every $r_{\bullet}\in R$ and $\gamma\in\Delta$.
-Elements in $\overline{\Gamma}\subseteq\mathcal{L}_{\text{sh}}$ are called (real-valued) deep computations or ultracomputations.
-For a CCS $(L,\mathcal{P},\Gamma)$, we regard the elements of $\Gamma$ as “standard” finitary computations and the elements of $\overline{\Gamma}$, i.e., deep computations, as possibly infinitary limits of standard computations.
+Elements of the semigroup $\Gamma$ are called the computations of the structure $(L,\mathcal{P},\Gamma)$.
+We assume that the identity map id on $L$ is an element of $\Gamma$ (which is thus not merely a semigroup but a monoid of transformations of $L$).
+We topologize $\Gamma$ as a subset of the topological product $L^{L}$, where the “exponent” $L$ serves merely as an index set, but the “base” $L$ is topologized by type; consequently, one may identify $\Gamma$ with a subset of the topological product $(\mathbb{R}^{\mathcal{P}})^{L}$.
+More specifically, $\Gamma$ is identified with a subset of $\mathcal{L}^{L}$, which is a closed subspace of $(\mathbb{R}^{\mathcal{P}})^{L}$.
+Therefore, we have an inclusion $\overline{\Gamma}\subseteq\mathcal{L}^{L}$.
+Elements $\xi\in\overline{\Gamma}$ are called (real-valued) deep computations or ultracomputations.
+A collection $R$ of sizers is exhaustive if $L=\bigcup_{r_{\bullet}\in R}L[r_{\bullet}]$ (shards $L[r_{\bullet}]$ exhaust $L$).
+A transformation $\gamma\in\Gamma$ is $R$-confined if $\gamma$ restricts to a map $\gamma|_{L[r_{\bullet}]}:L[r_{\bullet}]\to L[r_{\bullet}]$ (into $L[r_{\bullet}]$ itself) for every $r_{\bullet}\in R$.
+A subset $\Delta\subseteq\Gamma$ is $R$-confined if each $\gamma\in\Delta$ is.
+Proposition 2.4.
+If $\Delta\subseteq\Gamma$ is confined by an exhaustive sizer collection, then $\overline{\Delta}$ is a compact subset of $\mathcal{L}_{\mathrm{sh}}^{L}$.
+Proof.
+Assume that $R$ confines $\Delta$.
+For each $v\in L$, let $r_{\bullet}^{(v)}\in R$ be a sizer such that $v\in L[r_{\bullet}^{(v)}]$.
+An arbitrary $\gamma\in\Delta$ restricts to a map $\gamma\restriction L[r_{\bullet}^{(v)}]:L[r_{\bullet}^{(v)}]\to L[r_{\bullet}^{(v)}]$, so $\Gamma\subseteq K\coloneqq\prod_{v\in L}\mathcal{L}[r_{\bullet}^{(v)}]$.
+The space $K$ is a product of compact spaces, hence compact, so $\overline{\Gamma}$ is a closed, hence compact subset thereof, and a subset of $\mathcal{L}_{\mathrm{sh}}^{L}\supseteq K$ a fortiori.
+∎
+For a CCS $(L,\mathcal{P},\Gamma)$, we regard the elements of $\Gamma$ as “standard” finitary computations, and the elements of $\overline{\Gamma}$, i.e., deep computations, as possibly infinitary limits of standard computations.
 The main goal of this paper is to study the computability, definability and computational complexity of deep computations.
@@
 2.2. Computability and definability of deep computations and the Extendibility Axiom
-Let $f:L\to\mathcal{L}$ be a function that maps each input state type $(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ into an output state type $(P\circ f(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$.
+Let $f:L\to\mathcal{L}$ be a function that maps each input state type $(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ to an output state type $(P\circ f(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$.
   1. (1)
-    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{L}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ by a finite number of applications of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
+    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{L}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ and by a finite number of iterations of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
     $|\varphi_{Q,\,K,\varepsilon}(v)-Q{\circ}f(v)|\leq\epsilon,\qquad\text{for all $v\in K$}.$
-    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model the theory of real-valued structures first introduced in [Iovino:1994] for model theory of functional analysis and now standard in model theory (see [Keisler:2023]).
+    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model the theory of real-valued structures first introduced in [IOV94] for model theory of functional analysis and now standard in model theory (see [KEI03]).
 The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
@@
 This is quantifier-free definability (i.e., definability as given by the preceding paragraph, but without use of quantifiers), which, from a logic viewpoint, corresponds to computability (the presence of the quantifiers $\exists$ and $\forall$ are the reason behind the undecidability of first-order logic).
-It is shown in [alva2024approximability] that:
+It is shown in [ADI+24] that:
   1. (1)
@@
 Definition 2.5.
-We say that a CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendability Axiom if for all $\gamma\in\Gamma$, there is $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous.
+We say that a CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom if for all $\gamma\in\Gamma$, there is $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous.
 We refer to $\tilde{\gamma}$ as a free extension of $\gamma$.
-By the preceding remarks, the Extendability Axiom says that the elements of the semigroup $\Gamma$ are finitary computations.
+By the preceding remarks, the Extendibility Axiom says that the elements of the semigroup $\Gamma$ are finitary computations.
 For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$.
 For any $\Delta\subseteq\Gamma$, let $\tilde{\Delta}$ denote $\{\tilde{\gamma}\colon\gamma\in\Delta\}$.
-For a more detailed discussion of the Extendability Axiom, we refer the reader to [alva2024approximability].
-For an illustrative example, we can frame Newton’s polynomial root approximation method in the context of a CCS (see Example 5.6 of [alva2024approximability] for details) as follows.
-Begin by considering the extended complex numbers $\hat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ with the usual Riemann sphere topology that makes it into a compact space (where unbounded sequences converge to $\infty$).
-In fact, not only is this space compact, but it is covered by the shard given by the sizer $(1,1,1)$ (the unit sphere is contained in the cube $1,1]^{3}$).
-The space $\hat{\mathbb{C}}$ is homeomorphic to the usual unit sphere $S^{2}:=\{(x,y,z):x^{2}+y^{2}+z^{2}=1\}$ of $\mathbb{R}^{3}$, by means of the stereographic projection and its inverse $\hat{\mathbb{C}}\to S^{2}$.
-This function is regarded as a triple of predicates $x,y,z:\hat{\mathbb{C}}\to[-1,1]$ where each will map an extended complex number to its corresponding real coordinate on the cube $1,1]^{3}$.
-Now fix the cubic complex polynomial $p(s):=s^{3}-1$, and consider the map which performs one step in Newton’s method at a particular (extended) complex number $s$, for finding a root of $p$, $\gamma_{p}:\hat{\mathbb{C}}\to\hat{\mathbb{C}}$.
-The explicit inner workings of $\gamma_{p}$ are irrelevant for this example, except for the fact that it is a continuous mapping.
-It follows that $(S^{2},\{x,y,z\},\{\gamma_{p}^{k}:k\in\mathbb{N}\})$ is a CCS.
-The idea is that repeated applications of $\gamma_{p}(s),\gamma_{p}\circ\gamma_{p}(s),\gamma_{p}\circ\gamma_{p}\circ\gamma_{p}(s),\dots$ would approximate a root of $p$ provided $s$ was a good enough initial guess.
+For a more detailed discussion of the Extendibility Axiom, we refer the reader to [ADI+24].
+2.3. Newton’s method as a CCS
+Let $p(z)$ be a non-constant polynomial with complex coefficients. Newton’s method is an iterative method that is used to approximate a root of $p(z)$. Define the Newton map as:
+$N_{p}(z)=z-\frac{p(z)}{p^{\prime}(z)}$
+The method consists of taking an initial guess $z_{0}\in\mathbb{C}$ and iterating the rational map $N_{p}$ to obtain a sequence given by
+$z_{n+1}=N_{p}(z_{n})=z_{n}-\frac{p(z_{n})}{p^{\prime}(z_{n})}$
+For each root $r\in\mathbb{C}$ of $p(z)$, there exists $\varepsilon>0$ such that for any initial guess $z_{0}$ in the $\varepsilon$-ball centered at $r$, Newton’s iteration converges to $r$ (provided $p^{\prime}(r)\neq 0$) and the convergence is quadratic in that case, meaning the error at each step is roughly squared, causing the number of correct digits to double, leading to fast convergence.
+Denote by $N_{p}^{n}:=N_{p}\circ N_{p}\circ\cdots\circ N_{p}$ (nth iteration of $N_{p}$).
+Given a root $r$ of $p(z)$, the set $B_{r}=\{z\in\mathbb{C}:\lim_{n\rightarrow\infty}N_{p}^{n}(z)=r\}$ is an open set called the basin of $r$.
+However, Newton’s method can fail to converge to any root for some choices of $z_{0}$.
+For example, consider the polynomial $p(z)=z^{3}-2z+2$.
+The Newton map is given by
+$N_{p}(z)=z-\frac{z^{3}-2z+2}{3z^{2}-2}=\frac{2z^{3}-2}{3z^{2}-2}$
+Notice that taking $z_{0}=0$ as an initial guess will yield the sequence $0,1,0,1,0,1,\dots$ that oscillates between 0 and $1$ but none of them are roots of $p(z)$.
+Another more chaotic way Newton’s method can fail to converge is when the sequence of iterations has no convergent subsequence.
+The set of such points is known as the Julia set associated to $N_{p}$ and it is typically a fractal.
+This can be visualizaed by adding a dash of color: let us give each complex number $z_{0}$ a color $(R,G,B)$ where $R,G,B\in[0,1]$ (so that $(1,0,0)$ is red, $(0,1,0)$ is green, $(0,0,1)$ is blue and $(0.5,0,0.5)$ is a light purple, for example).
+The values of $R$, $G$ and $B$ are determined by looking at the image of said number at each stage of the iteration, $N_{p}^{n}(z_{0})$, and computing the current distance to each of the roots of $p(z)$; so $R=1/d_{r}$ where $d_{r}$ is the positive distance to the root which is colored red, and so on.
+In this way, the roots themselves are colored red, green, and blue, and every other point gets a mix of the three colors.
+As the number of iterations increases, each point gets a sharper color, as the sequence of images $\{N_{p}^{n}(z_{0})\}_{n=1}^{\infty}$ converges to one of the three roots.
+Each stage, the complex plane looks as if out of focus because the coloring function is continuous.
+As the reader can see in Figure 1, the points at the boundary of each color class form the famous Newton fractal (of which, interestingly, Newton was unaware).
+(a) After 1 iteration
+(b) After 3 iterations
+(c) After 100 iterations
+Figure 1. Newton’s method approximating $p(z)=z^{3}-2z+2$. Notice the regions of divergence.
+Another example of a Newton’s fractal is for $p(z)=z^{3}-1$. The roots of $p(z)$ are the 3rd roots of unity and the Newton map is given by:
+$N_{p}(z)=z-\frac{z^{3}-1}{3z^{2}}=\frac{2z^{3}+1}{3z^{2}}$
+In this case, there are three basins of attraction (one for each root) and the complement of their union is the Julia set, i.e., the common boundary.
+(a) After 1 iteration
+(b) After 3 iterations
+(c) After 100 iterations
+Figure 2. Newton’s method approximating $p(z)=z^{3}-1$.
+We can frame Newton’s method as a CCS (satisfying the Extendibility Axiom) as follows.
+Given a non-constant polynomial $p(z)$, let $L\subseteq\mathbb{C}$ be the complement of the Julia set (this is called the Fatou set) associated to $N_{p}$.
+This is the set where Newton’s method is tame.
+By [BLA84, Corollary 4.6], $L$ is open and dense in $\mathbb{C}$.
+Let $\mathcal{P}=\{P_{1},P_{2},P_{3}\}$ where the map $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$ is the stereographic projection into the Riemann Sphere $S^{2}$, i.e.,
+$P_{1}(z)=\frac{2{\rm Re}(z)}{|z|^{2}+1},$
+$P_{2}(z)=\frac{2{\rm Im}(z)}{|z|^{2}+1},$
+$P_{3}(z)=\frac{|z|^{2}-1}{|z|^{2}+1}.$
+Since $L$ is dense in the extended complex plane, then $\mathcal{L}=S^{2}$.
+Computations are the iterations of Newton’s method, i.e., $\Delta=\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$.
+Then, $(L,\mathcal{P},\Gamma)$ is a CCS.
+Since all computations $N_{p}^{n}$ are rational maps, they can be continuously extended to the extended complex plane $\overline{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, i.e., to $\mathcal{L}$.
+In particular, $(L,\mathcal{P},\Gamma)$ satisfy the Extendibility Axiom.
+The set of deep computations $\overline{\Delta}$ might behave different for various polynomials.
+Let us look at various examples:
+Example 2.6.
+[…]
+$P_{1}(x)=\frac{2x}{x^{2}+1},$
+$P_{2}(x)=\frac{x^{2}-1}{x^{2}+1}.$
+The Newton’s method map $N_{p}:L\rightarrow L$ is given by
+$N_{p}(x)=\frac{x^{2}+a}{2x}.$
+Note that $\mathcal{L}=S^{1}$ and that each iterate $N_{p}^{n}$ can be continuously extended to the extended real line $\mathbb{R}\cup\{\infty\}$, i.e., $\mathcal{L}$. For example,
+$\tilde{N_{p}}(x)=\begin{cases}\frac{x^{2}+a}{2x},&\text{if }x\in L;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
+For every initial guess $x\in L$, the limit $f(x)=\lim_{n\rightarrow\infty}N_{p}^{n}(x)$ converges pointwise to one of the roots. Moreover,
+$f(x)=\begin{cases}\sqrt{a},&\text{if }x>0;\\ -\sqrt{a},&\text{if }x<0.\end{cases}$
+Notice that $f$ can be extended to $\mathcal{L}$ by
+$\tilde{f}(x)=\begin{cases}\sqrt{a},&\text{if }x>0;\\ -\sqrt{a},&\text{if }x<0;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
+However, $\tilde{f}$$\mathcal{L}\rightarrow\mathcal{L}$ is not continuous.
+The set $\overline{\tilde{\Delta}}$ of deep computations is $\tilde{\Delta}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
+Example 2.7.
+Newton’s method for $p(z)=z^{3}-2z+2$.
+Let $r_{1}$, $r_{2}$ and $r_{3}$ be the three roots of $p(z)$.
+Let $B_{1}$, $B_{2}$ and $B_{3}$ be their respective basins.
+Let $B$ be the basin of the attractive cycle $0,1,0,1,\dots$.
+Let $L=B\cup\bigcup_{i=1}^{3}B_{i}$ and $J$ be its complement, i.e., the Julia set.
+Notice that $N_{p}^{n}$ does not converge pointwise.
+However, the subsequences $N_{p}^{2n}$ and $N_{p}^{2n+1}$ are pointwise convergent to functions $f_{1}$ and $f_{2}$ respectively.
+$f_{1}$ and $f_{2}$ are two distinct deep computations.
+Note that for $z\in J$, no subsequence of $\tilde{N_{p}^{n}}(z)$ converges to a complex number.
+However, since $\mathcal{L}=S^{2}$ is compact there is a subsequence of $\tilde{N_{p}^{n}}(z)$ that converges to $\infty$.
+We can extend $f_{i}:L\rightarrow\mathcal{L}$ to $\tilde{f_{i}}$$\mathcal{L}\rightarrow\mathcal{L}$ by:
+$\tilde{f_{i}}(z)=\begin{cases}f_{i}(z),&\text{if }z\in L;\\ \infty,&\text{if }z\in J.\end{cases}.$
+Again, note that $\tilde{f_{i}}$ for $i=1,2$ are not continuous and that $\tilde{f_{i}}\in\overline{\tilde{\Delta}}$.
+2.4. Finite precision threshold classifiers as a CCS
+Let $L=2^{\mathbb{N}}$, i.e., the set consisting of all infinite binary sequences with the topology of pointwise convergence.
+Let $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ is the collection of projections, i.e., $P_{n}(x)=x(n)$ for each $n$ and each $x\in L$.
+In other words, $P_{n}$ is the predicate that reads the $n$th digit of a sequence.
+Notice that $L\subseteq\mathbb{R}^{\mathcal{P}}$ is closed.
+Therefore, $\mathcal{L}=L$.
+We denote by $0^{\infty}$ the infinite binary sequence consisting of 0s, and by $1^{\infty}$ the infinite binary sequence consisting of $1$s.
+The set of finite binary strings is denoted by $2^{<\mathbb{N}}$.
+Given a finite binary string $w$, we consider the transition $\phi_{w}:L\to L$ given by the rule
+$\phi_{w}(x)=\begin{cases}1^{\infty},&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0^{\infty},&\text{otherwise};\end{cases}$
+where $|w|$ is the length of the string $w$ and $x|_{|w|}$ is the prefix of $x$ of length $|w|$.
+That is, $\phi_{w}(x)$ is equal to the constant sequence of ones if $x|_{|w|}$ comes before or is equal to $w$ in the lexicographic order of strings, and it is equal to the constant sequence of zeros otherwise.
+In words, $\phi_{w}$ checks if a number is less than or equal to the scalar value of threshold $w$ (the string $w$ is finite, hence the classifier has finite precision).
+Note that $P_{n}\circ\phi_{w}(x)=1$ if and only if $x|_{|w|}$ comes before $w$.
+Proposition 2.8.
+$\phi_{w}:2^{\mathbb{N}}\rightarrow 2^{\mathbb{N}}$ is continuous for all $w\in 2^{<\mathbb{N}}$.
+Proof.
+It suffices to prove that $P_{n}\circ\phi_{w}:2^{\mathbb{N}}\rightarrow\{0,1\}$ is continuous for all $n\in\mathbb{N}$. For simplicity, let us call $f:=P_{n}\circ\phi_{w}$, i.e.,
+$f(x)=\begin{cases}1,&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0,&\text{otherwise}.\end{cases}$
+We first prove that $f^{-1}(1)=\{x\in 2^{\mathbb{N}}:x|_{|w|}\leq_{\mathrm{lex}}w\}$ is an open set.
+Fix $x_{0}\in f^{-1}(1)$.
+Let $t:=x_{0}|_{|w|}$ and consider the basic open set $t]=\{x\in 2^{\mathbb{N}}:x|_{|t|}=t\}$.
+Then, it is not difficult to check that $x_{0}\in[t]\subseteq f^{-1}(1)$.
+The same reasoning shows that $f^{-1}(0)$ is open.
+∎
+Let $\Delta=\{\phi_{w}:w\in 2^{<\mathbb{N}}\}\cup\{\mathbf{0^{\infty}},\mathbf{1^{\infty}}\}$ where $\mathbf{0^{\infty}},\mathbf{1^{\infty}}:L\rightarrow L$ are the constant maps identical to $0^{\infty}$ and $1^{\infty}$, respectively.
+Let $\Gamma$ be the semigroup generated by $\Delta$.
+The preceding proposition shows that $\Delta$ (and hence $\Gamma$) consists of continuous functions.
+In particular, the CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom.
+In contrast with Newton’s method, the algebraic structure of $\Delta$ is quite simple: composing two classifiers results in something similar to a Boolean logic gate.
+The topological structure is far more interesting.
+Intuitively, the crucial difference between Newton’s method and threshold classifiers is that the complexity of the former comes from depth: the semigroup is generated by a single map but its iterates are highly complex.
+The complexity of threshold classification comes from width: the semigroup has infinitely many generators, but their compositions are simple.
+Intuitively, the closure of $\Delta$ consists of the set of all possible threshold classifiers on the real line, but there are two sorts: the ones that classify strict inequalities and those that classify $\leq$.
+The members of $\Delta$ are finite-precision approximations of classifiers that check all bits of information.
+But here it gets interesting: what is the difference, in terms of floating-point arithmetic, between $x<0.5$ and $x\leq 0.5$?
+Suppose that $f_{a}^{+}$ represents the $\leq$ classifier for a target $a\in L$.
+We identify the scalar truth values with constant sequences, formally $f_{a}^{+}:L\to\{0^{\infty},1^{\infty}\}$ is defined by $f_{a}^{+}(x)=1^{\infty}$ if $x\leq_{\mathrm{lex}}a$ and $f_{a}^{+}(x)=0^{\infty}$ otherwise.
+Note that if $a$ is the constant $1^{\infty}$, then $f_{a}^{+}=\mathbf{1^{\infty}}$.
+Similarly, let $f_{a}^{-}$ be the strict inequality << classifier, i.e., $f_{a}^{-}(x)=1^{\infty}$ if $x<_{\mathrm{lex}}a$ and $f_{a}^{-}(x)=0^{\infty}$ otherwise.
+Note that if $a$ is the constant zero, then $f_{a}^{-}=\mathbf{0^{\infty}}$.
+Proposition 2.9.
+$f_{a}^{+},f_{a}^{-}\in\overline{\Delta}$ for all $a\in 2^{\mathbb{N}}$.
+Proof.
+First, we show that $f_{a}^{+}\in\overline{\Delta}$.
+If $a=1^{\infty}$, then $f_{a}^{+}=\mathbf{1^{\infty}}\in\Delta$.
+If $a$ is not identically $1$, we argue that the pointwise limit of the threshold classifiers on $w_{n}:=a|_{n}^{\frown}1$ (that is, the sequence obtained from appending a $1$ to the first $n$ bits of $a$) is precisely $f_{a}^{+}$.
+Specifically, for every $x\in L$, we intend to prove that $\lim_{n\to\infty}\phi_{w_{n}}(x)=f_{a}^{+}(x)$.
+Assume that $x>_{\mathrm{lex}}a$.
+Let $m$ be the least index at which the two sequences differ.
+Then $a(m)=0<1=x(m)$, and for all $n\geq m$, $w_{n}$ agrees with $a$ up to $m$.
+Crucially, $w_{n}(m)=0<1=x(m)$, which implies that $w_{n}<_{\mathrm{lex}}x|_{n+1}$, and hence $\phi_{w_{n}}(x)=0^{\infty}=f_{a}^{+}(x)$ for large enough $n$.
+If $x\leq_{\rm lex}a$, then $x|_{n+1}\leq_{\rm lex}w_{n}$ for all $n\in\mathbb{N}$.
+Hence, $\phi_{w_{n}}(x)=1^{\infty}=f_{a}^{+}(x)$ for all $n\in\mathbb{N}$.
+Now, we prove that $f_{a}^{-}\in\overline{\Delta}$. If $a$ is the constant zero, then $f_{a}^{-}=\mathbf{0^{\infty}}\in\Delta$. Suppose that $a$ is not constantly zero; then we have two cases.
+  1. (1)
+    If $a$ is eventually zero ($a$ is often called a dyadic rational), that is $a=u^{\frown}1^{\frown}0^{\infty}$ (here denotes concatenation) for some finite $u$.
+Let $w_{n}:=u^{\frown}0^{\frown}1^{n}<_{\mathrm{lex}}a$.
+We claim that $\lim_{n\to\infty}\phi_{w_{n}}(x)=f_{a}^{-}(x)$.
+Assume that $x<_{\rm lex}a$.
+Then, $x|_{|w_{n}|}\leq_{\rm lex}w_{n}$ for large enough $n$.
+Hence, $\phi_{w_{n}}(x)=1^{\infty}=f_{a}^{-}(x)$ for large enough $n$.
+Now assume that $x\geq_{\rm lex}a$.
+Then, $w_{n}<_{\rm lex}a|_{|w_{n}|}\leq_{\rm lex}x|_{|w_{n}|}$ so $\phi_{w_{n}}(x)=0^{\infty}=f_{a}^{-}(x)$ for all $n\in\mathbb{N}$.
+  2. (2)
+    If $a$ is not eventually zero, enumerate the indices of all positive bits in $a$, $\{n\in\mathbb{N}:a(n)=1\}$, strictly increasingly as $\{n_{k}:k\in\mathbb{N}\}$ (this is possible as the former set is infinite by assumption).
+Let $w_{k}:=(a|_{n_{k}-1})^{\frown}0$; that is, $w_{k}$ is the result of flipping the $k$-th positive bit in $a$.
+Once again, observe that $w_{k}<_{\mathrm{lex}}a$ for all $k$.
+The crucial step follows from the fact that for any $x<_{\mathrm{lex}}a$, there is a large enough $K$ such that $x<_{\mathrm{lex}}w_{k}$ for all $k\geq K$.
+The preceding proposition shows that the topological structure of deep computations can be very complicated.
+Indeed, $\overline{P_{n}\circ\Delta}$ contains the Split Cantor space for all $n\in\mathbb{N}$.
+(see Examples 3.3(3)).
+2.5. Finite precision prefix test
+In this subsection we present another example of a CCS with a more complicated set of deep computations.
+Let $L=2^{\mathbb{N}}$ and $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ where $P_{n}(x)=x(n)$ are the projection maps so clearly $L\subseteq\mathbb{R}^{\mathcal{P}}$ and $\mathcal{L}=L$ (same computation states structure as subsection 2.4).
+For each $w\in 2^{<\mathbb{N}}$, let $\psi_{w}:L\rightarrow L$ be the transition given by:
+$\psi_{w}(x)=\begin{cases}1^{\infty},&\text{if }x|_{|w|}=w;\\ 0^{\infty},&\text{otherwise}.\end{cases}$
+In other words, $\psi_{w}$ determines whether the first $|w|$ bits of a binary sequence is exactly $w$.
+Let $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$ and $\Gamma$ be the semigroup generated by $\Delta$.
+Since the sets $\{x\in 2^{\mathbb{N}}:x|_{|w|}=w\}$ are open and closed in $2^{\mathbb{N}}$, then the transitions $\psi_{w}$ are all continuous.
+In particular, $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom.
+Let us analyze the set of deep computations of $\Delta$.
+The idea of these finite precision prefix tests $\psi_{w}$ is that they are approximating the equality relation on infinite binary sequences.
+For a given $a\in 2^{\mathbb{N}}$, let $\delta_{a}:L\rightarrow\{0^{\infty},1^{\infty}\}$ be the indicator function at $a$, i.e., $\delta_{a}(x)=1^{\infty}$ if $x=a$ and $\delta_{a}(x)=0^{\infty}$ otherwise.
+Proposition 2.10.
+$\delta_{a}\in\overline{\Delta}$ for all $a\in 2^{\mathbb{N}}$.
+Proof.
+Fix $a\in 2^{\mathbb{N}}$.
+Let $w_{n}:=a|_{n}$ for each $n\in\mathbb{N}$.
+We claim that $\lim_{n\rightarrow\infty}\psi_{w_{n}}(x)=\delta_{a}(x)$ for all $x\in L$.
+If $x=a$, then $x|_{|w_{n}|}=w_{n}$ for all $n$ and so $\psi_{w_{n}}(x)=1^{\infty}=\delta_{a}(x)$ for all $n$.
+If $x\neq a$, then $x|_{|w_{n}|}\neq w_{n}$ for large enough $n$.
+Hence, $\psi_{w_{n}}(x)=0^{\infty}=\delta_{a}(x)$ for large enough $n$.
+∎
+These equality tests $\delta_{a}$ are not all the deep computations. The other deep computation we are missing is the constant map $\mathbf{0^{\infty}}$.
+Proposition 2.11.
+$\mathbf{0^{\infty}}\in\overline{\Delta}$.
+Proof.
+To show that $\mathbf{0^{\infty}}\in\overline{\Delta}$ consider, for each $n\in\mathbb{N}$, ${w_{n}=1^{n}}^{\frown}0$, i.e., the string consisting of $n$ consecutive $1$s followed by a 0.
+If $x=1^{\infty}$, then $x|_{|w_{n}|}\neq w_{n}$ for all $n\in\mathbb{N}$.
+Hence, $\psi_{w_{n}}(x)=0^{\infty}$ for all $n\in\mathbb{N}$.
+If $x\neq 1^{\infty}$, let $N$ be the smallest such that $x(N)=0$.
+Then, $x|_{|w_{n}|}\neq w_{n}$ for all $n>N$.
+Hence, $\psi_{w_{n}}(x)=0^{\infty}$ for large enough $n$.
+∎
+In fact, $\overline{\Delta}=\Delta\cup\{\delta_{a}:a\in 2^{\mathbb{N}}\}\cup\{\mathbf{0^{\infty}}\}$ and this space is known as the Extended Alexandroff compactification of $2^{\mathbb{N}}$ (see Example 3.3(2)).
+One key topological property about this space is that $\mathbf{0^{\infty}}$ is not a $G_{\delta}$ point, i.e., $\{\mathbf{0^{\infty}}\}$ is not a countable intersection of open sets.
+Moreover, $\mathbf{0^{\infty}}$ is the only non-$G_{\delta}$ point.
+It is well-known that in a Hausdorff, first countable space every point is $G_{\delta}$.
+This shows that our space of deep computations is not first countable.
+This space also contains a discrete subspace of size continuum, namely $\{\delta_{a}:a\in 2^{\mathbb{N}}\}$.
 3. Classifying deep computations
@@
 Theorem 3.1.
-Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.3) satisfying the Extendability Axiom (Definition 2.4) with $\mathcal{P}$ countable.
+Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.3) satisfying the Extendibility Axiom (Definition 2.5) with $\mathcal{P}$ countable.
 Let $R$ be an exhaustive collection of sizers.
@@
 Since $\mathcal{P}$ is countable, $\mathcal{L}[r_{\bullet}]\subseteq\mathbb{R}^{\mathcal{P}}$ is Polish.
-Also, the Extendability Axiom implies that $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a pointwise bounded set of continuous functions for all $P\in\mathcal{P}$.
+Also, the Extendibility Axiom implies that $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a pointwise bounded set of continuous functions for all $P\in\mathcal{P}$.
 Hence, Theorem 1.11 and Lemma 1.12 prove the equivalence of (i) and (ii).
@@
 Hence, for all $r_{\bullet}\in R$ we have $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$.
-That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where topological closures coincide with sequential closures, see Theorem 3F in [BFT_1978_PCompactBaire] or Theorem 4.1 in [debs2013rosenthal]).
+That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where topological closures coincide with sequential closures, see Theorem 3F in [BFT78] or Theorem 4.1 in [DEB13]).
 ∎
 3.2. The Todorčević trichotomy and levels of PAC learnability
+In this subsection we study the case when the set of deep computations is a separable Rosenthal compactum. We are interested in the separable case for two reasons:
+  1. (1)
+    In practice, the set $\Delta$ of computations is countable. This implies that the set $\overline{\Delta}$ of deep computations is separable.
+  2. (2)
+    The non-separable case lacks some tools and nice examples, which makes their study more complicated.
+In the separable case we have two important results, which are introduced in this subsection (Todorčević’s Trichotomy) and the next subsection (Argyros-Dodos-Kanellopoulos heptachotomy).
+By introducing Todorčević’s Trichotomy into this framework, we obtain a classification of the complexity of deep computations.
 Given a countable set $\Delta$ of computations satisfying the NIP on features and shards (condition (ii) of Theorem 3.1) we have that $\overline{\tilde{\Delta}_{\mathcal{L}[r_{\bullet}]}}$ (for a fixed sizer $r_{\bullet}$) is a separable Rosenthal compactum (see Definition 1.9).
-Todorčević proved a remarkable trichotomy for Rosenthal compacta [Todorcevic_1999_CompactSubsetsBaire], and later Argyros, Dodos, Kanellopoulos [argyros2008rosenthal] proved an heptachotomy that refined Todorčević’s classification.
-In this section, inspired by the work of Glasner and Megrelishvili [glasner2022tame], we study ways in which this classification allows us to obtain different levels of PAC-learnability and NIP.
+Todorčević proved a remarkable trichotomy for Rosenthal compacta [TOD99], and later Argyros, Dodos, Kanellopoulos [ADK08] proved an heptachotomy that refined Todorčević’s classification.
+In this section, inspired by the work of Glasner and Megrelishvili [GM22], we study ways in which this classification allows us to obtain different levels of PAC-learnability and NIP.
 Recall that a topological space $X$ is hereditarily separable if every subspace is separable, and that $X$ is first countable if every point in $X$ has a countable local basis.
 Every separable metrizable space is hereditarily separable, and R.
-Pol proved that every hereditarily separable Rosenthal compactum is first countable (see section 10 of [debs2013rosenthal]).
+Pol proved that every hereditarily separable Rosenthal compactum is first countable (see section 10 of [DEB13]).
 This suggests the following definition:
 Definition 3.2.
-Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom and $R$ be an exhaustive collection of sizers.
+Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendibility Axiom and $R$ be an exhaustive collection of sizers.
 Let $\Delta\subseteq\Gamma$ be an $R$-confined countable set of computations satisfying the NIP on shards and features (condition (ii) in Theorem 3.1).
@@
 […]
-We now present some separable and non-separable examples of Rosenthal compacta.
+We now present some separable and non-separable examples of Rosenthal compacta. These show that the previously discussed classes NIPi are not equal.
 Examples 3.3.
@@
 Note that this space is a separable Rosenthal compactum which is not first countable.
+This is the example discussed in Section 2.5.
+It is an example of a CCS that is NIP but not NIP1.
   3. (3)
@@
 Let $f_{a}^{+}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{+}(x)=1$ if $x\leq a$ and $f_{a}^{+}(x)=0$ otherwise.
-The split Cantor is the space $S(2^{\mathbb{N}})=\{f_{a}^{-}:a\in 2^{\mathbb{N}}\}\cup\{f_{a}^{+}:a\in 2^{\mathbb{N}}\}$.
+The split Cantor is the space $S(2^{\mathbb{N}})=\{f_{a}^{-}:a\in 2^{\mathbb{N}}\}\cup\{f_{a}^{+}:a\in 2^{\mathbb{N}}\}$, which was obtained as the closure of the space discussed in Section 2.4, giving an example separating NIP2 from NIP3.
 This is a separable Rosenthal compactum.
@@
 Let $K$ be any compact metric space and consider the Polish space $X=C(K)\sqcup K$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$.
-For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={x(a),x∈C(K)0,x∈K g_a^1(x)={x(a),x∈C(K)δ_a(x),x∈K Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$.
+For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={x(a), if x∈C(K);0, if x∈K; g_a^1(x)={x(a), if x∈C(K);δa(x), if x∈K.
+Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$.
 Notice that $D(K)$ is a first countable Rosenthal compactum.
@@
 For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s.
-Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0,x¡a_t1/2,a_t≤x ≤b_t1,b_t¡x Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$.
+Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0, if x¡at;1/2, if at≤x ≤bt;1, if bt¡x.
+Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$.
 Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable.
-In fact, it contains an uncountable discrete subspace (see Theorem 5 in [Todorcevic_1999_CompactSubsetsBaire]).
-Theorem 3.4 (Todorčević’s Trichotomy, [Todorcevic_1999_CompactSubsetsBaire], Theorem 3 in [argyros2008rosenthal]).
+In fact, it contains an uncountable discrete subspace (see Theorem 5 in [TOD99]).
+Theorem 3.4 (Todorčević’s Trichotomy, [TOD99], Theorem 3 in [ADK08]).
 Let $K$ be a separable Rosenthal Compactum.
@@
 $K$ is separable Rosenthal compactumKK is metrizableKK is not metrizableKK is hereditarily separable (copy of $S(I)$)        $K$ is not hereditarily separableKK is first countable (copy of $D(2^{\mathbb{N}})$ or $\hat{D}(S(I))$)$K$ is not first countable (copy of $A(2^{\mathbb{N}})$)
+Todorčević’s Trichotomy suggests that in order to distinguish the classes NIPi, the examples in 3.3 are essential.
+The following examples show that the levels NIPi ($i=1,2,3$) may be distinguished by the topological complexity of deep computations.
+Examples 3.5.
+  1. (1)
+    Let $(L,\mathcal{P},\Gamma)$ be the computation of square root (example 2.6 with $\Delta=\Gamma$.
+We saw that $\overline{\tilde{\Delta}}=\tilde{\Delta}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
+This corresponds to the Alexandroff compactification of a countable discrete set, which is metrizable.
+Hence, $\Delta$ is NIP3 but it is not stable, in the sense that $\overline{\tilde{\Delta}}\not\subseteq C(\mathcal{L},\mathcal{L})$.
+  2. (2)
+    Let $(L,\mathcal{P},\Gamma)$ be the finite precision threshold classifiers (Section 2.4) with $\Delta=\{\phi_{w}:w\in 2^{<\mathbb{N}}\}\cup\{\mathbf{0^{\infty}},\mathbf{1^{\infty}}\}$.
+We saw that $\overline{\Delta}$ is homeomorphic to the Split Cantor space $S(2^{\mathbb{N}})$ (Example 3.3(3)), which is hereditarily separable but not metrizable.
+Hence, $\Delta$ is NIP2 but not NIP3.
+  3. (3)
+    Let $(L,\mathcal{P},\Gamma)$ be the finite precision prefix test (Section 2.5) with $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$.
+We saw that $\overline{\tilde{\Delta}}$ is homeomorphic to the Extended Alexandroff compactification $\hat{A}(2^{\mathbb{N}})$ (Example 3.3(3)), which is separable but not first countable.
+Hence, $\Delta$ is NIP but not NIP1.
 The definitions provided here for NIPi ($i=1,2,3$) are topological. This raises the following question:
-Question 3.5.
+Question 3.6.
 Is there a non-topological characterization for NIPi, $i=1,2,3$?
 3.3. The Argyros-Dodos-Kanellopoulos heptachotomy, and approximability of deep computation by minimal classes
-In the three separable three cases given in 3.3, namely, ($\hat{A}(2^{\mathbb{N}})$, $S(2^{\mathbb{N}})$ and $\hat{D}(S(2^{\mathbb{N}}))$), the countable dense subsets are indexed by the binary tree $2^{<\mathbb{N}}$.
+In the three separable cases given in 3.3, namely, $\hat{A}(2^{\mathbb{N}})$, $S(2^{\mathbb{N}})$ and $\hat{D}(S(2^{\mathbb{N}}))$, the countable dense subsets are indexed by the binary tree $2^{<\mathbb{N}}$.
 This choice of index is useful for two reasons:
@@
 Mathematically, deep computations are pointwise limits of standard computations.
-Howwever, computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
+However, computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
   2. (2)
-    The Ramsey theory of perfect subsets of the Cantor space $2^{\mathbb{N}}$ can be imported to analyze the behavior of the accumulation points.
-Since $2^{<\mathbb{N}}$ is countable, we can always choose this index for the countable dense subsets.
-This is done in [argyros2008rosenthal].
+    Infinite branches of the binary tree $2^{<\mathbb{N}}$ correspond to the Cantor space $2^{\mathbb{N}}$, the canonical perfect set (in the sense that any Polish space with no isolated points contains a copy of $2^{\mathbb{N}}$).
+The use of infinite dimensional Ramsey theory for trees (pioneered by the work of James D.
+Halpern, Hans Läuchli in [HL66] and also Keith Milliken in [MIL81], and Alain Louveau, Saharon Shelah, Boban Velickovic in [LSV93]) and perfect sets (Fred Galvin and Andreas Blass in [BLA81], Arnold W.
+Miller in [MIL89], and Stevo Todorčević in [TOD99]) allowed S.A.
+Argyros, P.
+Dodos and V.
+Kanellopoulos in [ADK08] to obtain an improved version of Theorem 3.4.
+It is no surprise that Ramsey Theory becomes relevant in the study of Rosenthal compacta as it was a key ingredient in Rosenthal’s $\ell_{1}$ Theorem.
+For this reason, the main results in [ADK08] (which we cite in this paper) are better explained by indexing Rosenthal compacta with the binary tree.
 Definition 3.7.
@@
     If $\{f_{t}:t\in 2^{<\mathbb{N}}\}$ is a pointwise bounded family, we say that $\{f_{t}:t\in 2^{<\mathbb{N}}\}$ is minimal if and only if for every dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$, $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to $\{f_{t}:t\in 2^{<\mathbb{N}}\}$.
-One of the main results in [argyros2008rosenthal] is that, up to equivalence, there are seven minimal families of Rosenthal compacta and that for every relatively compact $\{f_{t}:t\in 2^{<\mathbb{N}}\}\subseteq B_{1}(X)$ there is a dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ such that $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to one of the minimal families.
+One of the main results in [ADK08] is that, up to equivalence, there are seven minimal families of Rosenthal compacta and that for every relatively compact $\{f_{t}:t\in 2^{<\mathbb{N}}\}\subseteq B_{1}(X)$ there is a dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ such that $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to one of the minimal families.
 We shall describe the seven minimal families next.
-We follow the same notation as in [argyros2008rosenthal].
+We follow the same notation as in [ADK08].
 For any node $t\in 2^{<\mathbb{N}}$, let us denote by $t^{\frown}0^{\infty}$ ($t^{\frown}1^{\infty}$) the infinite binary sequence starting with $t$ and continuing will all 0’s (respectively, all $1$’s).
@@
     $D_{7}=\{(v_{s_{t}},x^{+}_{s_{t}^{\frown}0^{\infty}}):t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{7}}=\hat{D}(S(2^{\mathbb{N}}))$
-Theorem 3.7 (Heptachotomy of minimal families, Theorem 2 in [argyros2008rosenthal]).
+Theorem 3.8 (Heptachotomy of minimal families, Theorem 2 in [ADK08]).
 Let $X$ be Polish.
@@
 Moreover, all $D_{i}$ are minimal and mutually non-equivalent.
-4. Measure-theoretic versions of NIP and universal Monte Carlo computability of deep computations
-The countability assumption on $\mathcal{P}$ played a crucial role in the proof of Theorem 1.12 , as it makes $\mathbb{R}^{\mathcal{P}}$ a Polish space.
-In this section, we do not assume that $\mathcal{P}$ is countable.
-We replace deterministic computability by measure-theoretic (‘Monte Carlo’) computability.
-4.1. A measure-theoretic version of NIP
-Recall that the raison d’être of the class of Baire-1 functions is to have a class that contains the continuous functions but is closed under pointwise limits, and that for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$ (see Fact 1.2).
+The implications of this result for deep computations is the following: for any countable set of computations $\Delta$ satisfying the NIP (for some CCS $(L,\mathcal{P},\Gamma)$), we can always find a countable discrete set of deep computations that approximate all the other deep computations.
+For example: in the finite precision prefix test example (subsection 2.5), the prefix test computations (family $D_{5}$) approximate all other deep computations.
+However, note that this discrete set $D_{i}$ may not consist of continuous functions (i.e., they will not be computable in general).
+For example, functions in $D_{3}$ are not continuous.
+4. Randomized versions of NIP and Monte Carlo computability of deep computations
+In this section, we replace deterministic computability by probabilistic (‘Monte Carlo’) computability.
+We do not assume that $\mathcal{P}$ is countable.
+The main results of the section are Theorem 4.8 (connecting NIP and Monte Carlo computability) and 4.14 (connecting Talagrand stability and Monte Carlo computability).
+Fundamental in this section is a measure-theoretic version of Theorem 1.11, namely, Theorem 4.7.
+For the proof of Theorem 1.11, we assumed countability of $\mathcal{P}$ — this ensured that $\mathbb{R}^{\mathcal{P}}$ a Polish space.
+In this section, the countability assumption is not needed.
+4.1. NIP and Monte Carlo computability of deep computations
+The raison d’être of the Baire class-1 functions is have a class of functions that contains the continuous functions and is closed under pointwise limits.
+By Fact 1.2, for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$.
 This motivates the following definition:
 Definition 4.1.
-Given a Hausdorff space $X$ and a measurable space $(Y,\Sigma)$, we say that $f:X\rightarrow Y$ is universally measurable (with respect to $\Sigma$) if $f^{-1}(E)$ is Borel for every $E\in\Sigma$, i.e., $f^{-1}(E)$ is $\mu$-measurable for every Radon measure $\mu$ on $X$.
-When $Y=\mathbb{R}$ we will always take $\Sigma=\mathcal{B}(\mathbb{R})$, the Borel $\sigma$-algebra of $\mathbb{R}$.
+Given a Hausdorff space $X$ and a measurable space $(Y,\Sigma)$, we say that $f:X\rightarrow Y$ is universally measurable (with respect to $\Sigma$) if $f^{-1}(E)$ is $\mu$-measurable for every Radon measure $\mu$ on $X$ and every $E\in\Sigma$.
+When $Y=\mathbb{R}$, we will always take $\Sigma=\mathcal{B}(\mathbb{R})$, the Borel $\sigma$-algebra of $\mathbb{R}$.
 Remark 4.2.
@@
 Intuitively, a function is universally measurable if it is “measurable no matter which reasonable way you try to measure things on its domain”.
-The concept of universal measurability emerged from work of Kallianpur and Sazonov, in the late 1950’s and 1960s, with later developments by Blackwell, Darst, and others, building on earlier ideas of Gnedenko and Kolmogorov from the 1950s.
-See [Pap:2002, Chapters 1 and 2].
+The concept of universal measurability emerged from work of Kallianpur and Sazonov, in the late 1950’s and 1960s — with later developments by Blackwell, Darst and others — building on earlier ideas of Gnedenko and Kolmogorov from the 1950s.
+See [PAP02, Chapters 1 and 2].
 Notation 4.3.
-Following [BFT_1978_PCompactBaire], the collection of all universally measurable real-valued functions will be denoted by $M_{r}(X)$.
-In the context of deep computations, we will be interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself.
-There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra, i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$.
+Following [BFT78], the collection of all universally measurable real-valued functions on $X$ will be denoted by $M_{r}(X)$.
+Given a fixed Radon measure $\mu$ on $X$, the collection of all $\mu$-measurable real-valued functions on $X$ will be denoted by $\mathscr{M}^{0}(X,\mu)$.
+In the context of deep computations, we are interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself.
+There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the set $\pi_{i}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open).
 Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller.
@@
 The reason for this choice is the following characterization:
-Lemma 4.4.
+Proposition 4.4.
 Let $X$ be a Hausdorff space and $Y=\prod_{i\in I}Y_{i}$ be any product of measurable spaces $(Y_{i},\Sigma_{i})$ for $i\in I$.
 Let $\Sigma_{Y}$ be the cylinder $\sigma$-algebra generated by the measurable spaces $(Y_{i},\Sigma_{i})$.
-Let $f:X\rightarrow Y$.
-The following are equivalent:
+The following are equivalent for $f:X\rightarrow Y$.:
   1. (i)
@@
 […]
-The preceding lemma says that a transition map is universally measurable if and only if it is universally measurable on all its features; in other words, we can check measurability of a transition just by checking measurability feature by feature.
-We will denote by $M_{r}(X,\mathbb{R}^{\mathcal{P}})$ the collection of all universally measurable functions $f:X\rightarrow\mathbb{R}^{\mathcal{P}}$ (with respect to the cylinder $\sigma$-algebra), endowed with the topology of pointwise convergence.
-We will need the following result about NIP and universally measurable functions:
-Theorem 4.5 (Bourgain-Fremlin-Talagrand, Theorem 2F in [BFT_1978_PCompactBaire]).
+The preceding proposition says that a transition map is universally measurable if and only if it is universally measurable on all its features; in other words, we can check measurability of a transition just by checking measurability feature by feature.
+This is the same as in the Baire class-1 case (compare with Proposition 1.10).
+The main result in section 3 is that, as long as we work with countably many features, PAC-learning (or NIP) corresponds to relative compactness in the space of Baire class-1 functions.
+The following result (which does not assume countability of the number of features) gives an analogous characterization of the NIP in terms of universal measurability:
+Theorem 4.5 (Bourgain-Fremlin-Talagrand, Theorem 2F in [BFT78]).
 Let $X$ be a Hausdorff space and $A\subseteq C(X)$ be pointwise bounded. The following are equivalent:
@@
     For every Radon measure $\mu$ on $X$, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$, i.e., every countable subset of $A$ has an accumulation point in $\mathscr{M}^{0}(X,\mu)$.
-4.2. Universal Monte Carlo computability of deep computations
-We now wish to define the concept of a deep computation being computable except on a set of arbitrarily small measure “no matter which reasonable way you try to measure things on its domain” (see the remarks following definition 4.1).
-This is the concept of universal Monte Carlo computability defined below (Definition 4.6).
-To motivate the definition, we need to recall two facts:
+This result allows us to formalize the concept of a deep computation being Monte Carlo computable, which we define below (Definition 4.6). To motivate the definition, let us first recall two facts:
   1. (1)
     Littlewoood’s second principle states that every Lebesgue measurable function is “nearly continuous”.
-The formal version of this, which is Luzin’s theorem, states that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ be a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable ) equipped with a Borel algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|F$ is continuous.
+The formal statement of this, which is Luzin’s theorem, is that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ is a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable ) equipped with a Borel algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|_{F}$ is continuous and $\mu(E\backslash F)<\varepsilon$.
   2. (2)
-    Computability of deep computations can is characterized in terms of continuous extendibility of computations. This is at the core of [alva2024approximability].
+    Computability of deep computations is characterized in terms of continuous extendibility of computations. This is at the core of [ADI+24].
 These two facts motivate the following definition:
@@
 Let $(L,\mathcal{P},\Gamma)$ be a CCS.
-We say that a transition $f:L\rightarrow L$ is universally Monte Carlo computable if and only if there exists $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ extending $f$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that the restriction $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow\mathcal{L}[s_{\bullet}]$ is universally measurable, i.e., $\pi_{P}\circ\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow[-s_{P},s_{P}]$ is $\mu$-measurable for every Radon probability measure $\mu$ on $\mathcal{L}[r_{\bullet}]$.
-4.3. Bourgain-Fremlin-Talagrand, NIP, and universal Monte Carlo computability of deep computations
-Theorem 4.5 immediately yields the following.
+We say that a transition $f:L\rightarrow L$ is universally Monte Carlo computable if and only if there exists $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ extending $f$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that the restriction $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow\mathcal{L}[s_{\bullet}]$ is universally measurable, i.e., $\pi_{P}\circ\tilde{f}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\rightarrow[-s_{P},s_{P}]$ is $\mu$-measurable for every Radon probability measure $\mu$ on $\mathcal{L}[r_{\bullet}]$ and $P\in\mathcal{P}$.
+Remark 4.7.
+Condition (2) of Theorem  shows that to study measure-theoretic versions of NIP, we need only consider compact subsets of $X$.
+Now, every Radon measure on a compact space is finite; hence, by proper normalization, it can be treated as a probability measure.
+Therefore, in the context of Monte Carlo measurability, we focus on Radon probability measures rather than general Radon measures.
 Theorem 4.8.
-Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom.
+Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendibility Axiom.
 Let $R$ be an exhaustive collection of sizers.
 Let $\Delta\subseteq\Gamma$ be $R$-confined.
-If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and all $r_{\bullet}\in R$, then every deep computation is universally Monte Carlo computable.
+If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and all $r_{\bullet}\in R$, then every deep computation in $\Delta$ is universally Monte Carlo computable.
 Proof.
-By the Extendability Axiom, Theorem 4.5 and Lemma 1.13 we have $\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$.
+Fix $P\in\mathcal{P}$ and $r_{\bullet}\in R$.
+By the Extendibility Axiom, $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a set of pointwise bounded continuous functions on the compact set $\mathcal{L}[r_{\bullet}]$.
+Since $\pi_{P}\circ\tilde{\Delta}|_{L[r_{\bullet}]}=\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ has the NIP, so does $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ by Lemma 1.12.
+By Theorem 4.7, we have $\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$.
 Let $f\in\overline{\Delta}$ be a deep computation.
@@
 Define $\tilde{f}:=\mathcal{U}\lim_{i}\tilde{\gamma_{i}}$.
-Then, for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$ $\pi_{P}\circ\tilde{\gamma}_{i}|_{\mathcal{L}[r_{\bullet}]}\in M_{r}(\mathcal{L}[r_{\bullet}])$ for all $i$, so $\pi_{P}\circ f|_{\mathcal{L}[r_{\bullet}]}\in\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$.
+Then, $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ extends $f$.
+Since $\Delta$ is $R$-confined we have that $f:L[r_{\bullet}]\rightarrow L[r_{\bullet}]$ and $\tilde{f}:\mathcal{L}[r_{\bullet}]\rightarrow\mathcal{L}[r_{\bullet}]$ for all $r_{\bullet}\in R$.
+Lastly, note that for all $r_{\bullet}\in R$ and $P\in\mathcal{P}$ we have that $\pi_{P}\circ\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq M_{r}(\mathcal{L}[r_{\bullet}])$.
 ∎
-Question 4.8.
+Question 4.9.
 Under the same assumptions of the preceding theorem, suppose that every deep computation of $\Delta$ is universally Monte Carlo computable.
 Must $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ have the NIP for all $P\in\mathcal{P}$ and all $r_{\bullet}\in R$?
-4.4. Talagrand stability, Fremlin’s dichotomy, NIP, and universal Monte Carlo computability of deep computations
-There is another notion closely related to NIP, introduced by Talagrand in [talagrand1984pettis] while studying Pettis integration.
+4.2. Talagrand stability and Monte Carlo computability of deep computations
+There is another notion closely related to NIP, introduced by Talagrand in [TAL84] while studying Pettis integration.
 Suppose that $X$ is a compact Hausdorff space and $A\subseteq\mathbb{R}^{X}$.
@@
 we write:
-$D_{k}(A,E,a,b)=\bigcup_{f\in A}\{x\in E^{2k}:f(x_{2i})\leq a,\hskip 2.84526ptf(x_{2i+1})\geq b\hskip 2.84526pt{\text{ for all }i<k}\}$
+$D_{k}(A,E,a,b)=\bigcup_{f\in A}\{x\in E^{2k}:f(x_{2i})\leq a,\hskip 2.84526ptf(x_{2i+1})\geq b\hskip 2.84526pt{\text{ for all }i<k}\}.$
 We say that $A$ is Talagrand $\mu$-stable if and only if for every $\mu$-measurable set $E\subseteq X$ of positive measure and for every $a<b$ there is $k\geq 1$ such that
 $(\mu^{2k})^{*}(D_{k}(A,E,a,b))<(\mu(E))^{2k},$
-where $\mu*$ denotes the outer measure (we work with outer since the sets $D_{k}(A,E,a,b)$ need not be $\mu$-measurable).
-This is certainly the case when $A$ is a countable set of continuous (or $\mu$-measurable) functions.
-Notation 4.9.
-For a measure $\mu$ on a set $X$, the set of all $\mu$-measurable functions will denoted by $\mathscr{M}^{0}(X,\mu)$.
-The following lemma establishes that Talagrand stability is a way to ensure that deep computations are definable by measurable functions.
+where $\mu^{*}$ denotes the outer measure (we need to work with outer measure since the sets $D_{k}(A,E,a,b)$ need not be $\mu$-measurable).
+The inequality certainly holds when $A$ is a countable set of continuous (or $\mu$-measurable) functions.
+The main result of this section is that accumulation points of standard computations (i.e., deep computations) are Monte Carlo computable; this is Theorem 4.14 below.
+We now prove that deep computations that are Talagrand $\mu$-stable are $\mu$-measurable.
+But first, let us state the following useful characterization of measurable functions (compare with Fact 1.2):
+Fact 4.10 (Lemma 413G in [FRE03]).
+Suppose that $(X,\Sigma,\mu)$ is a measure space and $\mathcal{K}\subseteq\Sigma$ is a collection of measurable sets satisfying the following conditions:
+  1. (1)
+    $(X,\Sigma,\mu)$ is complete, i.e., for all $E\in\Sigma$ with $\mu(E)=0$ and $F\subseteq E$ we have $F\in\Sigma$.
+  2. (2)
+    $(X,\Sigma,\mu)$ is semi-finite, i.e., for all $E\in\Sigma$ with $\mu(E)=\infty$ there exists $F\subseteq E$ such that $F\in\Sigma$ and $0<\mu(F)<\infty$.
+  3. (3)
+    $(X,\Sigma,\mu)$ is saturated, i.e., $E\in\Sigma$ if and only if $E\cap F\in\Sigma$ for all $F\in\Sigma$ with $\mu(F)<\infty$.
+  4. (4)
+    $(X,\Sigma,\mu)$ is inner regular with respect to $\mathcal{K}$, i.e., for all $E\in\Sigma$ μ(E)=sup{μ(K):$K$ and KE}.
+(In particular, if $X$ is Hausdorff, $\mu$ is a Radon measure on $X$ and $\mathcal{K}$ is the collection of compact subsets of $X$, all these conditions hold).
+Then, $f:X\rightarrow\mathbb{R}$ is measurable if and only if for every $K\in\mathcal{K}$ with $0<\mu(K)<\infty$ and $a<b$, either $\mu^{*}(P)<\mu(K)$ or $\mu^{*}(Q)<\mu(K)$ where $P=\{x\in K:f(x)\leq a\}$ and $Q=\{x\in K:f(x)\geq b\}$.
+The following technical lemma will be instrumental to prove of Proposition 4.13, which, in turn, will yield the main result of the subsection, Theorem 4.14.
 Lemma 4.11.
@@
 Proof.
-First, observe that a subset of a $\mu$-stable set is $\mu$-stable.
-To show that $\overline{A}$ is $\mu$-stable, observe that $D_{k}(\overline{A},E,a,b)\subseteq D_{k}(A,E,a^{\prime},b^{\prime})$ where $a<a^{\prime}<b^{\prime}<b$ and $E$ is a $\mu$-measurable set with positive measure.
+First, we claim that a subset of a $\mu$-stable set is $\mu$-stable.
+To see this, suppose that $A\subseteq B$ and $B$ is $\mu$-stable.
+Fix any $\mu$-measurable $E\subseteq X$ of positive measure and $a<b$.
+Let $k\geq 1$ such that
+$(\mu^{2k})^{*}(D_{k}(B,E,a,b))<(\mu(E))^{2k}.$
+Since $A\subseteq B$, we have $D_{k}(A,E,a,b)\subseteq D_{k}(B,E,a,b)$; therefore,
+$(\mu^{2k})^{*}(D_{k}(A,E,a,b))\leq(\mu^{2k})^{*}(D_{k}(B,E,a,b))<(\mu(E))^{2k}.$
+We now show that $\overline{A}$ is $\mu$-stable.
+Fix $E\subseteq X$ measurable with positive measure and $a<b$.
+Let $a^{\prime}<b^{\prime}$ be such that $a<a^{\prime}<b^{\prime}<b$.
+Since $A$ is $\mu$-stable, let $k\geq 1$ be such that
+$(\mu^{2k})^{*}(D_{k}(A,E,a^{\prime},b^{\prime}))<(\mu(E))^{2k}.$
+If $x\in D_{k}(\overline{A},E,a,b)$, then there is $f\in\overline{A}$ such that $f(x_{2i})\leq a<a^{\prime}$ and $f(x_{2i+1})\geq b>b^{\prime}$ for all $i<k$.
+By definition of pointwise convergence topology, there exists $g\in A$ such that $g(x_{2i})<a^{\prime}$ and $g(x_{2i+1})>b^{\prime}$ for all $i<k$.
+Hence, $x\in D_{k}(A,E,a^{\prime},b^{\prime})$.
+We have shown that $D_{k}(\overline{A},E,a,b)\subseteq D_{k}(A,E,a^{\prime},b^{\prime})$; hence,
+$(\mu^{2k})^{*}(D_{k}(\overline{A},E,a,b))\leq(\mu^{2k})^{*}(D_{k}(A,E,a^{\prime},b^{\prime}))<(\mu(E))^{2k}.$
 It suffices to show that $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$.
 Suppose that there exists $f\in\overline{A}$ such that $f\notin\mathscr{M}^{0}(X,\mu)$.
-By a characterization of measurable functions (see 413G in [fremlin2003vol4]), there exists a $\mu$-measurable set $E$ of positive measure and $a<b$ such that $\mu^{*}(P)=\mu^{*}(Q)=\mu(E)$ where $P=\{x\in E:f(x)\leq a\}$ and $Q=\{x\in E:f(x)\geq b\}$.
-Then, for any $k\geq 1$: $(P\times Q)^{k}\subseteq D_{k}(\{f\},E,a,b)$ so $(\mu^{2k})^{*}(D_{k}(\{f\},E,a,b))=(\mu^{*}(P)\mu^{*}(Q))^{k}=(\mu(E))^{2k}$.
-Thus, $\{f\}$ is not $\mu$-stable, but we argued before that a subset of a $\mu$-stable set must be $\mu$-stable.
+By fact 4.10, there exists a $\mu$-measurable set $E$ of positive measure and $a<b$ such that $\mu^{*}(P)=\mu^{*}(Q)=\mu(E)$ where $P=\{x\in E:f(x)\leq a\}$ and $Q=\{x\in E:f(x)\geq b\}$.
+Then, for any $k\geq 1$: $(P\times Q)^{k}\subseteq D_{k}(\{f\},E,a,b)$, so $(\mu^{2k})^{*}(D_{k}(\{f\},E,a,b))=(\mu^{*}(P)\mu^{*}(Q))^{k}=(\mu(E))^{2k}$.
+Thus, $\{f\}$ is not $\mu$-stable.
+However, we argued above that a subset of a $\mu$-stable set must be $\mu$-stable, so we have a contradiction.
 ∎
-We say that $A$ is universally Talagrand stable if $A$ is Talagrand $\mu$-stable for every Radon probability measure $\mu$ on $X$. An argument similar to the proof of 4.5, yields the following:
-Theorem 4.11.
-Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom.
-If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ is universally Talagrand stable for all $P\in\mathcal{P}$ and all sizers $r_{\bullet}$, then every deep computation is universally Monte Carlo computable.
-It is then natural to ask: what is the relationship between Talagrand stability and the NIP? The following dichotomy will be useful.
-Lemma 4.12 (Fremlin’s Dichotomy, 463K in [fremlin2003vol4]).
-If $(X,\Sigma,\mu)$ is a perfect $\sigma$-finite measure space (in particular, for $X$ compact and $\mu$ a Radon probability measure on $X$) and $\{f_{n}:n\in\mathbb{N}\}$ be a sequence of real-valued measurable functions on $X$, then one (and only one) of the following conditions holds:
-  1. (i)
-    $\{f_{n}:n\in\mathbb{N}\}$ has a subsequence that converges $\mu$-almost everywhere,
-  2. (ii)
-    $\{f_{n}:n\in\mathbb{N}\}$ has a subsequence with no $\mu$-measurable accumulation point in $\mathbb{R}^{X}$.
-The preceding lemma can be considered as a measure-theoretic version of Rosenthal’s dichotomy. Combining this dichotomy with Theorem 4.5, we get the following result:
-Theorem 4.13.
-Let $X$ be a Hausdorff space and $A\subseteq C(X)$ be pointwise bounded. The following are equivalent:
-  1. (i)
-    $\overline{A}\subseteq M_{r}(X)$.
-  2. (ii)
-    For every compact $K\subseteq X$, $A|_{K}$ satisfies the NIP.
-  3. (iii)
-    For every Radon measure $\mu$ on $X$, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$, i.e., every countable subset of $A$ has an accumulation point in $\mathscr{M}^{0}(X,\mu)$.
-  4. (iv)
-    For every Radon measure $\mu$ on $X$ and every sequence $\{f_{n}:n\in\mathbb{N}\}\subseteq A$, there is a subsequence that converges $\mu$-almost everywhere.
-Proof.
-Notice that the equivalence (i)-(iii) is Theorem 4.5. Notice that the equivalence of (iii) and (iv) is Fremlin’s Dichotomy (Theorem 4.12). ∎
-Finally, it is natural to ask what the connection is between Talagrand stability and NIP.
+Definition 4.12.
+We say that $A$ is universally Talagrand stable if $A$ is Talagrand $\mu$-stable for every Radon probability measure $\mu$ on $X$.
+We first observe that universal Talagrand stability corresponds to a complexity class smaller than or equal to the NIP:
 Proposition 4.13.
@@
 Proof.
-By Theorem 4.5, it suffices to show that $A$ is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$ for all Radon probability measure $\mu$ on $X$.
-Since $A$ is Talagrand $\mu$-stable for any such $\mu$, we have $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$.
-In particular, $A$ is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$.
+By Theorem 4.7, it suffices to show that $A$ is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$ for every Radon probability measure $\mu$ on $X$.
+Since $A$ is Talagrand $\mu$-stable for any such $\mu$, we have $\overline{A}\subseteq\mathscr{M}^{0}(X,\mu)$ by Lemma 4.11.
+In particular, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$.
 ∎
-Question 4.15.
-Is the converse true?
-The following two results suggest that the precise connection between Talagrand stability and NIP may be sensitive to set-theoretic axioms (even assuming countability of $A$).
-Theorem 4.16 (Talagrand, Theorem 9-3-1(a) in [talagrand1984pettis]).
+Corollary 4.14.
+Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendibility Axiom.
+If $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ is universally Talagrand stable for all $P\in\mathcal{P}$ and all sizers $r_{\bullet}$, then every deep computation is universally Monte Carlo computable.
+Proof.
+This is a direct consequence of Proposition 4.13 and Theorem 4.8. ∎
+In the context of deep computations, we have identified two ways to obtain Monte Carlo computability, namely, NIP/PAC and Talagrand stability.
+It is natural to ask whether these two notions are equivalent.
+The following results show that, even in the simple case of countably many computations, this question is sensitive to the set-theoretic axioms.
+On the one hand, it is consistent (with respect to the standard ZFC axioms of set theory) that these two classes are the same:
+Theorem 4.15 (Talagrand, Theorem 9-3-1(a) in [TAL84]).
 Let $X$ be a compact Hausdorff space and $A\subseteq M_{r}(X)$ be countable and pointwise bounded.
@@
 If $A$ satisfies the NIP, then $A$ is universally Talagrand stable.
-Theorem 4.17 (Fremlin, Shelah, [fremlin1993pointwise]).
+(The assumption that $[0,1]$ is not the union of $\mathfrak{c}$ closed measure zero sets is a consequence of, for example, the Continuum Hypothesis.)
+On the other hand, by fixing a particular well-known measure, the Lebesgue measure, we see that the other case is also consistent:
+Theorem 4.16 (Fremlin, Shelah, [FS93]).
 It is consistent with the usual axioms of set theory that there exists a countable pointwise bounded set of Lebesgue measurable functions with the NIP which is not Talagrand stable with respect to Lebesgue measure.
+Notice that the preceding two results apply to sets of measurable functions, a class of functions larger than the class of continuous functions.
+It is possible that the continuity assumption, i.e., the Extendibility Axiom, settles the problem.
+This is a problem that remains open for future work.

### v3

@@
 Introduction
-In this paper we study asymptotic behavior of computations, e.g., the depth of a neural network tending to infinity, or the time interval between layers of a time-series network tending toward zero.
+In this paper we study asymptotic behavior of computations as parameters of the computation tend towards a limit, e.g., the depth of a neural network tending to infinity, or the time interval between layers of a time-series network tending toward zero.
 Recently, particular cases of this concept have attracted considerable attention in deep learning research (e.g., Neural Ordinary Differential Equations [8], Physics-Informed Neural Networks [RPK19], and deep equilibrium models [4]).
@@
 The viewpoint of spaces of types, which we have borrowed from model theory, thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: In topology, the classification of Rosenthal compacta pioneered by Todorčević [TOD99]; in logic, the classification of theories developed by Shelah [SHE90]; and in statistical learning, the notions of PAC learning and VC dimension pioneered by Vapnik and Chervonenkis [VC74, VC71].
-In a previous paper [ADI+24], we introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]).
+In a previous paper [ADI+24], the authors introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]).
 There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
-In [ADI+24], we proved a new “tame vs wild” (i.e., polynomially approximable vs nonpolynimially approximable) dichotomy for the complexity of deep computations by invoking a classical result of Grothendieck from the 1950s [GRO52].
-Under our model-theoretic Rosetta stone, the property of polynomial approximability of computations is identified with continuous extendibility in the sense of topology, and with the notions of stability and type definability in model theory.
-Simplest among deep computations are those arising as pointwise limits of (continuous) computations proper.
+In [ADI+24], a new “tame vs wild” (i.e., polynomially approximable vs nonpolynomially approximable) dichotomy for the complexity of deep computations was proved by invoking a classical result of Grothendieck from the 1950s [GRO52].
+Under our model-theoretic Rosetta stone, the property of polynomial approximability of computations is identified with continuous extendibility in the sense of topology, and with the notions of stability and type definability in the sense of model theory.
+Deep computations arise as topological limits of standard (continuous) computations.
+Simplest among these is when the limit is sequential.
 In topology, the first Baire class, or Baire class 1 consists of functions (also called simply “Baire-1”) arising as pointwise limits of sequences of continuous functions.
-Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity away from the Baire class 0 which (by definition) consists of continuous functions.
-We prove a new “tame vs wild” Ramsey-theoretic dichotomy for complexity of general deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT78], and a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [TOD99].
-Todorčević’s trichotomy regards Rosenthal compacta; these are special classes of topological spaces, defined as compact spaces that can be embedded (homeomorphically identified as a subset) within the space of Baire class 1 functions on some Polish (separable, completely metrizable) space, under the topology of pointwise convergence.
-Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways; since the late 70’s, they have played a crucial role in understanding the complexity of structures of functional analysis, especially Banach spaces.
-Todorčević’s trichotomy has been utilized to settle longstanding problems in topological dynamics and topological entropy [GM22].
+Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity above the Baire class 0 which (by definition) consists of continuous functions.
+In this paper, we apply the theory of spaces of Baire class 1 functions to identify and separate new complexity classes of deep computations.
+We obtain a new “tame vs wild” dichotomy for deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT78], and further, a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [TOD99].
+Todorčević’s trichotomy concerns Rosenthal compacta, which are topological spaces that fit our context superbly: they are defined as compact spaces that are homeomorphic to subspaces of the space of Baire class 1 functions on some Polish (separable, completely metrizable) space, under the topology of pointwise convergence.
+Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways.
+Since the late 70’s, they have played a crucial role in understanding the complexity of structures of functional analysis, especially Banach spaces.
+Todorčević’s trichotomy for Rosenthal compacta has been utilized to settle longstanding problems in topological dynamics and topological entropy [GM22].
 It is noteworthy that Todorčević’s proof relies on sophisticated set-theoretic forcing and infinite Ramsey theory.
 At the time of writing this paper, decades after his original argument, no elementary proof has been found [TOD23, HT19].
-Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of “Non Independence Property” (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concept of Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory identified by Valiant [VAL84].
-Going beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]; they identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
-They also showed that if a separable Rosenthal compactum is not hereditarily separable, then it must contain an uncountable discrete subspace of the size of the continuum.
+Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of Non Independence Property (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concepts VC-dimension and Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory [VC71, VC74, VAL84].
+Refining Todorčević’s trichotomy, we invoke a more recent heptachotomy for separable Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]; these authors identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
 We believe that the results presented in this paper show practitioners of computation, or topology, or descriptive set theory, or model theory, how classification invariants used in their field translate into classification invariants of other fields.
-In the interest of accessibility, we do not assume the reader to have previous familiarity with advanced topology, model theory, or computing.
+In the interest of accessibility, we do not assume that the reader has previous familiarity with advanced topology, model theory, or computing.
 The only technical prerequisites to read this paper are undergraduate-level topology and measure theory.
@@
 In this section we present some preliminaries from general topology and function space theory.
-In the interest of completeness, we some proofs that may be safely skipped by readers familiar with these topics.
+In the interest of completeness, we include some proofs that may be safely skipped by readers familiar with these topics.
 Recall that a subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets.
@@
 A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions.
+In symbols, $f:X\to Y$ is Baire class 1 if there is a sequence of continuous functions $f_{n}:X\to Y$ such that for all $x\in X$, $f(x)=\lim_{n\to\infty}f_{n}(x)$.
 If $X$ and $Y$ are topological spaces, the space of Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$ (as above, $B_{1}(X,\mathbb{R})$ is denoted $B_{1}(X)$).
-Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)$.
+Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)\subseteq Y^{X}$ and we give these the topology (called the topology of pointwise convergence) inherited from the product topology of $Y^{X}$.
 The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles.
 His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
+An elementary fact about Baire class 1 functions is that they are continuous except on a set of first category (also called a meager set, a set of first cateogory is the countable union of sets whose closure has empty interior; intuitively, these sets are “topologically small”).
+Thus, Baire class 1 functions are continuous on a “topologically large” subset of their domain.
 A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ (equivalently, every open subset of $X$ is a $F_{\sigma}$).
@@
 A topological space $X$ is Baire if every countable intersection of dense open sets is dense.
-The Baire Category Theorem states that every Hausdorff compact or completely metrizable space (hence, every Polish space) is Baire.
-The following fact was established by Baire in his 1899 thesis. A proof can be found in [TOD97, Section 10].
+The Baire Category Theorem states that every compact Hausdorff or completely metrizable space (hence, every Polish space) is Baire.
+The following fact was established by Baire in his groundbreaking thesis. A proof can be found in [TOD97, Section 10].
 Fact 1.2 (Baire).
@@
   • (2)
-    $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ whenever $U\subseteq Y$ is open.
-If, mooreover, $X$ is Baire, then (1) and (2) are equivalent to:
+    $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ whenever $U\subseteq\mathbb{R}$ is open.
+If, moreover, $X$ is Baire, then (1) and (2) are equivalent to:
   • (3)
@@
 A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact.
-Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) are of interest analysis and topological dynamics.
+Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) are of interest in analysis and topological dynamics.
 We begin with the following well-known result.
@@
   2. (ii)
-    AA is relatively countably compact in $B_{1}(X)$, i.e., every countable subset of $A$ has an accumulation point in $B_{1}(X)$.
+    AA is relatively countably compact in $B_{1}(X)$, i.e., every countable subset of $A$ has a limit point in $B_{1}(X)$.
   3. (iii)
@@
 For more complicated spaces, such as $C_{p}$-spaces, this fails in remarkable ways.
-The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$, and one can show (e.g., [TOD97, Chapter 1.1]) that $\{p_{n}\}_{n\in\mathbb{N}}$ is closed, but has no convergent subsequences, in $Y=C_{p}(X)$: in a sense, such example exhibits the worst failure of sequential convergence possible.
-The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) realizes the Stone-Čech compactification $\beta\mathbb{N}$ of the discrete space of natural numbers; it is an important object of study in general topology.
-The following theorem, proved by Haskell Rosenthal in 1974, is fundamental in functional analysis and captures a sharp division in the behavior of sequences in a Banach spaces.
+The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$, and one can show (e.g., [TOD97, Chapter 1.1]) that $\{p_{n}\}_{n\in\mathbb{N}}$ has no convergent subsequences, in $\mathbb{R}^{X}$.
+In a sense, this example exhibits the worst failure of sequential convergence possible.
+The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) is homeomorphic to the Stone-Čech compactification of the discrete space of natural numbers, usually denoted $\beta\mathbb{N}$, which is an important object of study in general topology.
+The following theorem, proved by Haskell Rosenthal in 1974, is fundamental in functional analysis and captures a sharp division in the behavior of sequences in a Banach space.
 Theorem 1.4 (Rosenthal’s Dichotomy, [ROS74]).
@@
 Fix $A\subseteq C_{p}(X)$, a sequence $\{f_{n}\}\subseteq A$ and reals $r<s$.
-For any $I,J\subseteq\mathbb{N}$ (eventually disjoint in applications), write $X_{I,J}$ for $X^{f_{I}}_{\leq r}\cap X^{f_{J}}_{\geq s}$.
+For any $I,J\subseteq\mathbb{N}$ (almost disjoint in applications), write $X_{I,J}$ for $X^{f_{I}}_{\leq r}\cap X^{f_{J}}_{\geq s}$.
 For $I\subseteq I^{\prime}\subseteq\mathbb{N}$ and $J\subseteq J^{\prime}\subseteq\mathbb{N}$, we have $X_{I,J}\supseteq X_{I^{\prime},J^{\prime}}$; moreover, $X_{I,J}=\bigcap_{E\subseteq I,F\subseteq J}X_{E,F}$, where the index variables $E\subseteq I$, $F\subseteq J$ range over finite subsets of $I,J$, respectively.
@@
 Theorem 1.8.
-Let $X$ be a metrizable compact (hence Polish) space. For every pointwise bounded $A\subseteq C_{p}(X)$, the following properties are all equivalent:
+Let $X$ be a compact metrizable (hence Polish) space. For every pointwise bounded $A\subseteq C_{p}(X)$, the following properties are all equivalent:
   1. (i)
@@
 The “polynomial” name is justified: indeed, for fixed $N>0$, as a function of the size $|A|=m>0$, we have
-$\sum_{i=0}^{N-1}\binom{m}{i}\leq\sum_{i=0}^{N-1}\frac{m^{i}}{i!}\leq\left(\sum_{i=0}^{N-1}\frac{1}{i!}\right)\cdot m^{N-1}\cdot<e\cdot m^{N-1}=O\left(m^{N}\right).$
+$\sum_{i=0}^{N-1}\binom{m}{i}\leq\sum_{i=0}^{N-1}\frac{m^{i}}{i!}\leq\left(\sum_{i=0}^{N-1}\frac{1}{i!}\right)\cdot m^{N-1}<e\cdot m^{N-1}=O\left(m^{N}\right).$
 (More precisely, the order of magnitude is $O(m^{N-1})$: polynomial in $m$ for $N$ fixed.)
@@
 Proof.
-The identification of $C_{p}$-spaces follows trivially from the definition of topological product and the fact that $\mathcal{P}$ is discrete: a continuous map $X\to\mathbb{R}^{\mathcal{P}}$ is precisely a $\mathcal{P}$-tuple of continuous functions $X\to\mathbb{R}$, and these correspond to continuous functions $X\times\mathcal{P}\to\mathbb{R}$.
+The identification of $C_{p}$-spaces follows trivially from the definition of topological product and the fact that $\mathcal{P}$ is discrete: a continuous map $X\to\mathbb{R}^{\mathcal{P}}$ is precisely a $\mathcal{P}$-indexed family of continuous functions $X\to\mathbb{R}$, and these correspond to continuous functions $X\times\mathcal{P}\to\mathbb{R}$.
 The identification of Baire-1 spaces follows immediately, since it is defined in terms of the purely topological notion of limit (in the ambient space) of sequences of continuous functions.
@@
   1. (i)
-    $A_{L}$ satisfies the NIP.
+    $A_{L}$ satisfies the NIP;
   2. (ii)
@@
 This contradicts (i). ∎
-2. Compositional Computation Structures: A structural approach to floating-point computations
-In this section, we connect function spaces with floating point computations. We start by summarizing some basic concepts from [ADI+24].
+2. Compositional Computation Structures: A structural approach to arbitrary-precision arithmetic
+In this section, we connect function spaces with arbitrary-precision arithmetic computations. We start by summarizing some basic concepts from [ADI+24].
 A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates.
@@
 Each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$.
-Important space states are $L=\mathbb{R}^{\mathbb{N}}$ and $L=\mathbb{R}^{n}$ for some positive integer $n$, endowed with predicate $P_{i}(v)=v_{i}$, one each for the $i$-th coordinate of $v$.
+Important state spaces are $L=\mathbb{R}^{\mathbb{N}}$ and $L=\mathbb{R}^{n}$ for some positive integer $n$, endowed with predicate $P_{i}(v)=v_{i}$, one each for the $i$-th coordinate of $v$.
 We regard the space of types as a topological space, endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$.
@@
 Elements of $\mathcal{L}\setminus L$ will be called unrealized types.
-Intuitively, state types capture a notion of “ultrastate”.
-The traditional setting of discrete first-order logic may be regarded as the special case in which predicates are binary (such predicates may take only the values 0 and $1$, say).
-In such a traditional setting, the space of types of a structure is necessarily compact; for some purposes, it serves as a sort of compactification of the structure itself.
-Compactness in such setting is critical.
-However, predicates of a computational states structure take values in the non-compact space $\mathbb{R}$, so the corresponding type space $\mathcal{L}$ need not be compact.
-To bypass this obstacle, we follow the idea, introduced in [ADI+24], of covering $\mathcal{L}$ by “thin” compact subspaces called shards.
+Intuitively, state types capture a notion of “limit state”.
+As we combine ideas of model theory [KEI03] and topology [BFT78], we are interested in families of real-valued functions that are pointwise bounded.
+This leads us to the concepts of sizer and shard introduced first in introduced in [ADI+24]:
 Definition 2.2.
-A sizer is a tuple $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$.
+A sizer is a family $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$.
 Given a sizer $r_{\bullet}$, let $\mathbb{R}^{[r_{\bullet}]}=\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of a states space $L$ be
@@
 For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$ (a closed, hence compact subset of $\mathbb{R}^{[r_{\bullet}]}$).
-Let also $\mathcal{L}_{\text{sh}}$ be the union of all type-shards (as the sizer $r_{\bullet}$ varies).
-In general, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$ may be a proper subset. Equality holds in the important special case when $\mathcal{P}$ is countable (see *** below).
+Let also $\mathcal{L}_{\text{sh}}$ be the union of all type-shards as the sizer $r_{\bullet}$ varies.
+In general, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$, and the inclusion may be proper. However, equality holds in the important special case when $\mathcal{P}$ is countable (see [ADI+24]).
 2.1. Compositional Computation Structures
@@
 The main goal of this paper is to study the computability, definability and computational complexity of deep computations.
-Since ultracomputations are defined through a combination of topological concepts (namely, topological closure) and structural and model-theoretic concepts (namely, models and types), we will import technology from both topology and model theory.
+Since deep computations are defined through a combination of topological concepts (namely, topological closure) and structural and model-theoretic concepts (namely, models and types), we will import technology from both topology and model theory.
 2.2. Computability and definability of deep computations and the Extendibility Axiom
@@
   1. (1)
-    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{L}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ and by a finite number of iterations of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
+    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{R}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ and by a finite number of iterations of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
     $|\varphi_{Q,\,K,\varepsilon}(v)-Q{\circ}f(v)|\leq\epsilon,\qquad\text{for all $v\in K$}.$
-    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model the theory of real-valued structures first introduced in [IOV94] for model theory of functional analysis and now standard in model theory (see [KEI03]).
+    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model theory of real-valued structures first introduced in [IOV94] for model theory of functional analysis and now standard in model theory (see [KEI03]).
 The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
@@
 To summarize, a function $f:L\to\mathcal{L}$ is computable if and only if it is definable if and only if it is polynomially approximable if and only it can be extended to a continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$.
-This motivates the following definition.
+This is the reason for the following definition.
 Definition 2.5.
@@
 We refer to $\tilde{\gamma}$ as a free extension of $\gamma$.
-By the preceding remarks, the Extendibility Axiom says that the elements of the semigroup $\Gamma$ are finitary computations.
 For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$.
@@
 2.3. Newton’s method as a CCS
-Let $p(z)$ be a non-constant polynomial with complex coefficients. Newton’s method is an iterative method that is used to approximate a root of $p(z)$. Define the Newton map as:
+Let $p(z)$ be a non-constant polynomial with complex coefficients. We say that $(L_{p},\mathcal{P},\Gamma_{p})$ is Newton’s method for $p(z)$ if:
+  • $L_{p}$ is the set of all $z\in\mathbb{C}$ such that there exists an open neighborhood $U$ of $z$ such that every sequence in $\{N_{p}^{n}:n\in\mathbb{N}\}$ has a subsequence that converges uniformly on compact subsets of $U$, where
     $N_{p}(z)=z-\frac{p(z)}{p^{\prime}(z)}$
+    and $N_{p}^{n}:=N_{p}\circ N_{p}\circ\cdots\circ N_{p}$ is the nth-iteration of $N_{p}$.
+  • $\mathcal{P}:=\{P_{1},P_{2},P_{3}\}$ where
+    $P_{1}(z)=\frac{2{\rm Re}(z)}{|z|^{2}+1},$
+    $P_{2}(z)=\frac{2{\rm Im}(z)}{|z|^{2}+1},$
+    $P_{3}(z)=\frac{|z|^{2}-1}{|z|^{2}+1}.$
+  • $\Gamma_{p}:=\{N_{p}^{n}:n\in\mathbb{N}\}$.
+Remarks 2.6.
+  1. (1)
+    The set $L_{p}$ is known as the Fatou set of $N_{p}$ (see section 2 in [BLA84]). Its complement $\mathbb{C}\backslash L_{p}$, is known as the Julia set of $N_{p}$.
+  2. (2)
+    The map $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$ is the stereographic projection into the Riemann Sphere $S^{2}$.
+  3. (3)
+    The set $L_{p}$ is open and dense in $\mathbb{C}$ ([BLA84], Corollary 4.6).
+Hence, its closure $\mathcal{L}$ in $\mathbb{R}^{\mathcal{P}}$ is the Riemann sphere $S^{2}$ (i.e., the extended complex plane).
+  4. (4)
+    The set $L_{p}$ is completely invariant under iterations of $N_{p}$, i.e., $N_{p}(L_{p})=L_{p}$ and $N_{p}^{-1}(L_{p})=L_{p}$.
+This implies that all iterations $N_{p}^{n}:L_{p}\rightarrow L_{p}$ are transition maps.
+  5. (5)
+    $\Gamma_{p}$ is the semigroup generated by $\{N_{p}\}$. Thus, $(L_{p},\mathcal{P},\Gamma_{p})$ is a CCS.
+Newton’s method is an iterative method that is used to approximate a root of $p(z)$.
+The map $N_{p}(z)$ defined above is known as Newton’s map.
 The method consists of taking an initial guess $z_{0}\in\mathbb{C}$ and iterating the rational map $N_{p}$ to obtain a sequence given by
 $z_{n+1}=N_{p}(z_{n})=z_{n}-\frac{p(z_{n})}{p^{\prime}(z_{n})}$
-For each root $r\in\mathbb{C}$ of $p(z)$, there exists $\varepsilon>0$ such that for any initial guess $z_{0}$ in the $\varepsilon$-ball centered at $r$, Newton’s iteration converges to $r$ (provided $p^{\prime}(r)\neq 0$) and the convergence is quadratic in that case, meaning the error at each step is roughly squared, causing the number of correct digits to double, leading to fast convergence.
-Denote by $N_{p}^{n}:=N_{p}\circ N_{p}\circ\cdots\circ N_{p}$ (nth iteration of $N_{p}$).
+For each root $r\in\mathbb{C}$ of $p(z)$, there exists an $\varepsilon>0$ such that for any initial guess $z_{0}$ in the $\varepsilon$-ball centered at $r$, Newton’s iteration converges to $r$ (provided $p^{\prime}(r)\neq 0$) and the convergence is quadratic in that case, meaning the error at each step is roughly squared, causing the number of correct digits to double, leading to fast convergence.
 Given a root $r$ of $p(z)$, the set $B_{r}=\{z\in\mathbb{C}:\lim_{n\rightarrow\infty}N_{p}^{n}(z)=r\}$ is an open set called the basin of $r$.
@@
 Another more chaotic way Newton’s method can fail to converge is when the sequence of iterations has no convergent subsequence.
-The set of such points is known as the Julia set associated to $N_{p}$ and it is typically a fractal.
-This can be visualizaed by adding a dash of color: let us give each complex number $z_{0}$ a color $(R,G,B)$ where $R,G,B\in[0,1]$ (so that $(1,0,0)$ is red, $(0,1,0)$ is green, $(0,0,1)$ is blue and $(0.5,0,0.5)$ is a light purple, for example).
+The set of such points, i.e.
+the Julia set associated to $N_{p}$, is typically a fractal.
+This can be visualized by adding a dash of color: let us give each complex number $z_{0}$ a color $(R,G,B)$ where $R,G,B\in[0,1]$ (so that $(1,0,0)$ is red, $(0,1,0)$ is green, $(0,0,1)$ is blue and $(0.5,0,0.5)$ is a light purple, for example).
 The values of $R$, $G$ and $B$ are determined by looking at the image of said number at each stage of the iteration, $N_{p}^{n}(z_{0})$, and computing the current distance to each of the roots of $p(z)$; so $R=1/d_{r}$ where $d_{r}$ is the positive distance to the root which is colored red, and so on.
@@
 As the number of iterations increases, each point gets a sharper color, as the sequence of images $\{N_{p}^{n}(z_{0})\}_{n=1}^{\infty}$ converges to one of the three roots.
-Each stage, the complex plane looks as if out of focus because the coloring function is continuous.
-As the reader can see in Figure 1, the points at the boundary of each color class form the famous Newton fractal (of which, interestingly, Newton was unaware).
+At each stage, the complex plane looks as if out of focus because the coloring function is continuous.
+As the reader can see in Figure 1, the points at the boundary of each color class form the famous Newton’s fractal (of which, interestingly, Newton was unaware).
 (a) After 1 iteration
@@
 Figure 2. Newton’s method approximating $p(z)=z^{3}-1$.
-We can frame Newton’s method as a CCS (satisfying the Extendibility Axiom) as follows.
-Given a non-constant polynomial $p(z)$, let $L\subseteq\mathbb{C}$ be the complement of the Julia set (this is called the Fatou set) associated to $N_{p}$.
-This is the set where Newton’s method is tame.
-By [BLA84, Corollary 4.6], $L$ is open and dense in $\mathbb{C}$.
-Let $\mathcal{P}=\{P_{1},P_{2},P_{3}\}$ where the map $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$ is the stereographic projection into the Riemann Sphere $S^{2}$, i.e.,
-$P_{1}(z)=\frac{2{\rm Re}(z)}{|z|^{2}+1},$
-$P_{2}(z)=\frac{2{\rm Im}(z)}{|z|^{2}+1},$
-$P_{3}(z)=\frac{|z|^{2}-1}{|z|^{2}+1}.$
-Since $L$ is dense in the extended complex plane, then $\mathcal{L}=S^{2}$.
-Computations are the iterations of Newton’s method, i.e., $\Delta=\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$.
-Then, $(L,\mathcal{P},\Gamma)$ is a CCS.
-Since all computations $N_{p}^{n}$ are rational maps, they can be continuously extended to the extended complex plane $\overline{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, i.e., to $\mathcal{L}$.
-In particular, $(L,\mathcal{P},\Gamma)$ satisfy the Extendibility Axiom.
-The set of deep computations $\overline{\Delta}$ might behave different for various polynomials.
-Let us look at various examples:
+Proposition 2.7.
+Let $p(z)$ be a non-constant polynomial. $(L_{p},\mathcal{P},\Gamma_{p})$ satisfies the Extendibility Axiom.
+Proof.
+$N_{p}:L_{p}\rightarrow L_{p}$ is a rational map.
+Rational maps can be continuously extended to the extended complex plane, i.e., to $\mathcal{L}$.
+Composition of rational maps is a rational map, so by the same reasoning, computations $N_{p}^{n}:L_{p}\rightarrow L_{p}$ can be continuously extended to $\mathcal{L}$.
+∎
+The set of deep computations $\overline{\Gamma}$ might behave different for various polynomials. Let us look at various examples:
 Example 2.8.
@@
 $P_{2}(x)=\frac{x^{2}-1}{x^{2}+1}.$
-The Newton’s method map $N_{p}:L\rightarrow L$ is given by
+Let $\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$ where
 $N_{p}(x)=\frac{x^{2}+a}{2x}.$
-Note that $\mathcal{L}=S^{1}$ and that each iterate $N_{p}^{n}$ can be continuously extended to the extended real line $\mathbb{R}\cup\{\infty\}$, i.e., $\mathcal{L}$. For example,
+As before, $(L,\mathcal{P},\Gamma)$ is a CCS.
+Note that $\mathcal{L}=S^{1}$ and that each iterate $N_{p}^{n}$ can be continuously extended to the extended real line $\mathbb{R}\cup\{\infty\}$, i.e., $\mathcal{L}$.
+For example,
 $\tilde{N_{p}}(x)=\begin{cases}\frac{x^{2}+a}{2x},&\text{if }x\in L;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
@@
 However, $\tilde{f}$$\mathcal{L}\rightarrow\mathcal{L}$ is not continuous.
-The set $\overline{\tilde{\Delta}}$ of deep computations is $\tilde{\Delta}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
+The set $\overline{\tilde{\Gamma}}$ of deep computations is $\tilde{\Gamma}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
 Example 2.9.
@@
 Let $B$ be the basin of the attractive cycle $0,1,0,1,\dots$.
-Let $L=B\cup\bigcup_{i=1}^{3}B_{i}$ and $J$ be its complement, i.e., the Julia set.
+Then, $L_{p}=B\cup\bigcup_{i=1}^{3}B_{i}$.
 Notice that $N_{p}^{n}$ does not converge pointwise.
@@
 $f_{1}$ and $f_{2}$ are two distinct deep computations.
-Note that for $z\in J$, no subsequence of $\tilde{N_{p}^{n}}(z)$ converges to a complex number.
+Note that for $z\notin L_{p}$, no subsequence of $\tilde{N_{p}^{n}}(z)$ converges to a complex number.
 However, since $\mathcal{L}=S^{2}$ is compact there is a subsequence of $\tilde{N_{p}^{n}}(z)$ that converges to $\infty$.
-We can extend $f_{i}:L\rightarrow\mathcal{L}$ to $\tilde{f_{i}}$$\mathcal{L}\rightarrow\mathcal{L}$ by:
-$\tilde{f_{i}}(z)=\begin{cases}f_{i}(z),&\text{if }z\in L;\\ \infty,&\text{if }z\in J.\end{cases}.$
-Again, note that $\tilde{f_{i}}$ for $i=1,2$ are not continuous and that $\tilde{f_{i}}\in\overline{\tilde{\Delta}}$.
+We can extend $f_{i}:L_{p}\rightarrow\mathcal{L}$ to $\tilde{f_{i}}$$\mathcal{L}\rightarrow\mathcal{L}$ by:
+[…]
+Again, note that $\tilde{f_{i}}$ for $i=1,2$ are not continuous and that $\tilde{f_{i}}\in\overline{\tilde{\Gamma}}$.
 2.4. Finite precision threshold classifiers as a CCS
 Let $L=2^{\mathbb{N}}$, i.e., the set consisting of all infinite binary sequences with the topology of pointwise convergence.
-Let $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ is the collection of projections, i.e., $P_{n}(x)=x(n)$ for each $n$ and each $x\in L$.
-In other words, $P_{n}$ is the predicate that reads the $n$th digit of a sequence.
+Let $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ equal the collection of projections, i.e., $P_{n}(x)=x(n)$ for each each $x\in L$ and $n\in\mathbb{N}$.
 Notice that $L\subseteq\mathbb{R}^{\mathcal{P}}$ is closed.
@@
 The set of finite binary strings is denoted by $2^{<\mathbb{N}}$.
+This set is naturally ordered by the lexicographic order $\leq_{\mathrm{lex}}$.
 Given a finite binary string $w$, we consider the transition $\phi_{w}:L\to L$ given by the rule
-$\phi_{w}(x)=\begin{cases}1^{\infty},&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0^{\infty},&\text{otherwise};\end{cases}$
+$\phi_{w}(x)=\begin{cases}1^{\infty},&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0^{\infty},&\text{otherwise},\end{cases}$
 where $|w|$ is the length of the string $w$ and $x|_{|w|}$ is the prefix of $x$ of length $|w|$.
@@
 $f(x)=\begin{cases}1,&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0,&\text{otherwise}.\end{cases}$
-We first prove that $f^{-1}(1)=\{x\in 2^{\mathbb{N}}:x|_{|w|}\leq_{\mathrm{lex}}w\}$ is an open set.
+We first observe that $f^{-1}(1)=\{x\in 2^{\mathbb{N}}:x|_{|w|}\leq_{\mathrm{lex}}w\}$ is an open set.
 Fix $x_{0}\in f^{-1}(1)$.
 Let $t:=x_{0}|_{|w|}$ and consider the basic open set $t]=\{x\in 2^{\mathbb{N}}:x|_{|t|}=t\}$.
-Then, it is not difficult to check that $x_{0}\in[t]\subseteq f^{-1}(1)$.
+Then it is not difficult to check that $x_{0}\in[t]\subseteq f^{-1}(1)$.
 The same reasoning shows that $f^{-1}(0)$ is open.
 ∎
-Let $\Delta=\{\phi_{w}:w\in 2^{<\mathbb{N}}\}\cup\{\mathbf{0^{\infty}},\mathbf{1^{\infty}}\}$ where $\mathbf{0^{\infty}},\mathbf{1^{\infty}}:L\rightarrow L$ are the constant maps identical to $0^{\infty}$ and $1^{\infty}$, respectively.
+Let $\Delta=\{\phi_{w}:w\in 2^{<\mathbb{N}}\}\cup\{\mathbf{0^{\infty}},\mathbf{1^{\infty}}\}$, where $\mathbf{0^{\infty}},\mathbf{1^{\infty}}:L\rightarrow L$ are the constant maps identical to $0^{\infty}$ and $1^{\infty}$, respectively.
 Let $\Gamma$ be the semigroup generated by $\Delta$.
@@
 The members of $\Delta$ are finite-precision approximations of classifiers that check all bits of information.
-But here it gets interesting: what is the difference, in terms of floating-point arithmetic, between $x<0.5$ and $x\leq 0.5$?
+But here it gets interesting: what is the difference, in terms of arbitrary-precision arithmetic, between $x<0.5$ and $x\leq 0.5$?
 Suppose that $f_{a}^{+}$ represents the $\leq$ classifier for a target $a\in L$.
@@
 The crucial step follows from the fact that for any $x<_{\mathrm{lex}}a$, there is a large enough $K$ such that $x<_{\mathrm{lex}}w_{k}$ for all $k\geq K$.
-The preceding proposition shows that the topological structure of deep computations can be very complicated.
+The preceding proposition shows that the topological structure of deep computations can be complicated.
 Indeed, $\overline{P_{n}\circ\Delta}$ contains the Split Cantor space for all $n\in\mathbb{N}$.
@@
 Proof.
-Fix $a\in 2^{\mathbb{N}}$.
-Let $w_{n}:=a|_{n}$ for each $n\in\mathbb{N}$.
+Fix $a\in 2^{\mathbb{N}}$, and let $w_{n}:=a|_{n}$ for each $n\in\mathbb{N}$.
 We claim that $\lim_{n\rightarrow\infty}\psi_{w_{n}}(x)=\delta_{a}(x)$ for all $x\in L$.
@@
 Proof.
-To show that $\mathbf{0^{\infty}}\in\overline{\Delta}$ consider, for each $n\in\mathbb{N}$, ${w_{n}=1^{n}}^{\frown}0$, i.e., the string consisting of $n$ consecutive $1$s followed by a 0.
+To show that $\mathbf{0^{\infty}}\in\overline{\Delta}$, for each $n\in\mathbb{N}$, consider, ${w_{n}=1^{n}}^{\frown}0$, i.e., the string consisting of $n$ consecutive $1$s followed by a 0.
 If $x=1^{\infty}$, then $x|_{|w_{n}|}\neq w_{n}$ for all $n\in\mathbb{N}$.
@@
   1. (i)
-    $\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$.
+    $\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$;
   2. (ii)
@@
 ∎
-3.2. The Todorčević trichotomy and levels of PAC learnability
+3.2. The Todorčević trichotomy, and levels of NIP and PAC learnability
 In this subsection we study the case when the set of deep computations is a separable Rosenthal compactum. We are interested in the separable case for two reasons:
@@
 By introducing Todorčević’s Trichotomy into this framework, we obtain a classification of the complexity of deep computations.
-Given a countable set $\Delta$ of computations satisfying the NIP on features and shards (condition (ii) of Theorem 3.1) we have that $\overline{\tilde{\Delta}_{\mathcal{L}[r_{\bullet}]}}$ (for a fixed sizer $r_{\bullet}$) is a separable Rosenthal compactum (see Definition 1.9).
-Todorčević proved a remarkable trichotomy for Rosenthal compacta [TOD99], and later Argyros, Dodos, Kanellopoulos [ADK08] proved an heptachotomy that refined Todorčević’s classification.
+Given a countable set $\Delta$ of computations satisfying the NIP on features and shards (condition (ii) of Theorem 3.1), the set $\overline{\tilde{\Delta}_{\mathcal{L}[r_{\bullet}]}}$ (for a fixed sizer $r_{\bullet}$) is a separable Rosenthal compactum (see Definition 1.9).
+Todorčević proved a remarkable trichotomy for Rosenthal compacta [TOD99] that was later refined through an heptachotomy proved by Argyros, Dodos, Kanellopoulos [ADK08].
 In this section, inspired by the work of Glasner and Megrelishvili [GM22], we study ways in which this classification allows us to obtain different levels of PAC-learnability and NIP.
@@
 […]
-We now present some separable and non-separable examples of Rosenthal compacta. These show that the previously discussed classes NIPi are not equal.
 Examples 3.3.
@@
 For each $a\in 2^{\mathbb{N}}$ consider the map $\delta_{a}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ given by $\delta_{a}(x)=1$ if $x=a$ and $\delta_{a}(x)=0$ otherwise.
-Let $A(2^{\mathbb{N}})=\{\delta_{a}:a\in 2^{\mathbb{N}}\}\cup\{0\}$, where 0 is the zero map.
-Notice that $A(2^{\mathbb{N}})$ is a compact subset of $B_{1}(2^{\mathbb{N}})$, in fact $\{\delta_{a}:a\in 2^{\mathbb{N}}\}$ is a discrete subspace of $B_{1}(2^{\mathbb{N}})$ and its pointwise closure is precisely $A(2^{\mathbb{N}})$.
-Hence, this is a Rosenthal compactum which is not first countable.
-Notice that this space is also not separable.
+Let $A(2^{\mathbb{N}})=\{\delta_{a}:a\in 2^{\mathbb{N}}\}\cup\{\mathbf{0}\}$, where $\mathbf{0}$ is the zero map.
+Notice that $A(2^{\mathbb{N}})$ is a compact subset of $B_{1}(2^{\mathbb{N}})$.
+In fact, $\{\delta_{a}:a\in 2^{\mathbb{N}}\}$ is an uncountable discrete subspace of $B_{1}(2^{\mathbb{N}})$, and its pointwise closure is precisely $A(2^{\mathbb{N}})$.
+Hence, this is a Rosenthal compactum which is not hereditarily separable (and therefore not first countable).
+In particular, this space is does not satisfy separability, but it can be made separable by adding a countable set as the next example shows.
   2. (2)
@@
 The Split Cantor is the space $S(2^{\mathbb{N}})=\{f_{a}^{-}:a\in 2^{\mathbb{N}}\}\cup\{f_{a}^{+}:a\in 2^{\mathbb{N}}\}$, which was obtained as the closure of the space discussed in Section 2.4, giving an example separating NIP2 from NIP3.
-This is a separable Rosenthal compactum.
+This is a well known separable Rosenthal compactum.
 One example of a countable dense subset is the set of all $f_{a}^{+}$ and $f_{a}^{-}$ where $a$ is an infinite binary sequence that is eventually constant.
 Moreover, it is hereditarily separable, but it is not metrizable.
+It is homeomorphic to the space $2^{\mathbb{N}}\times\{0,1\}$ with the lexicographic order topology via the identification $(a,1)\mapsto f_{a}^{+}$ and $(a,0)\mapsto f_{a}^{-}$.
   4. (4)
     Alexandroff Duplicate.
-Let $K$ be any compact metric space and consider the Polish space $X=C(K)\sqcup K$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$.
-For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={x(a), if x∈C(K);0, if x∈K; g_a^1(x)={x(a), if x∈C(K);δa(x), if x∈K.
+Let $K$ be any compact metric space and consider the Polish space $X=K\sqcup C(K)$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$.
+For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={0, if x∈Kx(a), if x∈C(K); g_a^1(x)={δa(x), if x∈K;x(a), if x∈C(K).
 Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$.
+Observe that all points $g_{a}^{1}$ are isolated and that open neighborhoods of $g_{a_{0}}^{0}$ are of the form $g_{a}^{i}:a\in U,i\in\{0,1\}\}\backslash\{g_{a}^{1}:a\in F\}$ where $U\subseteq K$ is an open neighborhood of $a_{0}$ and $F\subseteq K$ is a finite set.
+Another abstract way in which this space is presented is as the space $K\times\{0,1\}$ whose basic open neighborhoods are given as before, identifying $(a,0)\mapsto g_{a}^{0}$ and $(a,1)\mapsto g_{a}^{1}$.
+We can also embed $D(K)$ into the product $A(K)\times K$ by identifying $(a,0)\mapsto(\mathbf{0},a)$ and $(a,1)\mapsto(\delta_{a},a)$.
 Notice that $D(K)$ is a first countable Rosenthal compactum.
-It is not separable if $K$ is uncountable.
-The interesting case will be when $K=2^{\mathbb{N}}$.
+It is not separable if $K$ is uncountable, thus we typically study the interesting case when $K=2^{\mathbb{N}}$.
+As with the Alexandroff compactification $A(2^{\mathbb{N}})$, we can make the space $D(2^{\mathbb{N}})$ separable by adding a countable set.
+For example, the closure of the set $\{(v_{s},s^{\frown}0^{\infty}):s\in 2^{<\mathbb{N}}\}\subseteq\hat{A}(2^{\mathbb{N}})\times 2^{\mathbb{N}}$ is $\{(\mathbf{0},a):a\in 2^{\mathbb{N}}\}\cup\{(\delta_{a},a):a\in 2^{\mathbb{N}}\}\cup\{(v_{s},s^{\frown}0^{\infty}):s\in 2^{<\mathbb{N}}\}$, where $\{(\mathbf{0},a):a\in 2^{\mathbb{N}}\}\cup\{(\delta_{a},a):a\in 2^{\mathbb{N}}\}$ is homeomorphic to $D(2^{\mathbb{N}})$.
   5. (5)
@@
 Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$.
+The identification $h_{t}\mapsto(v_{t},f_{t^{\frown}0^{\infty}}^{+})$ lifts to a homeomorphism between $\hat{D}(S(2^{\mathbb{N}}))$ and the subspace of $\hat{A}(2^{\mathbb{N}})\times S(2^{\mathbb{N}})$ consisting of $(\mathbf{0},f_{a}^{+})$, $(\mathbf{0},f_{a}^{-})$, $(\delta_{a},f_{a}^{+})$ and $(v_{t},f_{t^{\frown}0^{\infty}}^{+})$ for $a\in 2^{\mathbb{N}}$ and $t\in 2^{<\mathbb{N}}$ (see 4.3.7 in [ADK08]).
 Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable.
-In fact, it contains an uncountable discrete subspace (see Theorem 5 in [TOD99]).
+In fact, it contains an uncountable discrete subspace.
 Theorem 3.4 (Todorčević’s Trichotomy, [TOD99], Theorem 3 in [ADK08]).
@@
   3. (3)
+    Let $(L,\mathcal{P},\Gamma)$ be the CCS given by $L=2^{\mathbb{N}}$, $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ and $\Gamma$ is the semigroup generated by $\Delta=\{\gamma_{t}:t\in 2^{<\mathbb{N}}\}$, where $P_{n}:2^{\mathbb{N}}\rightarrow\{0,1\}$ is the projection map $P_{n}(x)=x(n)$ and $\gamma_{t}:L\rightarrow L$ is given by γ_t(x)={0, if lext0;(01), if t0lexx ≤lext1;1, if t1¡lexx.
+where $(01)^{\infty}$ denotes the sequence of alternating bits: $010101\cdots$.
+As in the other examples, it is not difficult to see that $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom.
+For example, the condition $t^{\frown}0^{\infty}\leq_{\text{lex}}x\leq_{\text{lex}}t^{\frown}1^{\infty}$ is equivalent to $x$ extending $t$.
+Observe that the set of deep computations is homeomorphic to $\hat{D}(S(2^{\mathbb{N}}))$ (see Example 3.3(5)).
+This is an example of $\Delta$ which is NIP1 but not NIP2.
+  4. (4)
     Let $(L,\mathcal{P},\Gamma)$ be the finite precision prefix test (Section 2.5) with $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$.
-We saw that $\overline{\tilde{\Delta}}$ is homeomorphic to the Extended Alexandroff compactification $\hat{A}(2^{\mathbb{N}})$ (Example 3.3(3)), which is separable but not first countable.
+We saw that $\overline{\tilde{\Delta}}$ is homeomorphic to the Extended Alexandroff compactification $\hat{A}(2^{\mathbb{N}})$ (Example 3.3-(3)), which is separable but not first countable.
 Hence, $\Delta$ is NIP but not NIP1.
@@
     Our emphasis is computational.
-Elements of $2^{<\mathbb{N}}$ represent finite bitstrings, i.e., standard computations, while Rosenthal compacta represent deep computations, i.e., limits of finite computations.
-Mathematically, deep computations are pointwise limits of standard computations.
-However, computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
+Real numbers can be represented as infinite binary sequences, i.e., infinite branches of the binary tree $2^{<\mathbb{N}}$.
+We approximate real numbers or binary sequences with elements in $2^{<\mathbb{N}}$, i.e., finite bitstrings.
+Indexing standard computations with finite bitstrings allow us to better understand how deep computations arise and how they get approximated.
+Computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
   2. (2)
@@
   1. (1)
-    If $I$ is a countable and $\{f_{i}:i\in I\}\subseteq\mathbb{R}^{X}$, $\{g_{i}:i\in I\}\subseteq\mathbb{R}^{X}$ are two pointwise families by $I$, we say that $\{f_{i}:i\in\ I\}$ and $\{g_{i}:i\in I\}$ are equivalent if and only if the map $f_{i}\mapsto g_{i}$ is extended to a homeomorphism from $\overline{\{f_{i}:i\in I\}}$ to $\overline{\{g_{i}:i\in I\}}$.
+    If $I$ is countable and $\{f_{i}:i\in I\}\subseteq\mathbb{R}^{X}$, $\{g_{i}:i\in I\}\subseteq\mathbb{R}^{X}$ are two pointwise families by $I$, we say that $\{f_{i}:i\in\ I\}$ and $\{g_{i}:i\in I\}$ are equivalent if and only if the map $f_{i}\mapsto g_{i}$ is extended to a homeomorphism from $\overline{\{f_{i}:i\in I\}}$ to $\overline{\{g_{i}:i\in I\}}$.
   2. (2)
@@
 We follow the same notation as in [ADK08].
-For any node $t\in 2^{<\mathbb{N}}$, let us denote by $t^{\frown}0^{\infty}$ ($t^{\frown}1^{\infty}$) the infinite binary sequence starting with $t$ and continuing will all 0’s (respectively, all $1$’s).
+For any node $t\in 2^{<\mathbb{N}}$, let us denote by $t^{\frown}0^{\infty}$ ($t^{\frown}1^{\infty}$) the infinite binary sequence starting with $t$ and continuing with all 0’s (respectively, all $1$’s).
 Fix a regular dyadic subtree $R=\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$ (i.e., a dyadic subtree such that every level of $R$ is contained in a level of $2^{<\mathbb{N}}$) with the property that for all $s,s^{\prime}\in R$, $s^{\frown}0^{\infty}\neq s^{\prime\frown}0^{\infty}$ and $s^{\frown}1^{\infty}\neq s^{\prime\frown}1^{\infty}$.
@@
   3. (3)
-    $D_{3}=\{f^{+}_{s_{t}^{\frown}0^{\infty}}:t\in 2^{<\mathbb{N}}\}$. This is a discrete in $\overline{D_{3}}=S(2^{\mathbb{N}})$.
+    $D_{3}=\{f^{+}_{s_{t}^{\frown}0^{\infty}}:t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{3}}=S(2^{\mathbb{N}})$.
   4. (4)
@@
   7. (7)
-    $D_{7}=\{(v_{s_{t}},x^{+}_{s_{t}^{\frown}0^{\infty}}):t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{7}}=\hat{D}(S(2^{\mathbb{N}}))$
+    $D_{7}=\{(v_{s_{t}},x^{+}_{s_{t}^{\frown}0^{\infty}}):t\in 2^{<\mathbb{N}}\}$. This is discrete in $\overline{D_{7}}=\hat{D}(S(2^{\mathbb{N}}))$.
 Theorem 3.8 (Heptachotomy of minimal families, Theorem 2 in [ADK08]).
@@
 Moreover, all $D_{i}$ are minimal and mutually non-equivalent.
-The implications of this result for deep computations is the following: for any countable set of computations $\Delta$ satisfying the NIP (for some CCS $(L,\mathcal{P},\Gamma)$), we can always find a countable discrete set of deep computations that approximate all the other deep computations.
+The implication of this result for deep computations is the following: for any countable set of computations $\Delta$ satisfying the NIP (for some CCS $(L,\mathcal{P},\Gamma)$), we can always find a countable discrete set of deep computations that approximates all the other deep computations.
 For example: in the finite precision prefix test example (subsection 2.5), the prefix test computations (family $D_{5}$) approximate all other deep computations.
@@
 4.1. NIP and Monte Carlo computability of deep computations
-The raison d’être of the Baire class-1 functions is have a class of functions that contains the continuous functions and is closed under pointwise limits.
+The raison d’être of the Baire class-1 functions is to have with a class of functions that are obtained as equential limit points of continuous functions.
 By Fact 1.2, for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$.
-This motivates the following definition:
+Thus, for such $X$, functions in $B_{1}(X,Y)$ are not too far from being continuous.
+In this section we will study a more general class of functions, namely, the class of universally measurable functions, which we define next.
 Definition 4.1.
@@
 When $Y=\mathbb{R}$, we will always take $\Sigma=\mathcal{B}(\mathbb{R})$, the Borel $\sigma$-algebra of $\mathbb{R}$.
+If $X$ is a compact (Hausdorff) space, then every Radon measure $\mu$ on $X$ is finite.
+Then, the measure given by $\nu(A):=\mu(A)/\mu(X)$ is a probability measure on $X$ with the same null sets as $\mu$.
+Hence, Radon measures on compact spaces are equivalent to (Radon) probability measures.
+We summarize this fact in the next remark:
 Remark 4.2.
-A function $f:X\rightarrow\mathbb{R}$ is universally measurable if and only if $f^{-1}(U)$ is $\mu$-measurable for every Radon probability measure $\mu$ on $X$ and every open set $U\subseteq\mathbb{R}$.
+If $X$ is compact, then a function $f:X\rightarrow\mathbb{R}$ is universally measurable if and only if $f^{-1}(U)$ is $\mu$-measurable for every Radon probability measure $\mu$ on $X$ and every open set $U\subseteq\mathbb{R}$.
 Intuitively, a function is universally measurable if it is “measurable no matter which reasonable way you try to measure things on its domain”.
@@
 In the context of deep computations, we are interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself.
-There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the set $\pi_{i}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open).
+There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the sets $\pi_{P}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open and $P\in\mathcal{P}$).
 Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller.
@@
   3. (iii)
-    For every Radon measure $\mu$ on $X$, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$, i.e., every countable subset of $A$ has an accumulation point in $\mathscr{M}^{0}(X,\mu)$.
+    For every Radon measure $\mu$ on $X$, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$, i.e., every countable subset of $A$ has a limit point in $\mathscr{M}^{0}(X,\mu)$.
 This result allows us to formalize the concept of a deep computation being Monte Carlo computable, which we define below (Definition 4.6). To motivate the definition, let us first recall two facts:
@@
     Littlewoood’s second principle states that every Lebesgue measurable function is “nearly continuous”.
-The formal statement of this, which is Luzin’s theorem, is that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ is a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable ) equipped with a Borel algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|_{F}$ is continuous and $\mu(E\backslash F)<\varepsilon$.
+The formal statement of this, which is Luzin’s theorem, is that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ is a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable) equipped with the Borel $\sigma$-algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|_{F}$ is continuous and $\mu(E\backslash F)<\varepsilon$.
   2. (2)
@@
 Remark 4.7.
-Condition (2) of Theorem  shows that to study measure-theoretic versions of NIP, we need only consider compact subsets of $X$.
+Condition (2) of Theorem 4.5 shows that to study measure-theoretic versions of NIP, we need only consider compact subsets of $X$.
 Now, every Radon measure on a compact space is finite; hence, by proper normalization, it can be treated as a probability measure.
@@
 Let $\mu$ be a Radon probability measure on $X$.
-Given a $\mu$-measurable set $E\subseteq X$, a positive integer $k$ and real numbers $a<b$.
-we write:
+Given a $\mu$-measurable set $E\subseteq X$, a positive integer $k$ and real numbers $a<b$, we write:
 $D_{k}(A,E,a,b)=\bigcup_{f\in A}\{x\in E^{2k}:f(x_{2i})\leq a,\hskip 2.84526ptf(x_{2i+1})\geq b\hskip 2.84526pt{\text{ for all }i<k}\}.$
-We say that $A$ is Talagrand $\mu$-stable if and only if for every $\mu$-measurable set $E\subseteq X$ of positive measure and for every $a<b$ there is $k\geq 1$ such that
+We say that $A$ is Talagrand $\mu$-stable if and only if for every $\mu$-measurable set $E\subseteq X$ of positive measure and for every $a<b$ there is a $k\geq 1$ such that
 $(\mu^{2k})^{*}(D_{k}(A,E,a,b))<(\mu(E))^{2k},$
-where $\mu^{*}$ denotes the outer measure (we need to work with outer measure since the sets $D_{k}(A,E,a,b)$ need not be $\mu$-measurable).
+where $\mu^{*}$ denotes the outer measure (we need to work with outer measure since the sets $D_{k}(A,E,a,b)$ need not be $\mu^{2k}$-measurable).
 The inequality certainly holds when $A$ is a countable set of continuous (or $\mu$-measurable) functions.
-The main result of this section is that accumulation points of standard computations (i.e., deep computations) are Monte Carlo computable; this is Theorem 4.14 below.
-We now prove that deep computations that are Talagrand $\mu$-stable are $\mu$-measurable.
+The main result of this section is that deep computations, i.e., limit points of a Talagrand stable set of computations are Monte Carlo computable; this is Theorem 4.14 below.
+We now prove that limit points of a Talagrand $\mu$-stable set are $\mu$-measurable.
 But first, let us state the following useful characterization of measurable functions (compare with Fact 1.2):
@@
     $(X,\Sigma,\mu)$ is inner regular with respect to $\mathcal{K}$, i.e., for all $E\in\Sigma$ μ(E)=sup{μ(K):$K$ and KE}.
-(In particular, if $X$ is Hausdorff, $\mu$ is a Radon measure on $X$ and $\mathcal{K}$ is the collection of compact subsets of $X$, all these conditions hold).
+(In particular, if $X$ is compact Hausdorff, $\mu$ is a Radon probability measure on $X$, $\Sigma$ is the completion of the Borel $\sigma$-algebra by $\mu$, and $\mathcal{K}$ is the collection of compact subsets of $X$, all these conditions hold).
 Then, $f:X\rightarrow\mathbb{R}$ is measurable if and only if for every $K\in\mathcal{K}$ with $0<\mu(K)<\infty$ and $a<b$, either $\mu^{*}(P)<\mu(K)$ or $\mu^{*}(Q)<\mu(K)$ where $P=\{x\in K:f(x)\leq a\}$ and $Q=\{x\in K:f(x)\geq b\}$.
-The following technical lemma will be instrumental to prove of Proposition 4.13, which, in turn, will yield the main result of the subsection, Theorem 4.14.
+The following technical lemma will be instrumental for proving Proposition 4.13, which, in turn, will yield the main result of the subsection, namely Theorem 4.14.
 Lemma 4.11.
@@
 We say that $A$ is universally Talagrand stable if $A$ is Talagrand $\mu$-stable for every Radon probability measure $\mu$ on $X$.
-We first observe that universal Talagrand stability corresponds to a complexity class smaller than or equal to the NIP:
+We first observe that universal Talagrand stability corresponds to a complexity class smaller than or equal to the NIP class:
 Proposition 4.13.
@@
 (The assumption that $[0,1]$ is not the union of $\mathfrak{c}$ closed measure zero sets is a consequence of, for example, the Continuum Hypothesis.)
-On the other hand, by fixing a particular well-known measure, the Lebesgue measure, we see that the other case is also consistent:
+On the other hand, by fixing a particular well-known measure, namely the Lebesgue measure, we see that the other case is also consistent:
 Theorem 4.16 (Fremlin, Shelah, [FS93]).
@@
 Notice that the preceding two results apply to sets of measurable functions, a class of functions larger than the class of continuous functions.
-It is possible that the continuity assumption, i.e., the Extendibility Axiom, settles the problem.
-This is a problem that remains open for future work.
+However, by the Extendibility Axiom, finitary computations are continuous, i.e., if $A$ is a set of computations, then $A\subseteq C_{p}(X)$.
+The question of whether we can remove the set-theoretic assumption in Theorem 4.15 when $A\subseteq C_{p}(X)$ (instead of $A\subseteq M_{r}(X)$) remains open.

### v4

@@
 Abstract.
-We use topological methods to study the complexity of deep computations and limit computations.
-We use topology of function spaces, specifically, the classification of Rosenthal compacta, to identify new complexity classes.
-We use the language of model theory, specifically, the concept of independence from Shelah’s classification theory, to translate between topology and computation.
-We use the theory of Rosenthal compacta to characterize approximability of deep computations, both deterministically and probabilistically.
+We use topological methods and a suitable framework to study the complexity of limit computations (such as so-called “deep computations” implied by recurrent neural networks).
+By using topology of function spaces —specifically, the classification of Rosenthal compacta— we identify new computational complexity classes.
+The language of model theory, particularly, the concept of independence from Shelah’s classification theory, provides a medium to translate between topology and computation.
+The theory of Rosenthal compacta allows us to characterize approximability of deep computations, both deterministically and probabilistically.
 Key words and phrases:
@@
 The formal framework introduced here provides a unified setting to study these limit phenomena from a foundational viewpoint.
-Informed by model theory, to each computation in a given computation model, we associate a continuous real-valued function, called the type of the computation, that describes the logical properties of this computation with respect to the rest of the model.
-This allows us to view computations in any given computational model as elements of a space of real-valued functions, which is called the space of types of the model.
-The idea of embedding models of theories into their type spaces is central in model theory.
-In the context of this paper, the embedding of computations into spaces of types allows us to utilize the vast theory of topology of function spaces, known as $C_{p}$-theory, to obtain results about complexity of topological limits of computations.
+The main result of the paper is the identification of a hierarchy of four distinct complexity classes and seven minimal prototypes for limit computations.
+This is obtained by invoking results from topology of spaces of functions, namely, Todorčević’s trichotomy for Rosenthal compacta [TOD99] and the Argyros-Dodos-Kanellopoulos heptachotomy [ADK08], and transferring them into the realm of computation.
+The medium that we use to translate between topology and computation is model theory, specifically, Shelah’s classification theory [SHE90].
+The translation between topology and computation proceeds as follows.
+To every computation in a given computation model, we associate a continuous real-valued function, called the type of the computation, that describes the logical properties of this computation with respect to the rest of the model.
+This association allows us to view computations in any given computational model as elements of a space of real-valued functions, which is called the space of types of the model.
+The idea of embedding models of theories into their type spaces is borrowed from model theory, where types spaces play a central role.
+The embedding of computations into spaces of types then allows us to utilize the vast theory of topology of function spaces, known as $C_{p}$-theory, to obtain results about complexity of topological limits of computations.
 As we shall indicate next, recent classification results for spaces of functions provide an elegant and powerful machinery to classify computations according to their levels of “tameness” or “wildness”, with the former corresponding roughly to polynomial approximability and the latter to exponential approximability.
-The viewpoint of spaces of types, which we have borrowed from model theory, thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: In topology, the classification of Rosenthal compacta pioneered by Todorčević [TOD99]; in logic, the classification of theories developed by Shelah [SHE90]; and in statistical learning, the notions of PAC learning and VC dimension pioneered by Vapnik and Chervonenkis [VC74, VC71].
-In a previous paper [ADI+24], the authors introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]).
-There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
+The model-theoretic viewpoint of spaces of types thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: in topology, the classification of Rosenthal compacta pioneered by Todorčević [TOD99]; in logic, the classification of theories developed by Shelah [SHE90]; and in statistical learning, the notions of PAC learning and VC dimension pioneered by Vapnik and Chervonenkis [VC74, VC71].
+In a previous paper [ADI+24], the authors introduced the concept of limits of computations, which they called ultracomputations (given that they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]).
+There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use primarily the term “deep computation”.
 In [ADI+24], a new “tame vs wild” (i.e., polynomially approximable vs nonpolynomially approximable) dichotomy for the complexity of deep computations was proved by invoking a classical result of Grothendieck from the 1950s [GRO52].
 Under our model-theoretic Rosetta stone, the property of polynomial approximability of computations is identified with continuous extendibility in the sense of topology, and with the notions of stability and type definability in the sense of model theory.
+In this paper, we refine the dichotomy of [ADI+24], and obtain new complexity classes for deep computation.
+The “wild” part of the dichotomy from [ADI+24] is further subdivided between “tame” and “wild” classes; three new levels of tameness are identified within this new subdivision, and seven distinct minimal prototypes of tameness are detected.
+The finer classification results of this paper are obtained through more contemporary topological and model-theoretic tools than in the previous paper [ADI+24].
+From the topological side, instead of Grothendieck’s classical criterion, we use the modern theory of Rosenthal compacta, and from the model theory side, instead of stability, we use neostability.
+The main tool imported from neostability is Shelah’s Independence Property, or rather, its negation, which is commonly referred to as “NIP”.
+More on this below.
 Deep computations arise as topological limits of standard (continuous) computations.
-Simplest among these is when the limit is sequential.
-In topology, the first Baire class, or Baire class 1 consists of functions (also called simply “Baire-1”) arising as pointwise limits of sequences of continuous functions.
-Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity above the Baire class 0 which (by definition) consists of continuous functions.
-In this paper, we apply the theory of spaces of Baire class 1 functions to identify and separate new complexity classes of deep computations.
-We obtain a new “tame vs wild” dichotomy for deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT78], and further, a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [TOD99].
-Todorčević’s trichotomy concerns Rosenthal compacta, which are topological spaces that fit our context superbly: they are defined as compact spaces that are homeomorphic to subspaces of the space of Baire class 1 functions on some Polish (separable, completely metrizable) space, under the topology of pointwise convergence.
+In topology, the first Baire class, or Baire class 1 consists of functions (also called simply Baire-1) arising as pointwise limits of sequences of continuous functions.
+Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity above the Baire class 0, which (by definition) consists of continuous functions.
+We apply a result of Bourgain, Fremlin and Talagrand from the late 1970s on compact subsets of the first Baire class to obtain a new “tame vs wild” dichotomy for deep computations [BFT78]; then, we apply a trichotomy of Todorčević, from the late 90s, for Rosenthal compacta to identify new tame classes [TOD99].
+A Rosenthal compactum is a compact topological space that is homeomorphic to a subspace of the space of Baire class 1 functions on some Polish (i.e., separable, completely metrizable) space equipped with the topology of pointwise convergence.
 Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways.
@@
 It is noteworthy that Todorčević’s proof relies on sophisticated set-theoretic forcing and infinite Ramsey theory.
-At the time of writing this paper, decades after his original argument, no elementary proof has been found [TOD23, HT19].
-Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of Non Independence Property (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concepts VC-dimension and Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory [VC71, VC74, VAL84].
-Refining Todorčević’s trichotomy, we invoke a more recent heptachotomy for separable Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]; these authors identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
+At the time of writing this paper, decades after his original argument, there are no elementary proofs of the trichotomy; in fact, Todorčević’s methods have been analyzed and exploited further [TOD23, HT19].
+Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of No-Independence Property (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concepts VC-dimension and Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory [VC71, VC74, VAL84].
+Beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for separable Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08].
+These authors identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
+From the point of view of approximation theory, it is noteworthy that in the context of Grothendieck [GRO52] and Bourgain-Fremlin-Talagrand [BFT78], compactness and sequential compactness coincide; thus, the results of these papers allow us to ensure that deep computations, which in principle are abstract ultrafilter limits (over a possibly uncountable set), can always be approximated by concrete sequences of standard computations.
 We believe that the results presented in this paper show practitioners of computation, or topology, or descriptive set theory, or model theory, how classification invariants used in their field translate into classification invariants of other fields.
@@
 The necessary topological background beyond undergraduate topology is covered in section 1.
-In section 1, we present basic topological and combinatorial preliminaries, and in section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed).
+In section 1, we present basic topological and combinatorial preliminaries.
+In section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed).
 Section 3 is devoted to the classification of deep computations.
 Throughout the paper, our results pertain to classical models of computation (particularly computations involving real-valued quantities that are known and manipulated to a finite degree of precision).
-The final section, Section 4, introduces a probabilistic viewpoint, the development of which we intend to pursue in future research, extending the present framework to encompass non-deterministic and quantum computations.
-1. General topological preliminaries: From continuity to Baire class 1
-In this section we present some preliminaries from general topology and function space theory.
-In the interest of completeness, we include some proofs that may be safely skipped by readers familiar with these topics.
-Recall that a subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets.
-A space is metrizable if its topology agrees with the topology induced by some metric therein.
-Two such metrics inducing the same topology may induce quite different properties in the category of metric spaces.
-For example, the interval $(0,1)$ with the usual metric (as a subset) of the reals is not complete; however, $(0,1)$ is homeomorphic to the real line, which is complete with respect to the usual metric thereon.
-In a metrizable space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
-A Polish space is a separable and completely metrizable topological space, i.e., admitting some complete metric inducing its topology.
+The final section, Section 4, introduces a probabilistic viewpoint, the development of which we intend to extend in future research.
+1. Topological preliminaries:
+From continuity to Baire class 1
+In this section we recall notions and results from general topology and function space theory. For simplicity, we assume that all spaces are Hausdorff unless explicitly stated otherwise.
+A subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets.
+In a metric space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
+A Polish space is a separable and completely metrizable topological space, i.e., one admitting a complete metric inducing its topology.
 Although other (possibly incomplete) metrics may induce the same topology, being Polish is a purely topological property.
-One of the most important Polish spaces is the real line $\mathbb{R}$; the others include the Cantor space $2^{\mathbb{N}}$ and the Baire space $\mathbb{N}^{\mathbb{N}}$.
-The class of Polish spaces is closed under countable topological products; in particular, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology), and the space $\mathbb{R}^{\mathbb{N}}$ of sequences of real numbers are all Polish.
-Recall that the product topology on these spaces is the topology of pointwise convergence: a sequence converges in the space if and only if it converges at each coordinate index.
+The class of Polish spaces includes the real line $\mathbb{R}$ as an important example, and is closed under countable topological products.
+[^1] Recall that the product topology is the topology of pointwise convergence.
+In a product space, a sequence (or a net, or a filter) converges if and only each coordinate thereof converges.
+In particular, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology), and the space $\mathbb{R}^{\mathbb{N}}$ of sequences of real numbers are Polish spaces, each of considerable interest.
 Fact 1.1.
-A subset of a Polish space is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set. In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
-For a proof, see [ENG89, 4.3.24].
-Given two topological spaces $X$ and $Y$ we denote by $C_{p}(X,Y)$ the set of all continuous functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence.
-The space $C_{p}(X,\mathbb{R})$ of continuous real functions on $X$ is denoted simply $C_{p}(X)$.
+[ENG89, 4.3.24] A subset of a Polish space is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set.
+In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
+A subset $A$ of a topological space is said to be of the first category if $A$ is included in a countable union of closed nowhere dense sets; otherwise, $A$ is of the second category.
+A first-class subset $A\subseteq X$ is also called a mager set; its complement $X\setminus A$ is comeager.
+A topological space $X$ is Baire if every comeager subset thereof is dense.
+In Baire spaces, a meager subset is “everywhere small” in the sense that it includes no nonempty open set.
+The Baire Category Theorem states that every compact Hausdorff or completely metrizable space (hence, every Polish space) is Baire.
+[^2] More generally, any Čech-complete space is Baire [ENG89, Theorem 3.9.3].
+(All compact Hausdorff and all complete metric spaces are Čech-complete.)
+Given two topological spaces $X$ and $Y$ we denote by $C_{p}(X,Y)$ the set of all continuous functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence, i.e., with the relative (subspace) product topology $C_{p}(X,Y)\subseteq Y^{X}$—where $Y$ carries its topology, but $X$ serves merely as a (discrete) index set stripped of any topology.
+The space $C_{p}(X,\mathbb{R})$ of continuous real-valued functions on $X$ is denoted simply $C_{p}(X)$.
 A natural question is, how do topological properties of $X$ translate into $C_{p}(X)$ and vice versa?
-This general question, and the study of these spaces in general, is the concern of $C_{p}$-theory, an active field of research in general topology which was pioneered by A.
-V.
-Arhangel’skiĭ and his students in the 1970’s and 1980’s [ARK92].
-This field has found many applications in model theory and functional analysis.
-For a recent survey, see [TKA11].
-A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions.
-In symbols, $f:X\to Y$ is Baire class 1 if there is a sequence of continuous functions $f_{n}:X\to Y$ such that for all $x\in X$, $f(x)=\lim_{n\to\infty}f_{n}(x)$.
-If $X$ and $Y$ are topological spaces, the space of Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$ (as above, $B_{1}(X,\mathbb{R})$ is denoted $B_{1}(X)$).
-Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)\subseteq Y^{X}$ and we give these the topology (called the topology of pointwise convergence) inherited from the product topology of $Y^{X}$.
-The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles.
-His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
-An elementary fact about Baire class 1 functions is that they are continuous except on a set of first category (also called a meager set, a set of first cateogory is the countable union of sets whose closure has empty interior; intuitively, these sets are “topologically small”).
-Thus, Baire class 1 functions are continuous on a “topologically large” subset of their domain.
-A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ (equivalently, every open subset of $X$ is a $F_{\sigma}$).
+This general question, and the study of these spaces in general, is the concern of $C_{p}$-theory.
+A function $f\in Y^{X}$ between topological spaces $X$, $Y$ is said to be of Baire class 1, or a Baire-1 function, if there is a sequence $(f_{n})_{n\in\mathbb{N}}\subseteq C_{p}(X,Y)$ such that $f=\lim_{n\to\infty}f_{n}$ (i.e., $f(x)=\lim_{n\to\infty}f_{n}(x)$ for all $x\in X$).
+An elementary argument shows that Baire-1 functions are continuous everywhere except on a set of the first category.
+If $X$ and $Y$ are topological spaces, the space of Baire-1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$; thus, $C_{p}(X,Y)\subseteq B_{1}(X,Y)\subseteq Y^{X}$ are topological inclusions.
+We abbreviate $B_{1}(X,\mathbb{R})$ as $B_{1}(X)$.
+Relatively compact subsets of $C_{p}(X)$ and of $B_{1}(X)$ (particularly for $X$ Polish) are of interest in analysis and topological dynamics.
+A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ set (equivalently, every open subset of $X$ is an $F_{\sigma}$ set).
 Every metrizable space (hence, every Polish space) is perfectly normal.
-A topological space $X$ is Baire if every countable intersection of dense open sets is dense.
-The Baire Category Theorem states that every compact Hausdorff or completely metrizable space (hence, every Polish space) is Baire.
 The following fact was established by Baire in his groundbreaking thesis. A proof can be found in [TOD97, Section 10].
@@
 $D_{0}\subseteq f^{-1}(-\infty,a],\quad D_{1}\subseteq f^{-1}[b,\infty),\quad\overline{D_{0}}=\overline{D_{1}}.$
-A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact.
-Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) are of interest in analysis and topological dynamics.
-We begin with the following well-known result.
-Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}\geq 0$ (a pointwise bound at $x$) such that $|f(x)|\leq M_{x}$ for all $f\in A$.
-We include a proof for the reader’s convenience:
+The Lemma below is well known, but we include a proof for the reader’s convenience.
 Lemma 1.3.
-Let $X$ be a Polish space and $A\subseteq B_{1}(X)$ be pointwise bounded. The following are equivalent:
+Let $X$ be a Polish space and $A\subseteq B_{1}(X)$ be pointwise bounded.
+[^3] A set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there exists $M_{x}\geq 0$ (a pointwise bound at $x$) such that $|f(x)|\leq M_{x}$ for all $f\in A$.
+The following are equivalent:
   1. (i)
-    AA is relatively compact in $B_{1}(X)$.
+    AA is relatively compact in $B_{1}(X)$, i.e., the closure of $A$ in $B_{1}(X)$ is compact;
   2. (ii)
-    AA is relatively countably compact in $B_{1}(X)$, i.e., every countable subset of $A$ has a limit point in $B_{1}(X)$.
+    AA is relatively countably compact in $B_{1}(X)$, i.e., every (infinite) countable subset of $A$ has a limit point in $B_{1}(X)$;
   3. (iii)
-    $\overline{A}\subseteq B_{1}(X)$, where $\overline{A}$ denotes the closure in $\mathbb{R}^{X}$.
+    $\overline{A}\subseteq B_{1}(X)$, where $\overline{A}$ denotes the closure of $A$ in $\mathbb{R}^{X}$.
 Proof.
@@
 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
-In metrizable spaces, points of the closure of some subset can always be approximated by points in the set proper, via a convergent sequence.
-For more complicated spaces, such as $C_{p}$-spaces, this fails in remarkable ways.
-The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$, and one can show (e.g., [TOD97, Chapter 1.1]) that $\{p_{n}\}_{n\in\mathbb{N}}$ has no convergent subsequences, in $\mathbb{R}^{X}$.
-In a sense, this example exhibits the worst failure of sequential convergence possible.
-The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) is homeomorphic to the Stone-Čech compactification of the discrete space of natural numbers, usually denoted $\beta\mathbb{N}$, which is an important object of study in general topology.
+In metrizable spaces, the closure of a subset consists of limits of convergent sequences therein.
+In general spaces, including many $C_{p}$-spaces, such a characterization of closures fails in remarkable ways.
+The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$; however, $(p_{n})_{n\in\mathbb{N}}$ has no convergent subsequences in $\mathbb{R}^{X}$ ([TOD97, Chapter 1.1]).
+In a sense, this example exhibits the worst failure of sequential convergence possible: The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) is homeomorphic to the Stone-Čech compactification $\beta\mathbb{N}$ of the (discrete) space of natural numbers.
+The space $\beta\mathbb{N}$ has cardinality $2^{2^{\omega}}$ (largest among all separable topological spaces) and, effectively, is the set of ultrafilters on $\mathbb{N}$, which possesses complex and fascinating topological and combinatorial properties.
+[^4] Our earlier proof [ADI+24] of the existence of deep equilibria in compositional computational structures relies on a classical result on the existence of idempotent elements of $\beta\mathbb{N}$.
 The following theorem, proved by Haskell Rosenthal in 1974, is fundamental in functional analysis and captures a sharp division in the behavior of sequences in a Banach space.
 Theorem 1.4 (Rosenthal’s Dichotomy, [ROS74]).
-If $X$ is Polish and $\{f_{n}\}\subseteq C_{p}(X)$ is pointwise bounded, then $\{f_{n}\}_{n\in\mathbb{N}}$ has a convergent subsequence, or a subsequence whose closure in $\mathbb{R}^{X}$ is homeomorphic to $\beta\mathbb{N}$.
-Rosenthal’s Dichotomy states that a pointwise bounded set of continuous functions contains either a convergent subsequence, or a subsequence whose closure is essentially the same as the example mentioned in the previous paragraphs (i.e., “wildest” possible).
-The genesis of this theorem was Rosenthal’s $\ell_{1}$-Theorem”, which states that a Banach space includes an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences), or else every bounded sequence therein is weakly Cauchy.
+If $X$ is Polish and $(f_{n})_{n\in\mathbb{N}}\subseteq C_{p}(X)$ is pointwise bounded, then $(f_{n})$ has a convergent subsequence, or a subsequence whose closure in $\mathbb{R}^{X}$ is homeomorphic to $\beta\mathbb{N}$.
+The genesis of the dichotomy was Rosenthal’s $\ell_{1}$-Theorem”, which states that a Banach space includes an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences), or else every bounded sequence therein is weakly Cauchy.
 The $\ell_{1}$-Theorem connects diverse areas: Banach space geometry, Ramsey theory, set theory, and topology of function spaces.
-As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, a dichotomy paralleling the $\ell_{1}$-Theorem holds: Either every point of the closure of a set of functions is a Baire class 1 function, or there is a sequence in the set behaving in the wildest possible way.
-This result is usually not phrased as a dichotomy, but rather as an equivalence as in Theorem 1.5 below.
-First, we introduce some useful notation. For any set $A\subseteq\mathbb{R}^{X}$ and any real $a$, define
+As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, a dichotomy paralleling the $\ell_{1}$-Theorem holds: Either every point in the closure of a set of functions is a Baire class 1 function, or there is a sequence in the set behaving in the wildest possible way.
+This result is usually not phrased as a dichotomy in the above manner, but rather as an equivalence (see Theorem 1.5 below).
+First, we introduce some useful notation. For any $f\in\mathbb{R}^{X}$ and any real $a$, define
+$X^{f}_{\leq a}$ $\coloneqq f^{-1}(-\infty,a],$ $X^{f}_{\geq a}$ $\coloneqq f^{-1}[a,+\infty).$
+For any set $A\subseteq\mathbb{R}^{X}$ and any real $a$, define (by a slight abuse of notation)
 […]
 […]
-(In case $A=\emptyset$, we define $X^{\emptyset}_{\geq a}=X=X^{\emptyset}_{\leq a}$.) For any sequence $\{f_{n}\}\subseteq\mathbb{R}^{X}$ and $I\subseteq\mathbf{N}$, define $I^{\complement}\coloneqq\mathbb{N}\setminus I$ and $f_{I}\coloneqq\{f_{i}:i\in I\}$.
+(In case $A=\emptyset$, we define $X^{\emptyset}_{\geq a}=X=X^{\emptyset}_{\leq a}$.) For any sequence $\{f_{n}\}\subseteq\mathbb{R}^{X}$ and $I\subseteq\mathbb{N}$, define $I^{\complement}\coloneqq\mathbb{N}\setminus I$ and $f_{I}\coloneqq\{f_{i}:i\in I\}$.
 Theorem 1.5 (“The BFT Dichotomy”. Bourgain-Fremlin-Talagrand [BFT78]).
@@
   2. (ii)
-    For every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{I}}\cap X_{\geq b}^{f_{I^{\complement}}}=\emptyset$.
+    For every $(f_{n})_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that
+    $X_{\leq a}^{f_{I}}\cap X_{\geq b}^{f_{I^{\complement}}}=\emptyset.$
 […]
@@
 Definition 1.6.
-We say that a function collection $A\subseteq\mathbb{R}^{X}$ has the finitary No-Independence Property (NIP) if, for all sequences $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$, there exist finite disjoint sets $E,F\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}=\emptyset$.
-We say that such $E,F$ witness finitary NIP for $A$, $\{f_{n}\}$ and $a,b$.
-A set $A\subseteq\mathbb{R}^{X}$ has the finitary Independence Property (IP) if it does not have finitary NIP, i.e., if there exists a sequence $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$ such that for every pair of finite disjoint sets $E,F\subseteq\mathbb{N}$, we have $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}\neq\emptyset$.
+We say that a collection of functions $A\subseteq\mathbb{R}^{X}$ has the finitary No-Independence Property (NIP) if, for all sequences $(f_{n})_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$, there exist finite disjoint sets $E,F\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}=\emptyset$.
+We say that such $E,F$ witness finitary NIP for $(f_{n})$ and $a,b$.
+A set $A\subseteq\mathbb{R}^{X}$ has the finitary Independence Property (IP) if it does not have finitary NIP, i.e., if there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$ such that for every pair of finite disjoint sets $E,F\subseteq\mathbb{N}$, we have $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}\neq\emptyset$.
 If the word “finite” is omitted in the above definitions, we obtain the definitions of countable NIP (weaker than finitary NIP) and countable IP (stronger than finitary IP), respectively.
-If we insist on witnesses $E,F\subseteq\mathbb{N}$ such that $F=E^{\complement}$, we call the respective properties “BFT-NIP” (even weaker than countable NIP) and “BFT-IP” (even stronger than countable IP).
-Thus, Theorem 1.5 becomes that statement, for pointwise bounded function collections $A\subseteq C_{p}(X)$, that $A$ is relatively compact in $B_{1}(X)$ if and only if $A$ has BFT-NIP.
+If we insist on witnesses $E,F\subseteq\mathbb{N}$ such that $F=E^{\complement}$, we call the respective properties “BFT-NIP” and “BFT-IP”.
+Thus, Theorem 1.5 becomes (for pointwise bounded function collections $A\subseteq C_{p}(X)$) that $A$ is relatively compact in $B_{1}(X)$ if and only if $A$ has BFT-NIP.
 Unless otherwise unspecified, IP/NIP shall mean finitary IP/NIP henceforth.
@@
 Trivially (as per the preceding discussion), finitary NIP implies BFT-NIP.
-Reciprocally, assume that $X$ is compact and has finitary IP.
-Fix $A\subseteq C_{p}(X)$, a sequence $\{f_{n}\}\subseteq A$ and reals $r<s$.
-For any $I,J\subseteq\mathbb{N}$ (almost disjoint in applications), write $X_{I,J}$ for $X^{f_{I}}_{\leq r}\cap X^{f_{J}}_{\geq s}$.
-For $I\subseteq I^{\prime}\subseteq\mathbb{N}$ and $J\subseteq J^{\prime}\subseteq\mathbb{N}$, we have $X_{I,J}\supseteq X_{I^{\prime},J^{\prime}}$; moreover, $X_{I,J}=\bigcap_{E\subseteq I,F\subseteq J}X_{E,F}$, where the index variables $E\subseteq I$, $F\subseteq J$ range over finite subsets of $I,J$, respectively.
-Clearly, $E,F\subseteq\mathbb{N}$ witness finitary NIP for $\{f_{n}\}$ if and only if $X^{E,F}=\emptyset$.
+Reciprocally, assume that $X$ is compact and $A\subseteq C_{p}(X)$ has finitary IP.
+Fix a sequence $(f_{n})\subseteq A$ and reals $a<b$ such that $X^{f_{E}}_{\leq a}\cap X^{f_{F}}_{\geq b}\neq\emptyset$ for every finite disjoint $E,F\subseteq\mathbb{N}$.
 Fix $I\subseteq\mathbb{N}$.
-Since $\{f_{n}\}\subseteq A\subseteq C_{p}(X)$ is a sequence of continuous functions, and $X$ is compact, the nested family $\{X_{E,F}:E\subseteq I,F\subseteq I^{\complement}\}$ consists of closed, thus compact, sets.
-Since $A$ has finitary IP by hypothesis, the nested family consists of nonempty sets, hence its intersection $X_{I,I^{\complement}}\neq\emptyset$ by compactness.
-This holds for arbitrary $\{f_{n}\}\subseteq A$ and $r<s$, so $A$ has BFT-IP.
+Since $(f_{n})\subseteq A\subseteq C_{p}(X)$ is a sequence of continuous functions, the collection $\mathcal{F}_{I}=\{X^{f_{n}}_{\leq a}:n\in I\}\cup\{X^{f_{m}}_{\geq a}:m\in I^{\complement}\}$ consists of closed (non-empty) sets.
+Since all finite $E\subseteq I$ and $F\subseteq I^{\complement}$ witness finitary IP for $(f_{n})$ and $a,b$, we see that the collection $\mathcal{F}_{I}$ has the finite intersection property.
+By compactness, $X^{f_{I}}_{\leq a}\cap X^{f_{I^{\complement}}}_{\geq b}=\bigcap\mathcal{F}_{I}\neq\emptyset$.
 ∎
@@
 Corollary of Theorem 1.5 and Proposition 1.7. ∎
-Theorem 1.8 may be stated as the following dichotomy (under the assumptions): either $A$ is relatively compact in $B_{1}(X)$, or $A$ has IP (in either sense).
+Theorem 1.8 may be stated as the following dichotomy (under the assumptions) for any compact metrizable $X$: either $A$ is relatively compact in $B_{1}(X)$, or $A$ has IP (in either sense).
 The Independence Property was first isolated by Saharon Shelah in model theory as a dividing line between theories whose models are “tame” (corresponding to NIP) and theories whose models are “wild” (corresponding to IP).
@@
 In combinatorics:
-Sauer [SAU72] and Shelah [SHE72] proved the following independently: Let $\mathscr{F}$ be a family of subsets of some set $S$.
-Either: for every $n\in\mathbb{N}$ there is a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A:i\in\mathbb{N}\}|=2^{n}$ ($\mathscr{F}$ has “exponential complexity”); or: there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
+Sauer [SAU72] and Shelah [SHE72] proved the following independently: Let $\mathscr{F}$ be a collection of subsets of some set $S$.
+Either: for every $n\in\mathbb{N}$ there is a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A:i\in\mathbb{N}\}|=2^{n}$ ($\mathscr{F}$ is a collection having “exponential complexity”); or: there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
 $|\{S_{i}\cap A:i\in\mathbb{N}\}|\leq\sum_{i=0}^{N-1}\binom{|A|}{i}$.
-($\mathscr{F}$ has “polynomial complexity”).
-Clearly, any family $\mathscr{F}$ of subsets of a finite set $S$ has polynomial complexity.
+($\mathscr{F}$ is a collection having “polynomial complexity”).
+Clearly, any collection $\mathscr{F}$ of subsets of a finite set $S$ has polynomial complexity.
 The “polynomial” name is justified: indeed, for fixed $N>0$, as a function of the size $|A|=m>0$, we have
@@
 In this paper, we use them to study deep computations.
-For this, we shall first focus on countable languages, which is the theme of the next subsection.
 1.4. The special case $B_{1}(X,\mathbb{R}^{\mathcal{P}})$ with $\mathcal{P}$ countable.
 Fix an arbitrary (at most) countable set $\mathcal{P}$ whose elements $P\in\mathcal{P}$ will be called predicate symbols or formal predicates.
-Our present goal is to characterize relatively compact subsets of $B_{1}(X,\mathbb{R}^{\mathcal{P}})$, where $X$ is always assumed to be a perfectly normal space (often a Polish space).
+Our present goal is to characterize relatively compact subsets of $B_{1}(X,\mathbb{R}^{\mathcal{P}})$, where $X$ is always assumed to be a perfectly normal space (typically a Polish space).
 The set $\mathcal{P}$ shall be considered discrete whenever regarded as a topological space.
@@
 For $x\in X$, define the “left projection” map
-$\lambda_{x}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{\mathcal{P}}:\varphi\mapsto\lambda_{x}(\varphi)\coloneqq\varphi(x,\cdot);$
+$\lambda_{x}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{\mathcal{P}}:\varphi\mapsto\lambda_{x}(\varphi)\coloneqq\hat{}\varphi(x);$
 for $P\in\mathcal{P}$, the “right projection” map
-$\rho_{P}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{X}:\varphi\mapsto\varphi(\cdot,P).$
+$\rho_{P}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{X}:\varphi\mapsto\rho_{P}(\varphi)\coloneqq\varphi\,\hat{}(P).$
 For fixed $x\in X$ and $P\in\mathcal{P}$, we also have canonical projection maps
-$\pi_{x}:\mathbb{R}^{X}\to\mathbb{R}:f\mapsto f(x),$ $\pi_{P}:\mathbb{R}^{\mathcal{P}}\to\mathbb{R}:f\mapsto f(P).$
+$\lambda_{x}:\mathbb{R}^{X}\to\mathbb{R}:f\mapsto f(x),$ $\rho_{P}:\mathbb{R}^{\mathcal{P}}\to\mathbb{R}:f\mapsto f(P).$
 […]
@@
 Compact subsets $K\subseteq X$ are closed, hence also Polish. Therefore, the asserted equivalence follows from Theorems 1.5 and 1.7. ∎
-Lastly, a simple but useful lemma that helps understand when we restrict a set of functions to a specific subspace of the domain space, we may always assume that the subspace is closed, as replacing the subspace by its closure has no effect on NIP.
+Lastly, a simple but useful lemma to the effect that a set of continuous functions has NIP over a subset of the domain if and only if it has NIP over the closure of the subset.
 Lemma 1.12.
@@
 This contradicts (i). ∎
-2. Compositional Computation Structures: A structural approach to arbitrary-precision arithmetic
-In this section, we connect function spaces with arbitrary-precision arithmetic computations. We start by summarizing some basic concepts from [ADI+24].
-A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates.
-For a state $v\in L$, the type of a state $v$ is the indexed family
-$\operatorname{tp}(v)=(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}.$
-For each $P\in\mathcal{P}$, we call the value $P(v)$ the $P$-th feature of $v$. A transition of a computation states structure $(L,\mathcal{P})$ is a map $f:L\to L$.
+Historical remarks
+The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles.
+His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
+$C_{p}$-theory is an active field of research in general topology pioneered by A.
+V.
+Arhangel’skiĭ and his students in the 1970s and 1980s [ARK92].
+$C_{p}$-theory has found many applications in model theory and functional analysis.
+For a recent survey, see [TKA11].
+2. Compositional Computation Structures:
+A structural approach to arbitrary-precision computations
+In this section, we connect function spaces with arbitrary-precision computations. We start by summarizing some basic concepts from [ADI+24].
+2.1. Computation States Structures
+A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates, also called features.
+For each $P\in\mathcal{P}$, we call the value $P(v)$ the $P$-th (observable) feature of $v$.
+A transition of a computation states structure $(L,\mathcal{P})$ is a map $f:L\to L$.
+The type of a state $v\in L$ is the indexed family [^5] As in §1.10 above, we regard each predicate on $L$ as named by a specific formal label $P$ (a predicate symbol) which is given the semantic meaning of a bona fide real function $P(\cdot):v\mapsto P(v)$ (the predicate proper) in the computation states structure.
+Such formal symbol $P$ becomes de facto the name of an observable feature having the real value $P(v)$ at any state $v\in L$.
+Whenever $\mathcal{P}$ is used as an index set (e.g., as in the topological product $\mathbb{R}^{\mathcal{P}}$ below), it is regarded as a set of formal symbols.
+This syntactic vs.
+semantic abuse of nomenclature is commonplace in mathematics, but the model-theoretic framework requires contextual awareness of the distinction.
+To disambiguate when necessary, we shall write $P(\cdot)$, implying the semantic meaning of a real function on $L$, in contrast to $P$, meant as a purely syntactic symbol.
+Nevertheless, by a standard abuse of nomenclature in appropriate contexts, we may also treat $P$ as a function on states (i.e., take $P$ to mean $P(\cdot)$ proper).
+$\operatorname{tp}(v)=(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$
+of its features.
 Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable.
-Each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$.
-Important state spaces are $L=\mathbb{R}^{\mathbb{N}}$ and $L=\mathbb{R}^{n}$ for some positive integer $n$, endowed with predicate $P_{i}(v)=v_{i}$, one each for the $i$-th coordinate of $v$.
-We regard the space of types as a topological space, endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$.
-Via the identification $v\mapsto\operatorname{tp}(v)$, the states space $L$ is correspondingly topologized; in particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
+(I.e., each symbol $P\in\mathcal{P}$ denotes a function $P(\cdot)$ taken as a computational primitive on $L$.)
+We assume that each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a topological subspace of $\mathbb{R}^{\mathcal{P}}$ and topologized as such (i.e., $L$ is identified with its forward image $\operatorname{tp}[L]\subseteq\mathbb{R}^{\mathcal{P}}$ under the type map).
+As a topological space, $L$ is thus endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$.
+In particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
+Important state spaces include $L=\mathbb{R}^{\mathbb{N}}$ (resp., $L=\mathbb{R}^{n}$ for some positive integer $n$), endowed with one predicate $P_{i}(v)=v_{i}$ (giving the $i$-th coordinate of $v$) for each $i\in\mathbb{N}$ (resp., for each $0\leq i<n$).
+State subspaces $L\subseteq\mathbb{R}^{\mathbb{N}}$ and $L\subseteq\mathbb{R}^{n}$ are of primary interest in this paper.
+(Any states structure having a finite or countable family of predicates is effectively such a subspace.)
 Definition 2.1.
-Given a computation states structure $(L,\mathcal{P})$, any element of $\mathbb{R}^{\mathcal{P}}$ in the image of $L$ under the map $v\mapsto\operatorname{tp}(v)$ will be called a realized (state) type.
-The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of types of $(L,\mathcal{P})$, denoted $\mathcal{L}$; elements $\xi\in\mathcal{L}$ are called state types.
-Elements of $\mathcal{L}\setminus L$ will be called unrealized types.
+Given a computation states structure $(L,\mathcal{P})$, any type $\operatorname{tp}(v)\in\mathbb{R}^{\mathcal{P}}$ of a state $v\in L$ will be called a realized (state) type.
+The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of (state) types of $(L,\mathcal{P})$, denoted $\mathcal{L}$; elements $\mathbf{v}\in\mathcal{L}$ are called state types.
+Elements $\mathbf{v}\in\mathcal{L}\setminus L$ are unrealized state types.
 Intuitively, state types capture a notion of “limit state”.
-As we combine ideas of model theory [KEI03] and topology [BFT78], we are interested in families of real-valued functions that are pointwise bounded.
-This leads us to the concepts of sizer and shard introduced first in introduced in [ADI+24]:
+For each $P\in\mathcal{P}$, the $P$-th coordinate map (natural projection) $\mathbb{R}^{\mathcal{P}}\to\mathbb{R}:\mathbf{v}=(\mathbf{v}_{Q})_{Q\in\mathcal{P}}\mapsto\mathbf{v}_{P}$, gives a continuous extension to $\mathcal{L}$ of the predicate $P(\cdot)$ on $L$; abusing notation, the extension will still be denoted $P(\cdot)$ (or just $P$).
+Remark 2.2.
+Just as $L$ is identified with a subset of $\mathcal{L}$, but going in the opposite direction, one may regard unrealized state types as “hyper-states” (when such type is obtained as a sequential limit—or ultralimit—of realized state types, unrealized state types may be called “deep states”).
+A comment for readers versed in model theory: at least for the purpose of realizing state types (with parameters in the original states space $L$), the type space $(\mathcal{L},\mathcal{P})$ serves as a “poor man’s” ultraproduct of $(L,\mathcal{P})$ wherein, by the very definition of $\mathcal{L}$, each type $\mathbf{v}$ is realized.
+2.1.1. Bounds on precision and magnitude. Shards
+In order for computation state structures to serve their purpose as models of real-world computing, we must keep in mind that computations involving real-valued quantities can only ever carried out approximately, to a preset degree of precision (i.e., to within acceptable error bounds).
+More subtly, yet equally important in practice, one must also bound the magnitude of real quantities involved in all computations: one cannot possibly store—let alone compute with—arbitrarily large numbers!
+Thus, although the collection $\mathcal{P}$ of predicates may be infinite (possibly, uncountable even), “true” computations must imply choices, at the outset, of bounds for errors and for absolute magnitudes of quantities involved.
+Heuristically, different parts of a computation may require suitably matching (different) error bounds, as well as dealing with quantities of very different sizes.
+Focusing for now on the second of these issues, our results shall generally apply to sets of computations for which arbitrary but fixed bounds for each predicate are set in advance.
+This leads to the concept of shard first introduced in [ADI+24], as per the following definition.
 Definition 2.3.
+Fix a computation states structure $\underline{L}=(L,\mathcal{P})$, with $L\subseteq\mathbb{R}^{\mathcal{P}}$ implicitly identified with its type space.
 A sizer is a family $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$.
-Given a sizer $r_{\bullet}$, let $\mathbb{R}^{[r_{\bullet}]}=\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of a states space $L$ be
-$L[r_{\bullet}]=L\cap\mathbb{R}^{[r_{\bullet}]}.$
-For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$ (a closed, hence compact subset of $\mathbb{R}^{[r_{\bullet}]}$).
-Let also $\mathcal{L}_{\text{sh}}$ be the union of all type-shards as the sizer $r_{\bullet}$ varies.
-In general, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$, and the inclusion may be proper. However, equality holds in the important special case when $\mathcal{P}$ is countable (see [ADI+24]).
+Given a sizer $r_{\bullet}$, let $\mathbb{R}[r_{\bullet}]\coloneqq\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of the states space $L$ be
+$L[r_{\bullet}]=L\cap\mathbb{R}[r_{\bullet}].$
+For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$ (a closed, hence compact subset of $\mathbb{R}[r_{\bullet}]$).
+Let also $\mathcal{L}_{\text{sh}}$ (the space of shard-supported types) be the union of all type-shards as the sizer $r_{\bullet}$ varies.
+Evidently, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$, although the inclusion may be proper.
+However, the equality $\mathcal{L}_{\text{sh}}=\mathcal{L}$ holds in the important special case when $\mathcal{P}$ is (at most) countable (see [ADI+24]).
 2.2. Compositional Computation Structures
@@
 Elements of the semigroup $\Gamma$ are called the computations of the structure $(L,\mathcal{P},\Gamma)$.
-We assume that the identity map id on $L$ is an element of $\Gamma$ (which is thus not merely a semigroup but a monoid of transformations of $L$).
-We topologize $\Gamma$ as a subset of the topological product $L^{L}$, where the “exponent” $L$ serves merely as an index set, but the “base” $L$ is topologized by type; consequently, one may identify $\Gamma$ with a subset of the topological product $(\mathbb{R}^{\mathcal{P}})^{L}$.
-More specifically, $\Gamma$ is identified with a subset of $\mathcal{L}^{L}$, which is a closed subspace of $(\mathbb{R}^{\mathcal{P}})^{L}$.
-Therefore, we have an inclusion $\overline{\Gamma}\subseteq\mathcal{L}^{L}$.
+The state space $L$ is implicitly identified with a subset of the type space $\mathcal{L}\subseteq\mathbb{R}^{\mathcal{P}}$ of $(L,\mathcal{P})$.
+We also assume that the identity map id on $L$ is an element of $\Gamma$ (which is thus not merely a semigroup but a monoid of transformations of $L$).
+We topologize $\Gamma$ as a subset of the topological product $L^{L}$, where (as usual) the “exponent” $L$ serves merely as an index set, but the “base” $L\subseteq\mathbb{R}^{\mathcal{P}}$ has the relative product topology (of pointwise convergence).
+In this manner, the semigroup $\Gamma$ is identified with a subset of the topological product $\mathcal{L}^{L}\subseteq(\mathbb{R}^{\mathcal{P}})^{L}$, which leads to an inclusion $\overline{\Gamma}\subseteq\mathcal{L}^{L}$.
 Elements $\xi\in\overline{\Gamma}$ are called (real-valued) deep computations or ultracomputations.
@@
 A subset $\Delta\subseteq\Gamma$ is $R$-confined if each $\gamma\in\Delta$ is.
+Unlike its subspace $L^{L}$ (a topological semigroup), the space $\mathcal{L}^{L}$ does not have a natural semigroup structure, so ultracomputations are not generally composable.
+Intuitively, ultracomputations are “final” or “ultimate”.
 Proposition 2.5.
@@
 Since deep computations are defined through a combination of topological concepts (namely, topological closure) and structural and model-theoretic concepts (namely, models and types), we will import technology from both topology and model theory.
-2.2. Computability and definability of deep computations and the Extendibility Axiom
-Let $f:L\to\mathcal{L}$ be a function that maps each input state type $(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ to an output state type $(P\circ f(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$.
+Remarks 2.6.
+(For readers versed in model theory.) [^6] Refer to [KEI23] where a general framework for real-valued structures is introduced; see also [DI17, section 6] for a basic tutorial.
   1. (1)
-    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{R}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ and by a finite number of iterations of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
-    $|\varphi_{Q,\,K,\varepsilon}(v)-Q{\circ}f(v)|\leq\epsilon,\qquad\text{for all $v\in K$}.$
-    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model theory of real-valued structures first introduced in [IOV94] for model theory of functional analysis and now standard in model theory (see [KEI03]).
-The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
+    Compositional computational structures as defined above are not real-valued structures in a literal sense, because there is no intrinsic structural manner to ensure that $\Gamma$ is realized as a literal subset of the product space $L^{L}$.
+However, in a semantically transparent manner, any CSS as above may alternatively be seen (per its original definition in [ADI+24]) as a real-valued structure $(L,\mathcal{P},\Gamma,\circ,\operatorname{ev})$, where
+    • $(L,\mathcal{P})$ is a computation states structure,
+    • $(\Gamma,\circ)$ is a semigroup,
+    • $\operatorname{ev}:\Gamma\times L\to L:(\gamma,v)\mapsto\operatorname{ev}(\gamma,v)$ is a semigroup action of $\Gamma$ on $L$ (i.e., $\operatorname{ev}(\gamma\circ\delta,v)=\operatorname{ev}(\gamma,\operatorname{ev}(\delta,v))$ for $\gamma,\delta\in\Gamma$ and $v\in L$).
+    Both definitions are reconciled upon identifying $\gamma\in\Gamma$ with the map $(\gamma,v)\mapsto\operatorname{ev}(\gamma,v)$.
   2. (2)
-    We will say that $f$ is computable if it is definable in the sense defined above under (1), but without the use of the sup/inf operators; in other words, if for every choice of $Q,K,\varepsilon$, the approximation function $\varphi_{Q,K,\varepsilon}:L\to\mathcal{L}$ can be constructed without any use of sup or inf operators.
-This is quantifier-free definability (i.e., definability as given by the preceding paragraph, but without use of quantifiers), which, from a logic viewpoint, corresponds to computability (the presence of the quantifiers $\exists$ and $\forall$ are the reason behind the undecidability of first-order logic).
-It is shown in [ADI+24] that:
+    In the preceding sense of real-valued structures, the class of compositional computation structures (for a fixed predicate symbol collection $\mathcal{P}$) is first-order elementary.
+The limited notion of state type we have introduced (essentially, quantifier-free state types with parameters of sort $L$ only) may be enriched to state types involving more formulas (e.g., with parameters in $\Gamma$, or including quantified formulas), and extended to “transform-types” (i.e., types of sort $\Gamma$).
+In particular, given a structure $\underline{C}=(L,\mathcal{P},\Gamma,\circ,\operatorname{ev})$, all types are realized in some ultrapower $\underline{C}^{(\mathcal{U})}$ of $\underline{C}$ (constructed in the usual manner detailed in the references above).
+[^7] By contrast (cf., Remark 2.2), the ultrapower construction is not necessary to realize state types in a computation states structures.
+This approach to spaces of types was introduced in Banach space theory by Krivine [KRI74, KM81, KRI72].
+2.3. Computability of deep computations and the Extendibility Axiom
+Let $(L,\mathcal{P})$ be a computation structure with countable predicate collection $\mathcal{P}$.
   1. (1)
-    For a definable $f:L\to\mathcal{L}$, the approximating functions $\varphi_{Q,\,K,\varepsilon}$ may be taken to be polynomials of the input features, and
+    A computable predicate is any real function $\varphi:L\to\mathbb{R}$ whose restriction to an arbitrary shard is uniformly approximable by polynomials (with real coefficients) in the features $P(\cdot)$.
   2. (2)
-    Definable transforms $f:L\to\mathcal{L}$ are precisely those that extend to continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$.
-To summarize, a function $f:L\to\mathcal{L}$ is computable if and only if it is definable if and only if it is polynomially approximable if and only it can be extended to a continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$.
-This is the reason for the following definition.
-Definition 2.5.
-We say that a CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom if for all $\gamma\in\Gamma$, there is $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous.
-We refer to $\tilde{\gamma}$ as a free extension of $\gamma$.
-For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$.
-For any $\Delta\subseteq\Gamma$, let $\tilde{\Delta}$ denote $\{\tilde{\gamma}\colon\gamma\in\Delta\}$.
+    A transform $f:L\to\mathcal{L}$ is computable if, for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a computable predicate.
+It is shown in [ADI+24] that computable transforms $f:L\to\mathcal{L}$ are precisely those that extend to continuous maps $\tilde{f}$$\mathcal{L}\to\mathcal{L}$.
+To summarize, for a function $f:L\to\mathcal{L}$, the following conditions are equivalent:
+  • $f$ is computable
+  • $f$ has is uniformly polynomially approximable features on shards
+  • $f$ extends to a continuous map $\mathcal{L}\to\mathcal{L}$.
+These equivalences motivate the following definition:
+A computation $\gamma\in\Gamma$ is extendable if it has some extension $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that, for every sizer $r_{\bullet}$, there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous.
+Because $\tilde{\gamma}$ extends $\gamma$ continuously on each compact type-shard, we call it the $K$-extension of $\gamma$; such extension is evidently unique (if it exists).
+Abusing notation, for any $\Delta\subseteq\Gamma$ consisting of extendable computations, the set $\{\tilde{\gamma}\colon\gamma\in\Delta\}$ will be denoted $\tilde{\Delta}$.
+Extendibility Axiom.
+Every computation $\gamma\in\Gamma$ is extendable.
 For a more detailed discussion of the Extendibility Axiom, we refer the reader to [ADI+24].
+For the rest of the paper, the Extendibility Axiom is imposed on all CSSs.
 2.4. Newton’s method as a CCS
-Let $p(z)$ be a non-constant polynomial with complex coefficients. We say that $(L_{p},\mathcal{P},\Gamma_{p})$ is Newton’s method for $p(z)$ if:
-  • $L_{p}$ is the set of all $z\in\mathbb{C}$ such that there exists an open neighborhood $U$ of $z$ such that every sequence in $\{N_{p}^{n}:n\in\mathbb{N}\}$ has a subsequence that converges uniformly on compact subsets of $U$, where
-    $N_{p}(z)=z-\frac{p(z)}{p^{\prime}(z)}$
-    and $N_{p}^{n}:=N_{p}\circ N_{p}\circ\cdots\circ N_{p}$ is the nth-iteration of $N_{p}$.
-  • $\mathcal{P}:=\{P_{1},P_{2},P_{3}\}$ where
-    $P_{1}(z)=\frac{2{\rm Re}(z)}{|z|^{2}+1},$
-    $P_{2}(z)=\frac{2{\rm Im}(z)}{|z|^{2}+1},$
-    $P_{3}(z)=\frac{|z|^{2}-1}{|z|^{2}+1}.$
-  • $\Gamma_{p}:=\{N_{p}^{n}:n\in\mathbb{N}\}$.
-Remarks 2.6.
-  1. (1)
-    The set $L_{p}$ is known as the Fatou set of $N_{p}$ (see section 2 in [BLA84]). Its complement $\mathbb{C}\backslash L_{p}$, is known as the Julia set of $N_{p}$.
-  2. (2)
-    The map $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$ is the stereographic projection into the Riemann Sphere $S^{2}$.
-  3. (3)
-    The set $L_{p}$ is open and dense in $\mathbb{C}$ ([BLA84], Corollary 4.6).
-Hence, its closure $\mathcal{L}$ in $\mathbb{R}^{\mathcal{P}}$ is the Riemann sphere $S^{2}$ (i.e., the extended complex plane).
-  4. (4)
-    The set $L_{p}$ is completely invariant under iterations of $N_{p}$, i.e., $N_{p}(L_{p})=L_{p}$ and $N_{p}^{-1}(L_{p})=L_{p}$.
-This implies that all iterations $N_{p}^{n}:L_{p}\rightarrow L_{p}$ are transition maps.
-  5. (5)
-    $\Gamma_{p}$ is the semigroup generated by $\{N_{p}\}$. Thus, $(L_{p},\mathcal{P},\Gamma_{p})$ is a CCS.
-Newton’s method is an iterative method that is used to approximate a root of $p(z)$.
-The map $N_{p}(z)$ defined above is known as Newton’s map.
-The method consists of taking an initial guess $z_{0}\in\mathbb{C}$ and iterating the rational map $N_{p}$ to obtain a sequence given by
-$z_{n+1}=N_{p}(z_{n})=z_{n}-\frac{p(z_{n})}{p^{\prime}(z_{n})}$
-For each root $r\in\mathbb{C}$ of $p(z)$, there exists an $\varepsilon>0$ such that for any initial guess $z_{0}$ in the $\varepsilon$-ball centered at $r$, Newton’s iteration converges to $r$ (provided $p^{\prime}(r)\neq 0$) and the convergence is quadratic in that case, meaning the error at each step is roughly squared, causing the number of correct digits to double, leading to fast convergence.
-Given a root $r$ of $p(z)$, the set $B_{r}=\{z\in\mathbb{C}:\lim_{n\rightarrow\infty}N_{p}^{n}(z)=r\}$ is an open set called the basin of $r$.
-However, Newton’s method can fail to converge to any root for some choices of $z_{0}$.
-For example, consider the polynomial $p(z)=z^{3}-2z+2$.
-The Newton map is given by
-$N_{p}(z)=z-\frac{z^{3}-2z+2}{3z^{2}-2}=\frac{2z^{3}-2}{3z^{2}-2}$
-Notice that taking $z_{0}=0$ as an initial guess will yield the sequence $0,1,0,1,0,1,\dots$ that oscillates between 0 and $1$ but none of them are roots of $p(z)$.
-Another more chaotic way Newton’s method can fail to converge is when the sequence of iterations has no convergent subsequence.
-The set of such points, i.e.
-the Julia set associated to $N_{p}$, is typically a fractal.
-This can be visualized by adding a dash of color: let us give each complex number $z_{0}$ a color $(R,G,B)$ where $R,G,B\in[0,1]$ (so that $(1,0,0)$ is red, $(0,1,0)$ is green, $(0,0,1)$ is blue and $(0.5,0,0.5)$ is a light purple, for example).
-The values of $R$, $G$ and $B$ are determined by looking at the image of said number at each stage of the iteration, $N_{p}^{n}(z_{0})$, and computing the current distance to each of the roots of $p(z)$; so $R=1/d_{r}$ where $d_{r}$ is the positive distance to the root which is colored red, and so on.
-In this way, the roots themselves are colored red, green, and blue, and every other point gets a mix of the three colors.
-As the number of iterations increases, each point gets a sharper color, as the sequence of images $\{N_{p}^{n}(z_{0})\}_{n=1}^{\infty}$ converges to one of the three roots.
-At each stage, the complex plane looks as if out of focus because the coloring function is continuous.
-As the reader can see in Figure 1, the points at the boundary of each color class form the famous Newton’s fractal (of which, interestingly, Newton was unaware).
+We formalize the Newton-Raphson root-finding method as a CCS [GAL00], [BFB16, Section 2.3].
+2.4.1. The extended complex plane
+Let $\hat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ be the extended complex plane (the Alexandroff —one-point— compactification of $\mathbb{C}$).
+We regard the unit sphere $S^{2}\subseteq\mathbb{R}^{3}$ as the Riemann sphere and endowed with the stereographic projection map $S^{2}\to\hat{\mathbb{C}}$ —a topological homeomorphism whose inverse $\hat{\mathbb{C}}\rightarrow S^{2}$ is given by $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$, where $(P_{1}(\infty),P_{2}(\infty),P_{3}(\infty))=(0,0,1)$ and, for $z\neq\infty$,
+$P_{1}(z)$ $=\frac{2{\Re}(z)}{|z|^{2}+1},$ $P_{2}(z)$ $=\frac{2{\Im}(z)}{|z|^{2}+1},$ $P_{3}(z)$ $=\frac{|z|^{2}-1}{|z|^{2}+1}.$
+The extended complex plane $\hat{\mathbb{C}}$ is thus endowed with the real predicates $P_{1},P_{2},P_{3}$ with values in [1,1][-1,1]; its topology is metrized by the pullback of the usual Euclidean metric on $S^{2}$.
+2.4.2. Newton maps
+Let $p=p(z)$ be a non-constant polynomial with complex coefficients.
+[^8] The entire discussion of Newton CCSs applies, mutatis mutandis, to rational maps $p$.
+The Newton method step (or Newton map) for $p$ is the map
+(2.7) $N_{p}:z\mapsto z-\frac{p(z)}{p^{\prime}(z)}.$
+Now, if one wishes to interpret such map as a (complex-valued) function of a complex variable, its natural domain $\mathbb{C}\setminus\mathcal{Z}(p^{\prime})$ excludes all roots of $p^{\prime}$.
+If one wishes to iterate the Newton map indefinitely, the natural domain is
+$L_{p}\coloneqq\mathbb{C}\setminus\bigcup_{n\in\mathbb{N}}N_{p}^{-n}[\mathcal{Z}(p^{\prime})]=\{z\in\mathbb{C}:\text{$N_{p}^{(n)}(z)\notin\mathcal{Z}(p^{\prime})$ for all $n\in\mathbb{N}$}\}.$
+On the other hand, any rational map extends continuously to an analytic (meromorphic) function on the extended complex plane, so $N_{p}$ naturally defines a map $\hat{\mathbb{C}}\to\hat{\mathbb{C}}$.
+Successive iterations of the Newton-Raphson method (i.e., compositional powers $N_{p}^{(n)}$) may be regarded as computations of a suitable CCS whose states space is perhaps $L_{p}\subseteq\mathbb{C}$, or perhaps the full extended complex plane $\hat{\mathbb{C}}$.
+In the former case, it is natural to use the predicates $X:z\mapsto\Re(z)$ and $Y:z\mapsto\Im z$; in the latter, the predicates $P_{1},P_{2},P_{3}$ are necessary because $X,Y$ are not defined at $\infty$.
+An important (if somewhat obvious) point is that the choice of predicates on a CCS greatly affect the topologies of itself and of its type space.
+Thus, the type space of $(L_{p},\{X,Y\})$ above is $\mathbb{R}^{2}$ ($\cong\mathbb{C}$), while the type space of $(\hat{\mathbb{C}},\{P_{1},P_{2},P_{3}\})$ is $S^{2}$ $\cong\hat{\mathbb{C}}$.
+All $\{P_{1},P_{2},P_{3}\}$-types are realized in $\hat{\mathbb{C}}$, but not all $\{X,Y\}$-types are realized in $L_{p}$ (let alone all $\{P_{1},P_{2},P_{3}\}$-types).
+2.4.3. Newton CCSs
+Per the discusion above, we choose to discuss the Newton-Raphson method as a CCS whose states space is the extended complex plane $\hat{\mathbb{C}}$.
+Let $p=p(z)$ be a nonconstant complex polynomial. We say that $(\hat{\mathbb{C}},\mathcal{P},\Gamma_{p})$ is Newton’s method for $p(z)$ if:
+  • $\mathcal{P}:=\{P_{1},P_{2},P_{3}\}$ are the coordinates of inverse steregraphic projection (as above);
+  • $\Gamma_{p}:=\{N_{p}^{(n)}:n\in\mathbb{N}\}$ consists of iterates of the Newton map $N_{p}$ in equation (2.7).
+Given an initial estimate $z_{0}$, one hopes for the sequence $N_{p}^{(\bullet)}(z_{0})\coloneqq\bigl(N_{p}^{(n)}(z_{0})\bigr)_{n}$ of Newton steps applied to $z_{0}$ to converge to a root $r$ of $p$.
+However, Newton’s method fails to converge to any root in many cases.
+For example, consider the polynomial $p(z)=z^{3}-2z+2$ with Newton map
+$N_{p}(z)=z-\frac{z^{3}-2z+2}{3z^{2}-2}=\frac{2z^{3}-2}{3z^{2}-2}.$
+[…]
+For any $n\in\mathbb{N}$ we define a coloring of $\mathbb{C}$ in the natural way: color each $z_{0}\in\mathbb{C}$ using the standard color of $z_{n}\coloneqq N_{p}^{(n)}(z_{0})$.
+At any stage, particularly early ones, the coloring looks “fuzzy” (out of focus) because the coloring function is continuous; however, as $n$ increases, the coloring becomes sharper.
+For one thing, for many $z$ in the Fatou set, the sequence $(z_{n})$ converges to one of the three roots; the colors of $(z_{n})$ tends towards (a single one of) red, green or blue correspondingly.
+As immediately recognizable in Figure 1, such $z_{0}$ lie in large (open connected) regions.
+However, the Fatou set also includes points $z_{0}$ for which $(z_{n})$ diverges, but has some subsequence converging uniformly in some compact neighborhood of $z_{0}$; those points form dark teal regions.
+The Julia set at boundary of the Fatou set is Newton’s fractal proper.
 (a) After 1 iteration
@@
 (c) After 100 iterations
-Figure 1. Newton’s method approximating $p(z)=z^{3}-2z+2$. Notice the regions of divergence.
-Another example of a Newton’s fractal is for $p(z)=z^{3}-1$. The roots of $p(z)$ are the 3rd roots of unity and the Newton map is given by:
-$N_{p}(z)=z-\frac{z^{3}-1}{3z^{2}}=\frac{2z^{3}+1}{3z^{2}}$
-In this case, there are three basins of attraction (one for each root) and the complement of their union is the Julia set, i.e., the common boundary.
+Figure 1.
+Behavior of iterates of the Newton map for $p(z)=z^{3}-2z+2$.
+In (C), the “color fuzzines” of (A) and (B) has already visually disappeared; only sharply defined open red, green and blue regions (of convergence towards a specific root), as well as dark teal open regions (of divergence) remain visually identifiable: their union (the Fatou set of $p$) is dense in the limit, and the Newton fractal of $p$ is the closed nowhere dense complement (the Julia set of $p$).
+Consider next the $3$-cyclotomic polynomial $p(z)=z^{3}-1$. The roots of $p(z)$ are the cube roots of unity; the Newton map is
+$N_{p}:z\mapsto z-\frac{z^{3}-1}{3z^{2}}=\frac{2z^{3}+1}{3z^{2}}.$
+In Figure 2(C), one sees a Fatou set consisting of solid-color (disconnected) open sets, one in each of the colors red (for the root $r_{0}=1$), green and blue (for $r_{\pm 1}=\exp(\pm 2\pi i/3)$, resp.), where the iterates sequence $(z_{n})=(p^{(n)}(z_{0}))_{n}$ starting with any colored point $z_{0}$ converges to the corresponding root.
+(Thus, each solid-color region is the basin of attraction of the respective root).
 (a) After 1 iteration
@@
 Proposition 2.8.
-Let $p(z)$ be a non-constant polynomial. $(L_{p},\mathcal{P},\Gamma_{p})$ satisfies the Extendibility Axiom.
+If $p(z)$ is a non-constant polynomial, then $(L_{p},\mathcal{P},\Gamma_{p})$ satisfies the Extendibility Axiom.
 Proof.
@@
 Example 2.9.
-[…]
-$P_{1}(x)=\frac{2x}{x^{2}+1},$
-$P_{2}(x)=\frac{x^{2}-1}{x^{2}+1}.$
-Let $\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$ where
-$N_{p}(x)=\frac{x^{2}+a}{2x}.$
-As before, $(L,\mathcal{P},\Gamma)$ is a CCS.
-Note that $\mathcal{L}=S^{1}$ and that each iterate $N_{p}^{n}$ can be continuously extended to the extended real line $\mathbb{R}\cup\{\infty\}$, i.e., $\mathcal{L}$.
-For example,
-$\tilde{N_{p}}(x)=\begin{cases}\frac{x^{2}+a}{2x},&\text{if }x\in L;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
-For every initial guess $x\in L$, the limit $f(x)=\lim_{n\rightarrow\infty}N_{p}^{n}(x)$ converges pointwise to one of the roots. Moreover,
-$f(x)=\begin{cases}\sqrt{a},&\text{if }x>0;\\ -\sqrt{a},&\text{if }x<0.\end{cases}$
-Notice that $f$ can be extended to $\mathcal{L}$ by
-$\tilde{f}(x)=\begin{cases}\sqrt{a},&\text{if }x>0;\\ -\sqrt{a},&\text{if }x<0;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
-However, $\tilde{f}$$\mathcal{L}\rightarrow\mathcal{L}$ is not continuous.
-The set $\overline{\tilde{\Gamma}}$ of deep computations is $\tilde{\Gamma}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
+Computation of square roots.
+Let $a$ be a positive real number and $p(x)=x^{2}-a$.
+Let $\mathbb{R}\to S^{1}:x\mapsto(P_{1}(x),P_{2}(x))$ be the (inverse) stereographic projection of the real line onto $S^{1}\subseteq\mathbb{R}^{2}$, i.e.,
+$P_{1}(x)$ $=\frac{2x}{x^{2}+1},$ $P_{2}(x)$ $=\frac{x^{2}-1}{x^{2}+1},$
+extended to a projective transformation of $\mathbb{R}\mathbb{P}^{1}=\mathbb{R}\cup\{\infty\}$ by $\infty\mapsto(P_{1}(\infty),P_{2}(\infty))=(0,1)$.
+Let $\mathcal{P}=(P_{1},P_{2})$ and $\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$, where $N_{p}:\mathbb{R}\mathbb{P}^{1}\rightarrow\mathbb{R}\mathbb{P}^{1}$ is given by
+$x\mapsto x-\frac{x^{2}-a}{2x}=\begin{cases}\frac{x^{2}+a}{2x}&(x\neq 0,\infty),\\ \infty&(x=0,\infty).\end{cases}$
+As before, $(L,\mathcal{P},\Gamma)$ is a CCS with $L\coloneqq\mathbb{R}\mathbb{P}^{1}\setminus\{0,\infty\}=\mathbb{R}\setminus\{0\}$ whose type space is $\mathcal{L}=S^{1}\cong\mathbb{R}\mathbb{P}^{1}$.
+In fact (with the identification $\mathcal{L}\cong\mathbb{R}\mathbb{P}^{1}$), the (full) sequence if iterates of the Newton map has a pointwise limit $\mathfrak{f}:\mathcal{L}\to\mathcal{L}$, namely
+$\mathfrak{f}(x)=\lim_{n\rightarrow\infty}N_{p}^{n}(x)=\begin{cases}\sqrt{a}&(x>0),\\ -\sqrt{a}&(x<0),\\ \infty&(x=0,\infty).\end{cases}$
+Thus, $\mathfrak{f}$ is discontinuous at $x=0,\infty$. The set of deep computations is $\{N_{p}^{n}\}_{n}\cup\{\mathfrak{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
 Example 2.10.
 Newton’s method for $p(z)=z^{3}-2z+2$.
-Let $r_{1}$, $r_{2}$ and $r_{3}$ be the three roots of $p(z)$.
-Let $B_{1}$, $B_{2}$ and $B_{3}$ be their respective basins.
-Let $B$ be the basin of the attractive cycle $0,1,0,1,\dots$.
-Then, $L_{p}=B\cup\bigcup_{i=1}^{3}B_{i}$.
-Notice that $N_{p}^{n}$ does not converge pointwise.
-However, the subsequences $N_{p}^{2n}$ and $N_{p}^{2n+1}$ are pointwise convergent to functions $f_{1}$ and $f_{2}$ respectively.
-$f_{1}$ and $f_{2}$ are two distinct deep computations.
-Note that for $z\notin L_{p}$, no subsequence of $\tilde{N_{p}^{n}}(z)$ converges to a complex number.
-However, since $\mathcal{L}=S^{2}$ is compact there is a subsequence of $\tilde{N_{p}^{n}}(z)$ that converges to $\infty$.
-We can extend $f_{i}:L_{p}\rightarrow\mathcal{L}$ to $\tilde{f_{i}}$$\mathcal{L}\rightarrow\mathcal{L}$ by:
-[…]
-Again, note that $\tilde{f_{i}}$ for $i=1,2$ are not continuous and that $\tilde{f_{i}}\in\overline{\tilde{\Gamma}}$.
+Let $B_{i}$, $B_{\pm 1}$ be the basin of the root $r_{i}$ ($i=0,\pm 1$), and let $B$ be the basin of the attractive cycle 0, 1, 0, 1, ….
+It may be shown that the Fatou set of $N_{p}$ is $F_{p}=B\cup\bigcup_{i=0,\pm 1}B_{i}$, each of these four sets being nonempty open.
+Since $B$ is nonempty, $N_{p}^{n}$ does not converge pointwise; however, the subsequences $N_{p}^{2n}$ and $N_{p}^{2n+1}$ are pointwise convergent to two distinct deep computations $f_{0}$ (constant 0 on $B$) and $f_{1}$ (constant $1$ on BB).
+[^10] As remarked above, $\infty$ is a repelling fixed point of $N_{p}$, hence $f_{0}(\infty)=\infty=f_{1}(\infty)$.
+The behavior of the iterates $N_{p}$ on the Julia set $\mathbb{C}\setminus L_{p}$ is chaotic
 2.5. Finite precision threshold classifiers as a CCS
@@
 Let $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$ and $\Gamma$ be the semigroup generated by $\Delta$.
-Since the sets $\{x\in 2^{\mathbb{N}}:x|_{|w|}=w\}$ are open and closed in $2^{\mathbb{N}}$, then the transitions $\psi_{w}$ are all continuous.
+Since the sets $\{x\in 2^{\mathbb{N}}:x|_{|w|}=w\}$ are open and closed in $2^{\mathbb{N}}$, we see that the transitions $\psi_{w}$ are all continuous.
 In particular, $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom.
@@
 Theorem 3.1.
-Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.3) satisfying the Extendibility Axiom (Definition 2.5) with $\mathcal{P}$ countable.
+Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.4) satisfying the Extendibility Axiom (Definition LABEL:D:extendibility_axiom) with $\mathcal{P}$ countable.
 Let $R$ be an exhaustive collection of sizers.
@@
   2. (ii)
-    $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and $r_{\bullet}\in R$; that is, for all $P\in\mathcal{P}$, $r_{\bullet}\in R$, $a<b$, $\{\gamma_{n}\}_{n\in\mathbb{N}}\subseteq\Delta$ there are finite disjoint $E,F\subseteq\mathbb{N}$ such that L[r_]∩⋂_nE(π_Pγ_n)^-1(-,a]∩⋂_nF(π_Pγ_n)^-1[b,)=.
+    $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and $r_{\bullet}\in R$; that is, for all $P\in\mathcal{P}$, $r_{\bullet}\in R$, $a<b$, $\{\gamma_{n}\}_{n\in\mathbb{N}}\subseteq\Delta$ there are finite disjoint $E,F\subseteq\mathbb{N}$ such that
+    $L[r_{\bullet}]\cap\bigcap_{n\in E}(\pi_{P}\circ\gamma_{n})^{-1}(-\infty,a]\cap\bigcap_{n\in F}(\pi_{P}\circ\gamma_{n})^{-1}[b,\infty)=\emptyset.$
 Moreover, if any (hence all) of the preceding conditions hold, then every deep computation $f\in\overline{\Delta}$ can be extended to a Baire-1 function on shards, i.e., there is $\tilde{f}:\mathcal{L}_{\text{sh}}\rightarrow\mathcal{L}_{\text{sh}}$ such that $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$ for all $r_{\bullet}\in R$.
@@
 Hence, for all $r_{\bullet}\in R$ we have $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$.
-That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where topological closures coincide with sequential closures, see Theorem 3F in [BFT78] or Theorem 4.1 in [DEB13]).
+That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where any point in the closure of a set $A$ is the limit of a sequence of points in $A$; see Theorem 3F in [BFT78] or Theorem 4.1 in [DEB13]).
 ∎
@@
     Alexandroff Duplicate.
-Let $K$ be any compact metric space and consider the Polish space $X=K\sqcup C(K)$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$.
-For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={0, if x∈Kx(a), if x∈C(K); g_a^1(x)={δa(x), if x∈K;x(a), if x∈C(K).
+Let $K$ be any compact metric space and consider the Polish space $X=K\sqcup C(K)$, i.e., the disjoint union of $C(K)$ (continuous functions on $K$ with the supremum norm topology) and $K$.
+For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows:
+    $g_{a}^{0}(x)=\begin{cases}0,&\text{if }x\in K\\ x(a),&\text{if }x\in C(K);\end{cases}$
+    $g_{a}^{1}(x)=\begin{cases}\delta_{a}(x),&\text{if }x\in K;\\ x(a),&\text{if }x\in C(K).\end{cases}$
     Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$.
-Observe that all points $g_{a}^{1}$ are isolated and that open neighborhoods of $g_{a_{0}}^{0}$ are of the form $g_{a}^{i}:a\in U,i\in\{0,1\}\}\backslash\{g_{a}^{1}:a\in F\}$ where $U\subseteq K$ is an open neighborhood of $a_{0}$ and $F\subseteq K$ is a finite set.
+Observe that all points $g_{a}^{1}$ are isolated and that open neighborhoods of $g_{a_{0}}^{0}$ are of the form $\{g_{a}^{i}:a\in U,i\in\{0,1\}\}\setminus\{g_{a}^{1}:a\in F\}$ where $U\subseteq K$ is an open neighborhood of $a_{0}$ and $F\subseteq K$ is a finite set.
 Another abstract way in which this space is presented is as the space $K\times\{0,1\}$ whose basic open neighborhoods are given as before, identifying $(a,0)\mapsto g_{a}^{0}$ and $(a,1)\mapsto g_{a}^{1}$.
@@
 Notice that $D(K)$ is a first countable Rosenthal compactum.
-It is not separable if $K$ is uncountable, thus we typically study the interesting case when $K=2^{\mathbb{N}}$.
+It is not separable if $K$ is uncountable; thus, we typically study the interesting case when $K=2^{\mathbb{N}}$.
 As with the Alexandroff compactification $A(2^{\mathbb{N}})$, we can make the space $D(2^{\mathbb{N}})$ separable by adding a countable set.
@@
 For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s.
-Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0, if x¡at;1/2, if at≤x ≤bt;1, if bt¡x.
+Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by
+    $h_{t}(x)=\begin{cases}0,&\text{if }x<a_{t};\\ 1/2,&\text{if }a_{t}\leq x\leq b_{t};\\ 1,&\text{if }b_{t}<x.\end{cases}$
     Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$.
@@
   3. (3)
-    Let $(L,\mathcal{P},\Gamma)$ be the CCS given by $L=2^{\mathbb{N}}$, $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ and $\Gamma$ is the semigroup generated by $\Delta=\{\gamma_{t}:t\in 2^{<\mathbb{N}}\}$, where $P_{n}:2^{\mathbb{N}}\rightarrow\{0,1\}$ is the projection map $P_{n}(x)=x(n)$ and $\gamma_{t}:L\rightarrow L$ is given by γ_t(x)={0, if lext0;(01), if t0lexx ≤lext1;1, if t1¡lexx.
+    Let $(L,\mathcal{P},\Gamma)$ be the CCS given by $L=2^{\mathbb{N}}$, $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ and $\Gamma$ is the semigroup generated by $\Delta=\{\gamma_{t}:t\in 2^{<\mathbb{N}}\}$, where $P_{n}:2^{\mathbb{N}}\rightarrow\{0,1\}$ is the projection map $P_{n}(x)=x(n)$ and $\gamma_{t}:L\rightarrow L$ is given by
+    $\gamma_{t}(x)=\begin{cases}0^{\infty},&\text{if }x<_{\text{lex}}t^{\frown}0^{\infty};\\ (01)^{\infty},&\text{if }t^{\frown}0^{\infty}\leq_{\text{lex}}x\leq_{\text{lex}}t^{\frown}1^{\infty};\\ 1^{\infty},&\text{if }t^{\frown}1^{\infty}<_{\text{lex}}x.\end{cases}$
     where $(01)^{\infty}$ denotes the sequence of alternating bits: $010101\cdots$.
@@
 We do not assume that $\mathcal{P}$ is countable.
-The main results of the section are Theorem 4.8 (connecting NIP and Monte Carlo computability) and 4.14 (connecting Talagrand stability and Monte Carlo computability).
+(The countability assumption on $\mathcal{P}$ played a crucial role in the proof of Theorem 1.11, as it makes $\mathbb{R}^{\mathcal{P}}$ a Polish space.) The main results of the section are Theorem 4.8 (connecting NIP and Monte Carlo computability) and 4.14 (connecting Talagrand stability and Monte Carlo computability).
 Fundamental in this section is a measure-theoretic version of Theorem 1.11, namely, Theorem 4.5.
@@
 4.1. NIP and Monte Carlo computability of deep computations
-The raison d’être of the Baire class-1 functions is to have with a class of functions that are obtained as equential limit points of continuous functions.
+The raison d’être of the Baire class 1 functions is to have a class of functions that are obtained as sequential limit points of continuous functions.
 By Fact 1.2, for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$.
@@
 In the context of deep computations, we are interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself.
-There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the sets $\pi_{P}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open and $P\in\mathcal{P}$).
-Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller.
+In the product space $\mathbb{R}^{\mathcal{P}}$, we have two natural $\sigma$-algebras: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the set $\pi_{i}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open).
+In general, the cylinder $\sigma$-algebra is strictly smaller, but when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide.
 We will use the cylinder $\sigma$-algebra to define universally measurable maps $f:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$.
@@
 The preceding proposition says that a transition map is universally measurable if and only if it is universally measurable on all its features; in other words, we can check measurability of a transition just by checking measurability feature by feature.
-This is the same as in the Baire class-1 case (compare with Proposition 1.10).
-The main result in section 3 is that, as long as we work with countably many features, PAC-learning (or NIP) corresponds to relative compactness in the space of Baire class-1 functions.
+This is the same as in the Baire class 1 case (compare with Proposition 1.10).
+The main result in section 3 is that, as long as we work with countably many features, PAC-learning (or NIP) corresponds to relative compactness in the space of Baire class 1 functions.
 The following result (which does not assume countability of the number of features) gives an analogous characterization of the NIP in terms of universal measurability:
@@
 Remark 4.7.
-Condition (2) of Theorem 4.5 shows that to study measure-theoretic versions of NIP, we need only consider compact subsets of $X$.
-Now, every Radon measure on a compact space is finite; hence, by proper normalization, it can be treated as a probability measure.
-Therefore, in the context of Monte Carlo measurability, we focus on Radon probability measures rather than general Radon measures.
+Every Radon measure on a compact space (e.g., $\mathcal{L}_{\text{sh}}$) is finite; hence, by proper normalization, it can be treated as a probability measure.
+For this reason, in the context of Monte Carlo computability (4.6), we restrict our attention to Radon probability measures instead of general Radon measures.
 Theorem 4.8.
@@
   4. (4)
-    $(X,\Sigma,\mu)$ is inner regular with respect to $\mathcal{K}$, i.e., for all $E\in\Sigma$ μ(E)=sup{μ(K):$K$ and KE}.
+    $(X,\Sigma,\mu)$ is inner regular with respect to $\mathcal{K}$, i.e., for all $E\in\Sigma$
+    $\mu(E)=\sup\{\mu(K):K\in\mathcal{K}\text{ and }K\subseteq E\}.$
 (In particular, if $X$ is compact Hausdorff, $\mu$ is a Radon probability measure on $X$, $\Sigma$ is the completion of the Borel $\sigma$-algebra by $\mu$, and $\mathcal{K}$ is the collection of compact subsets of $X$, all these conditions hold).
@@
 However, by the Extendibility Axiom, finitary computations are continuous, i.e., if $A$ is a set of computations, then $A\subseteq C_{p}(X)$.
-The question of whether we can remove the set-theoretic assumption in Theorem 4.15 when $A\subseteq C_{p}(X)$ (instead of $A\subseteq M_{r}(X)$) remains open.
+5. Conclusion
+We have seen how the tools of model theory and the topology of function spaces can give us new insights into classifying the difficulty of computations.
+In future papers, we plan to extend this research in a variety of directions.
+A direction of particular interest to us is replacing classical computation by quantum computation.
+Topologically, our results so far are not optimal because of their restriction to Polish spaces.
+We expect to extend most of them to $K$-countably determined (Lindelöf $\Sigma$) spaces or $K$-analytic spaces, which have previously been used in functional analysis — see e.g., Kakol et al [KKL+02].
+Our restricting our considerations to continuous and Baire class 1 functions can be relaxed by applying more high-powered model-theoretic and topological techniques.
+This is work in progress.
+We have hinted at applications to deep equilibria; see [ADI+24].
+In another direction, we have hinted at the role additional set-theoretic axioms may play.
+Since real world computations are finite and hence do not depend on other than the usual set-theoretic axioms, in principle we can assume these axioms and then dispense with them; however, this is delicate, since many of our techniques involve infinite sets so we have to check that our assertions are sufficiently simple so as to be absolute, i.e., unaffected by additional axioms.
+We hope that this paper, the preceding one [ADI+24], and the future ones will stimulate interactions among these three seemingly unrelated fields of research.